### v1

# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

Abstract.
Efficient use of GPU memory is essential for high throughput LLM inference. Prior systems reserved memory for the KV-cache ahead-of-time, resulting in wasted capacity due to internal fragmentation. Inspired by OS-based virtual memory systems, vLLM proposed PagedAttention to enable dynamic memory allocation for KV-cache. This approach eliminates fragmentation, enabling high-throughput LLM serving with larger batch sizes. However, to be able to allocate physical memory dynamically, PagedAttention changes the layout of KV-cache from contiguous virtual memory to non-contiguous virtual memory. This change requires attention kernels to be rewritten to support paging, and serving framework to implement a memory manager. Thus, the PagedAttention model leads to software complexity, portability issues, redundancy and inefficiency.
In this paper, we propose vAttention for dynamic KV-cache memory management. In contrast to PagedAttention, vAttention retains KV-cache in contiguous virtual memory and leverages low-level system support for demand paging, that already exists, to enable on-demand physical memory allocation. Thus, vAttention unburdens the attention kernel developer from having to explicitly support paging and avoids re-implementation of memory management in the serving framework. We show that vAttention enables seamless dynamic memory management for unchanged implementations of various attention kernels. vAttention also generates tokens up to 1.97×1.97$\times$1.97 × faster than vLLM, while processing input prompts up to 3.92×3.92$\times$3.92 × and 1.45×1.45$\times$1.45 × faster than the PagedAttention variants of FlashAttention and FlashInfer.
copyright: none
1. Introduction
Large Language Models (LLMs) are being deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (OpenAI, 2023; Chowdhery et al., 2022; cla, [n. d.]; bin, [n. d.]; bar, [n. d.]; git, [n. d.]; ama, [n. d.]; rep, [n. d.]). Optimizing LLM inference has thus become important (Agrawal et al., 2023; Holmes et al., 2024; Patel et al., 2023; Zhong et al., 2024; Hu et al., 2024; Kwon et al., 2023; Yu et al., 2022). One of the key techniques used to improve LLM serving throughput is batching (Yu et al., 2022; Kwon et al., 2023; Patel et al., 2023; Agrawal et al., 2024). Out of the two phases of LLM inference – namely prefill and decode – the decode phase is memory-bound because it processes a single token at-a-time per request. Batching amortizes the cost of fetching model weights from GPU memory and boosts throughput by improving memory bandwidth utilization (Agrawal et al., 2023).
System/library and issues related to PagedAttention
vLLM: Pioneered PagedAttention. Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.85×2.85$\times$2.85 × slower than FlashAttention (9(b)). Further, block size has as much as 3×3$\times$3 × effect on the execution time of the PagedAttention kernel (Figure 5).
FlashAttention: PagedAttention version of the kernel is up to 12% slower than the vanilla version. Initial attempts to add paging support failed unit tests (fa-, 2023).
TensorRT-LLM: Serving throughput dropped by more than 10% in a Python front-end (ten, 2023). Recommends using the C++ frontend. Even with C++, we observe up to 5% higher latency in some cases with PagedAttention.
FlashInfer: Using PagedAttention-based prefill kernel increases time-to-first-token (TTFT) by up to 45% (Table 6).
Table 1. The PagedAttention model requires an application to explicitly manage dynamically allocated physical memory. These examples highlight the complexity, performance and maintenance challenges associated with this approach.
Efficient inference also requires careful allocation of GPU memory. For every request, an LLM maintains an in-memory state known as the KV-cache and re-uses it in every iteration for the lifetime of the request (Yu et al., 2022; Agrawal et al., 2023; Patel et al., 2023). Achieving high memory capacity utilization during inference is challenging for two reasons: 1) per-request KV-cache grows slowly i.e., one token per iteration (few 10s of milliseconds) and 2) the number of generated tokens, and hence the total size of a request’s KV-cache, is typically not known ahead-of-time.
Prior systems like Orca (Yu et al., 2022) and FasterTransformer (fas, 2024b) allocate a contiguous chunk of virtual memory (backed by pre-allocated physical memory) for the KV-cache. The allocated size corresponded to model’s maximum supported sequence length, e.g., 32K. Since models often generate far fewer tokens than the maximum limit, significant GPU memory was wasted due to internal fragmentation. Consequently, these systems exhibit poor throughput as they are unable to support large batch sizes.
System Memory management in user code KV-cache fragmentation KV-cache allocation Portable wrt GPU code
Programming Effort
GPU CPU
vLLM (Kwon et al., 2023) Block-Table(s) Low Fast No High High
FlashInfer (Ye et al., 2024) Compressed Block-Table(s) Low Fast No High High
vAttention None Low Slow Yes None Low
Table 2. Dynamic memory management in various systems. TensorRT-LLM (trt, 2023), LightLLM (lig, 2023) and FlashAttention (Dao, 2023) also adopt vLLM-style PagedAttention model. FlashInfer uses a compressed representation and prefetching to limit the overhead of paging. We introduce various LLM-specific optimizations to hide the impact of high allocation latency in vAttention.
Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (Kwon et al., 2023) to mitigate KV-cache related memory fragmentation. Instead of reserving the maximum sequence length of KV-cache memory ahead-of-time, vLLM allocates small blocks of virtual memory (backed by physical memory) on-demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens. However, dynamically allocated blocks are not guaranteed to be contiguous in virtual memory (the system may have allocated those to other requests). Thus, PagedAttention accepts that KV-cache allocations will be non-contiguous and implements a block-table to stitch together these non-contiguous allocations (§3.2).
Today, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems e.g., in TensorRT-LLM (trt, 2023), HuggingFace TGI (hft, [n. d.]), FlashInfer (Ye et al., 2024), LightLLM (lig, 2023) etc. The most notable aspect of the PagedAttention approach is that it stores KV-cache in non-contiguous virtual memory to be able to allocate physical memory dynamically. While this approach provides an adequate solution for KV-cache fragmentation, we argue that it has several pitfalls (see Table 1 for empirical evidence and real-world experiences):
1. Requires re-writing the attention kernel (GPU code). The elements of a virtually contiguous object can be accessed using index-based lookup which is both simple and efficient [^1] A popular example of this is arrays vs. linked lists.. By storing KV-cache in non-contiguous virtual memory, PagedAttention mandates re-writing GPU code so that the attention kernel can de-reference all the elements of KV-cache. The need to re-write code is a major barrier to using new attention optimizations in production settings.
2. Adds software complexity and redundancy (CPU code). PagedAttention also forces the developers to implement a memory manager inside the serving framework, making it responsible for (de)allocating KV-cache and tracking the location of dynamically allocated KV-cache blocks. This approach essentially translates to re-implementing demand paging – which is an OS functionality – in user code.
3. Introduces performance overhead. PagedAttention can add runtime overhead in the critical path of execution in two ways. First, it requires GPU kernels to execute extra code related to fetching KV-cache from non-contiguous memory blocks. We show that this can slow down attention computation by more than 10% in many cases. Second, the user space memory manager can add CPU overhead, contributing up to another 10% cost (§3.3).
Figure 1. Illustration of memory waste due to internal fragmentation. Orca (top) reserves memory for the KV-cache ahead-of-time. vLLM (bottom) mitigates fragmentation with dynamic memory allocation. Shaded boxes represent memory occupied by the KV tokens whereas the white boxes represent allocated but unused memory.
In this paper, we argue that retaining the virtual memory contiguity of KV-cache is critical for reducing software complexity and redundancy in LLM deployments. Instead of re-implementing paging at the user-level, we contend that existing virtual memory abstractions in the OS can be re-purposed for dynamic KV-cache memory management, resulting in simplified deployments as well as higher performance.
To demonstrate this, we propose vAttention – a system that stores KV-cache in contiguous virtual memory without committing physical memory ahead-of-time. We achieve this by leveraging CUDA support of low-level virtual memory APIs which expose distinct interfaces for allocating virtual and physical memory (§5). vAttention exposes a set of simple APIs using which a serving framework reserves contiguous virtual space for the KV-cache and allocates physical memory on-demand. This approach lends several benefits as listed in Table 2. vAttention also improves portability by enabling a seamless re-use of readily available GPU kernels while eliminating the need to implement a memory manager in a serving system.
Challenges and Optimizations: vAttention solves two key challenges in enabling efficient dynamic memory management without PagedAttention (§6). First, the minimum physical memory allocation granularity supported by CUDA APIs is 2MB. This size can result in significant wasted capacity depending on the model and workload characteristics (Table 8). To address this, we modify the open-source CUDA unified virtual memory driver to add support for finer-grained physical memory allocations of 64KB to 256KB. Second, memory allocation using CUDA APIs incurs high latency because each allocation involves a round-trip to the kernel. To hide the latency of memory allocation from end-users, we introduce several LLM-specific optimizations such as overlapping memory allocation with compute, executing some operations ahead of time and deferring memory reclamation. We show that our optimizations make vAttention an efficient KV-cache memory manager.
Overall, we make the following contributions in this paper:
  • We present vAttention, a system that retains the virtual contiguity of KV-cache while enabling dynamic allocation of physical memory.
  • We show that vAttention is able to seamlessly add dynamic memory management support to unmodified attention kernels of FlashAttention (fag, 2022) and FlashInfer (fig, 2023) while also being performant.
  • We evaluate Yi-6B, Llama-3-8B and Yi-34B on 1-2 A100 GPUs and show that using FlashAttention’s original kernel, vAttention outperforms vLLM by up to 1.97×1.97$\times$1.97 ×, while reducing the time-to-first-token (TTFT) by up to 1.45×1.45$\times$1.45 × over the PagedAttention variant of FlashInfer.
2. Background
2.1. Large Language Models
Given an input sequence, an LLM predicts the probability of an output sequence wherein a sequence is a set of tokens (Kwon et al., 2023). Each inference request begins with a prefill phase that processes all its prompt tokens in parallel. The prefill phase produces the first output token of a request. Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (Agrawal et al., 2023).
LLMs are built atop one of variants of the transformer architecture (Vaswani et al., 2017). A transformer block contains two types of operators: position-wise and sequence-wise. The former category includes feed-forward network, layer normalization, activation, embedding layer, output sampling layer, and residual connections whereas attention is a sequence-level operator. In this paper, we primarily focus on attention since it is the primary consumer of GPU memory in LLM inference.
For the attention operator, the model first computes the query, key and value vectors from a given sequence of tokens $(x_{1},x_{2},....,x_{K})\in\mathbb{R}^{K\times E}$ where E represents the embedding size of the model. For each $x_{i}$, query, key and value vectors are computed as follows:
(1) $q_{i}=W_{q}x_{i},\;\;\;k_{i}=W_{k}x_{i},\;\;\;v_{i}=W_{v}x_{i}$
The resulting $k_{i}$ and $v_{i}$ are appended to the key and value vectors of the prior tokens of the corresponding request, producing two matrices $K,V\in\mathbb{R}^{L^{\prime}\times(H\times D)}$ where $L^{\prime}$ represents the context length of the request seen so far, H is the number of KV heads and D is the dimension of each KV head. Then, attention is computed as follows:
(2) $Attention(q_{i},K,V)=softmax(\frac{q_{i}K^{T}}{scale})V$
The attention score is computed separately for each request in the batch. Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention [^2] An exception to this is sliding window attention that fixes an upper bound on how many of the prior tokens are used in computing attention.. Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations: we refer to this state as KV-cache. A request is executed until the model generates a special end-of-sequence token or reaches the maximum context length for the request.
Structure of the KV-cache and terminology: An LLM consists of multiple layers of the transformer block and each layer maintains its own cache of keys and values. In this paper, we refer to the cache of all transformer blocks collectively as KV-cache while using the term K-cache or V-cache for keys and values, respectively. In deep learning frameworks, the K-cache (or V-cache) at each layer is typically represented as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request. We refer to a kernel implementation that computes attention scores over contiguously stored K and V as a vanilla kernel.
2.2. Fragmentation and PagedAttention
To improve serving throughput, production systems rely on batching which requires careful allocation of GPU memory. This is challenging because the total context length of a request is not known in advance. Serving systems worked around this challenge by pre-reserving KV-cache space assuming that each context is as long as the maximum length supported by the model (e.g.,200K for Yi-6B-200K). vLLM shows that this strategy is prone to severe internal fragmentation. In fact, vLLM showed that prior reservation wastes memory even if the context lengths are known in advance. This is because the per-request KV-cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
Inspired by the OS-based virtual memory systems, vLLM proposed PagedAttention to mitigate fragmentation by dynamically allocating memory for the KV-cache. PagedAttention splits KV-cache into fixed-sized blocks and allocates memory for one block at a time. This way, vLLM allocates only as much memory as a request needs, and only when required – not ahead-of-time. Figure 1 shows an example of how reservation-based systems such as Orca (Yu et al., 2022) can waste significant memory due to fragmentation and how vLLM avoids it with dynamic memory allocation.
Figure 2. PagedAttention involves two layers of address translation. First, the framework tracks the virtual memory addresses of KV-cache blocks via Block-Tables. Second, the GPU performs virtual-to-physical address translation via OS-managed Page-Tables.
3. Issues with the PagedAttention Model
Despite being inspired by demand paging, PagedAttention adopts an approach that is different from conventional demand paging: it requires an application’s code to be modified to adapt to dynamically allocated physical memory whereas conventional demand paging is transparent to applications. This section elaborates on some of the issues that arise with such an approach.
3.1. Requires re-writing the attention kernel
PagedAttention necessitates re-writing the attention kernel. This is because conventional implementations of the attention operator assume that the two input tensors K and V (Equation 2) are stored in contiguous memory. By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV-cache blocks. Writing correct and performant GPU kernels can be challenging for most programmers (fa-, 2023).
Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (Dao et al., 2022; Dao, 2023; fla, 2023; Hong et al., 2024; Ye et al., 2024; Zhang et al., 2023; Bikshandi and Shah, 2023; Ainslie et al., 2023; Shazeer, 2019; Child et al., 2019; Beltagy et al., 2020), and this trend is likely to continue. In the PagedAttention model, keeping up with new research requires continued efforts in porting new optimizations to a PagedAttention-aware implementation. Production systems can therefore easily fall behind research, potentially losing performance and competitive advantage. To provide an example, 9(b) shows that the paged kernel of vLLM is already up to 2.85×2.85$\times$2.85 × slower than the FlashAttention counterpart for grouped-query attention (Ainslie et al., 2023).
Figure 3. Overhead of PagedAttention in FlashAttention’s decode kernel (model: Yi-6B). Legend shows the batch size.
3.2. Adds redundancy in the serving framework
PagedAttention makes an LLM serving system responsible for managing the mappings between KV-cache and dynamically allocated memory blocks. For example, consider a request that allocates four KV-cache blocks over time (left half of Figure 2). These blocks are usually non-contiguous in virtual memory. During the computation of Equation 2, PagedAttention kernel needs to access all the elements of the four KV-cache blocks. To facilitate this, the serving system needs to track the virtual memory addresses of KV-cache blocks and pass them to the attention kernel at runtime. This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half Figure 2).
3.3. Performance Overhead
PagedAttention also leads to potential performance issues both on GPU and CPU. We investigate them separately.
3.3.1. Runtime overhead on the GPU
PagedAttention slows down attention computation by adding extra code in the critical path. For example, vLLM acknowledges that their PagedAttention-based implementation was $20-26\%$ slower than the original FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (Kwon et al., 2023). Figure 3 shows that the paged decode kernel in FlashAttention is also slower than the vanilla kernel. Our further analysis reveals that the number of instructions executed in PagedAttention is up to $13\%$ higher than the vanilla kernel. We also find that the overhead of paging reduces at high batch sizes or long context lengths. This is because computing attention for decodes is memory bound and when the KV-cache size is large, memory stalls hide the instruction overhead.
Figure 4. Overhead of PagedAttention in FlashInfer’s prefill kernel (model: Yi-6B). Legend shows the batch size.
However, hiding instruction overhead is more challenging in a prefill attention kernel which is compute-bound with $N^{2}$ complexity. For example, Figure 4 shows that the paged version of FlashInfer (Ye et al., 2024) prefill kernel is up to 14% slower than the vanilla kernel. The paged kernel is also missing some well-known optimizations (fip, 2024).
To highlight another example of difficulty involved in writing an efficient attention kernel, Figure 5 shows that the performance of vLLM’s paged decode kernel is significantly worse with larger block sizes of 64 and 128. Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in L1 cache.
3.3.2. Runtime overhead on the CPU
Implementing an additional memory manager can add performance issues in the CPU runtime of the serving system. We refer to a few real-world examples and our own observations on vLLM to corroborate this argument.
To enable PagedAttention, a serving system needs to supply Block-Tables to the attention kernel. In vLLM, the latency of preparing a Block-Table depends on batch composition and grows proportional to max_num_blocks $\times$ batch_size where max_num_blocks refers to the number of KV-cache blocks in the longest request of the batch. This is because vLLM manages a Block-Table as a 2D tensor and aligns the number of KV-cache blocks in each request by padding unoccupied slots with zeros. If a batch contains a few long and many short requests, such padding results in a significant overhead. In our earlier experiments, we observed that Block-Table preparation in vLLM was contributing 30% latency in decode iterations. While a recent fix (vll, 2024a) has mitigated some of this overhead, we find that it can still be as high as $10\%$. High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11%, from 412 tokens/sec to 365 tokens/sec (trt, 2023). This issue was attributed to the Python runtime of TensorRT-LLM and moving to a C++ runtime can mitigate the CPU overhead.
Overall, this section shows that the PagedAttention model adds a significant programming burden while also being inefficient. vAttention introduces a more systematic approach to dynamic KV-cache memory management by leveraging the existing system support for demand paging. However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads in terms of memory management.
Figure 5. Latency of vLLM’s paged decode kernel is sensitive to block size (model: Yi-6B). Legend shows the block size.
(a) Output tokens per second.
(b) Decode iteration latency.
(c) Memory allocation bandwidth.
Figure 6. LLM inference throughput plateaus at large batch sizes (a) beyond which latency suffers (b) without improving throughput. This also means that the rate at which physical memory needs to be allocated also saturates (c).
4. Insights into LLM Serving Systems
To highlight the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single NVIDIA A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism. We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure latency, throughput and memory requirement of the decode phase (see §6 for our discussion and optimizations for the prefill phase).
Observation-1: On a per-iteration basis, KV-cache memory requirement is known in advance. This is due to auto-regressive decoding that generates one token at a time, per-request. Therefore, with every iteration, the KV-cache memory footprint of a request grows uniformly by one token until the request completes.
Observation-2: KV-cache does not require high memory allocation bandwidth. The memory footprint of a single token across all layers is typically few 10s-100s of kilobytes of memory. For example, the per-token memory footprint of Yi-6B, Llama-3-8B and Yi-34B is 64KB, 128KB and 240KB, respectively. Further, each iteration runs for 10s-100s of milliseconds (6(b)) implying that a request requires at most a few megabytes of memory per second. While batching improves system throughput, the number of tokens generated per second plateaus beyond a certain batch size (6(a)). This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 128 for Yi-34B). For all the models we studied, we observe that the highest memory allocation rate is at most 600MB per second (6(c)).
In vAttention, we leverage these observations to implement an efficient dynamic memory management system for KV-cache. In the next section, we begin with a high-level design description of vAttention (§5.1), then discuss how vAttention is used for serving LLMs (§5.3) and finally describe our optimizations (§6).
5. vAttention: System Design
Our goal is to improve efficiency and portability by adding dynamic memory allocation support to existing kernels. To achieve this goal, vAttention leverages system support for dynamic memory allocation instead of implementing paging in user space.
5.1. Design Overview
vAttention builds on the ability to allocate virtual memory and physical memory separately. Specifically, we allocate a large contiguous buffer for the KV-cache in virtual memory ahead-of-time (similar to reservation-based allocators) while deferring the allocation of physical memory to runtime i.e., allocate physical memory only when required (similar to PagedAttention). This way, vAttention preserves virtual contiguity of KV-cache without wasting physical memory. This approach is feasible because memory capacity and fragmentation are limiting factors only for physical memory whereas virtual memory is abundant e.g., modern 64-bits systems provide a 128TB user-managed virtual address space to each process33364-bits systems typically utilize 48 bits for virtual addresses, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space..
Figure 7. Dynamic memory management in vAttention for a single K-cache (or V-cache) tensor. (a) shows a virtual tensor for a batch of two requests with no physical memory allocation yet. (b) R1 is allocated one physical page. (c) R1 is allocated two pages and R2 is allocated one page. (d) R1 has completed but vAttention does not reclaim its memory (deferred reclamation). (e) when R3 arrives, vAttention assigns R1’s tensor to it which is already backed by physical memory.
5.1.1. Pre-reserving virtual memory
Since virtual memory is abundant, we pre-allocate enough virtual memory space that is large enough to hold the KV-cache of the maximum batch size (configurable) that needs to be supported.
Number of virtual memory buffers: Each layer in an LLM maintains its own K and V tensors: we refer to them individually as K-cache and V-cache. We allocate separate virtual memory buffers for K-cache and V-cache. For a single GPU job, this requires pre-reserving $2\times N$ buffers where $N$ is the number of layers in the model. In a multi-GPU job, each worker reserves $2\times N^{\prime}$ buffers where $N^{\prime}$ is the number of layers managed by that worker ($N^{\prime}=N$ with tensor-parallelism whereas $N^{\prime}<N$ with pipeline-parallelism).
Size of a virtual memory buffer: The maximum size of a buffer is $BS=B\times L\times S$ where B is the maximum batch size, L is the maximum context length supported by the model, and $S$ is the size of a single token’s per-layer K-cache (or V-cache) on a worker. Further, $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16). Note that $S$ is constant for a given model configuration.
Consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2). In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$. For this model, the maximum size of K-cache (or V-cache) per-worker per-layer is $S=200MB$ $(200K*4*128*2)$. Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$). Therefore, the total virtual memory requirement for 60 layers of Yi-34B is 120 buffers of 100GB each (12TB total). Note that the amount of virtual address space available grows with the number of GPUs e.g., with two TP workers, the amount of virtual address space available is 256TB. Therefore, virtual memory allocations can be satisfied easily.
5.1.2. On-demand physical memory allocation
vAttention preferentially allocates physical memory one page at a time and only when a request has used all of its previously allocated physical memory pages. To show how it works, we refer to a simple example in Figure 7. The example shows how vAttention manages the K-cache (or V-cache) at one layer of the model assuming maximum batch size of two. Rest of the K-cache and V-cache buffers are managed similarly at all layers.
5.2. Leveraging Low-level CUDA Support
The standard GPU memory allocation interface cudaMalloc does not support demand paging i.e., it allocates virtual memory and physical memory at the same time. However, recent CUDA versions provide programmers a fine-grained control over virtual and physical memory (cud, 2024; Guo et al., 2024). We leverage these low-level APIs in vAttention.
5.2.1. CUDA virtual memory APIs
Table 3 provides a high-level overview of CUDA APIs that allow separating the allocation of virtual memory from physical memory (see the leftmost column). The allocation granularity depends on the page size used by the GPU and the size of virtual memory buffer or a physical memory handle must be a multiple of the allocation granularity. Different sub-regions of a virtual memory buffer can be backed by physical memory independently of other sub-regions in that buffer (see Figure 7c for an example). For simplicity, we refer to the granularity at which physical memory is allocated as page size.
Latency (microseconds)
CUDA APIs vAttention APIs Description 64KB 128KB 256KB 2MB
cuMemAddressReserve * vMemReserve * Allocate a buffer in virtual memory 18 17 16 2
cuMemCreate * vMemCreate * Allocate a handle in physical memory 1.7 2 2.1 29
cuMemMap vMemMap Map a physical handle to a virtual buffer 8 8.5 9 2
cuMemSetAccess - Enable access to a virtual buffer - - - 38
[…]
cuMemRelease * vMemRelease * Free physical pages of a handle 2 3 4 23
cuMemAddressFree * vMemFree * Free a virtual memory buffer 35 35 35 1
Table 3. Low-level APIs for virtual memory management and their latency with different allocation sizes. * represents APIs that we use once while instantiating or terminating the serving framework. Rest of the APIs are used for (un)mapping physical memory pages at runtime. CUDA APIs (prefixed with cu) support only 2MB allocation sizes, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
5.2.2. Extending PyTorch caching allocator:
KV-cache is a collection of tensors. In current deep learning frameworks such as PyTorch, a tensor allocated via APIs such as torch.empty comes with pre-allocated physical memory. This is because the PyTorch caching allocator uses the cudaMalloc interface to allocate GPU memory (both virtual and physical). Relying on the low-level API support from CUDA, we extend the PyTorch caching allocator to allow an application to reserve a virtual memory buffer for a tensor without committing physical memory ahead-of-time. We refer to tensors allocated via these APIs as virtual tensors.
5.2.3. Request-level KV-cache indexing:
Note that each virtual tensor represents the K-cache (or V-cache) of a layer for the maximum batch size B. In these tensors, different requests occupy different non-overlapping sub-regions (say sub-tensors). We locate the sub-tensor of a request with a unique integer identifier reqId that lies in the range of 00 to $B-1$ (note that at most $B$ requests run simultaneously). The K-cache (or V-cache) offset of a request’s sub-tensor in the virtual tensor of the entire batch is $\texttt{reqId}\times S$ where $S$ is the maximum K-cache (or V-cache) size of a request on a worker. The request identifier reqId is allocated by vAttention.
5.3. Serving LLMs with vAttention
We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers. Our library exposes a set of simple APIs listed in Table 4 to the serving framework.
5.3.1. Initial setup:
When the serving framework starts, each model worker loads the vAttention library and configures it with model parameters $N^{\prime},H,D,P$, $B$ and a preferred page size via the init API. Internally, vAttention reserves $2\times N^{\prime}$ virtual tensors (using our modified PyTorch caching allocator) for the KV-cache at each worker. These virtual tensors are reserved for the lifetime of the serving application. In addition, vAttention also pre-allocates physical memory pages during initialization. However, these pages are not mapped into the KV-cache yet.
APIs Description
init Initializes vAttention with model parameters.
arguments: $N^{\prime},B,L,H^{\prime},P$, page_size.
return value: a list of KV-cache tensors.
alloc_reqid Allocates an unused reqId and marks it active
arguments: None
return value: an integer reqId
free_reqid Frees a reqId and marks it inactive
arguments: an integer reqId
return value: None
step Ensures physical memory pages are mapped
arguments: an array of size B containing sequence lengths of each reqId
return value: 0 (success), -1 (failure).
Table 4. APIs exposed to a serving framework for dynamic KV-cache memory management with vAttention.
5.3.2. Scheduling a new request:
When a new request is scheduled for the first time, the serving framework obtains a new reqId from vAttention via alloc_reqid. All subsequent memory management operations of the request are tagged with this reqId.
5.3.3. Model execution:
Before scheduling a batch for execution, the framework needs to ensure that the KV-cache sub-tensors of each active request are backed by physical memory. For this purpose, before dispatching the first kernel of an iteration to the GPU, the framework invokes the step API, specifying the current context length of each request (context length is set to 0 for each inactive reqId). Internally, vAttention ensures that enough physical pages are mapped for each active reqId before returning execution back to the framework. If vAttention cannot satisfy the memory demand, it returns with a failure in response to which a serving framework can preempt one or more requests to allow forward progress (this is similar to vLLM’s default behavior). We leave more sophisticated policies such as swapping out KV-cache to CPU memory as future work.
Depending on whether a request is in the prefill phase or decode phase, different number of physical memory pages may need to be mapped for a given iteration. The prefill phase processes the input tokens of given prompt in parallel and populates one slot in the K-cache (and V-cache) of the request at each layer of the model. Therefore, the number of pages needed to be mapped depends on the number of prompt tokens being scheduled. If the total K-cache size of all prompt tokens at one layer of the model is $s$ and page size is $t$, then each worker needs to ensure that at least $(s+t-1)/t$ physical memory pages are mapped in each of the $2\times N^{\prime}$ KV-cache sub-tensors of the given reqId.
For a request in the decode phase, the number of new pages required is at most one per request. This is because each iteration produces only one output token for a request. vAttention internally tracks the number of pages mapped for each request and maps a new page only when the last page allocated to that request is fully utilized.
5.3.4. Request completion.
A request terminates when a user specified context length or the maximum context length supported by the model is reached, or when the model produces a special end-of-sequence token. The framework notifies vAttention of a request’s completion with free_reqid. Internally, vAttention may unmap the pages of a completed request or defer them to be freed later.
6. vAttention: Optimizations
There are two primary challenges in using CUDA’s virtual memory support for serving LLMs. First, cuMemCreate currently allocates a minimum of 2MB physical memory page. Large pages can waste physical memory due to internal fragmentation. Second, invoking CUDA APIs incurs high latency. This section details a set of simple-yet-effective optimizations that we introduce to overcome these limitations.
6.1. Mitigating internal fragmentation
We mitigate internal fragmentation by reducing the granularity of physical memory allocation. NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB. Therefore, in principal, physical memory can be allocated in any multiple of 4KB sizes. The simplest way to achieve this would be to extend the existing CUDA virtual memory APIs (listed in Table 3) to also support allocating smaller pages (similar to how mmap in Linux supports multiple page sizes). Unfortunately, the CUDA APIs are implemented in the closed-source NVIDIA drivers which makes it impossible for us to modify their implementation.
Fortunately, some part of NVIDIA drivers (particularly related to unified memory management) is open-source. Therefore, we implement a new set of APIs in the open-source NVIDIA drivers to mimic the same functionality that existing CUDA APIs provide but with support for multiple page sizes. The second column in Table 3 shows our new APIs: most of our APIs have a one-to-one relationship with existing CUDA APIs except for vMemMap that combines the functionality of cuMemMap and cuMemSetAccess, and vMemRelease that combines the functionality of cuMemUnmap and cuMemRelease for simplicity. In contrast to CUDA APIs, our APIs can allocate memory in 64KB, 128KB and 256KB page sizes. A serving framework can configure a desired page size in vAttention while initializing it: we recommend using 256KB pages by default. The last set of columns in Table 3 shows the latency of each API with different page sizes.
6.2. Hiding memory allocation latency
The serving framework invokes the step API in every iteration. The latency of step depends on how many new pages need to be mapped in the virtual tensors of KV-cache. Consider, for example, that the KV-cache of one request needs to be extended for Yi-34B which has 60 layers. This requires 120 calls to vMemMap each of which takes about 9 microseconds. Therefore, growing the KV-cache of one request by one page will add about 1 millisecond latency to the corresponding iteration and would grow proportional to amount of physical memory that needs to be mapped. We propose the following optimizations to hide the latency of allocation:
6.2.1. Overlapping memory allocation with compute
We leverage the predictability of memory demand to overlap memory allocation with computation. In particular, note that each iteration produces a single output token for every decode request. Therefore, memory demand for a decode iteration is known ahead-of-time. Further, in the decode phase, a request requires at most one new page. vAttention keeps track of the current context length and how many physical memory pages are already mapped for each request. Using this information, it determines when a request would need a new page and uses a background thread to allocate a new page when the preceding iteration is executing. For example, consider that a request R1 would require a new page in iteration i. When the serving framework invokes step API in iteration i-1, vAttention launches a background thread that maps physical memory pages for iteration i. Since iteration latency is typically in the range of 10s-100s of milliseconds, the background thread has enough time to prepare physical memory mappings for an iteration before it starts executing. This way, vAttention hides the latency of CUDA APIs by mapping physical pages in the KV-cache tensors out of the critical path. Note that in every iteration, step API still needs to ensure that physical pages required for the current iteration are actually mapped. If not, required pages are mapped synchronously.
6.2.2. Deferred reclamation + eager allocation
We observe that allocating physical memory for a new request can be avoided in many cases. Consider that a request R1 completed in iteration i and a new request R2 joins the running batch in iteration i+1. To avoid allocating new pages to R2 from scratch, vAttention simply defers the reclamation of R1’s pages and assigns R1’s reqId to R2. This way, R2 uses the same tensors for its KV-cache that R1 was using which are already backed by physical pages. Therefore, new pages for R2 are required only if its context length is bigger than that of R1.
We further optimize memory allocation by proactively mapping physical pages before they are needed. We do so by using one of the inactive reqId’s KV-cache. When a new request arrives, we can allocate this reqId without mapping any physical pages. We then select a new reqId that would be allocated next and map physical pages for it. In most cases, these eager optimizations obviate the need to allocate new physical pages even for the prefill phase of new requests. Finally, we trigger memory reclamation only when the number of physical memory pages cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory). We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
Model Hardware ##\## Q Heads ##\## KV Heads ##\## Layers
Yi-6B 1 A100 32 4 32
Llama-3-8B 2 A100s 32 8 32
Yi-34B 2 A100s 56 8 60
Table 5. Models and hardware used for evaluation.
7. Evaluation
Our evaluation seeks to answer the following questions:
  • How does vAttention perform for prefill and decode phases in LLM inference? What are the portability and performance advantages of vAttention.
  • How efficiently can vAttention allocate GPU memory for LLM serving workloads, and how effectively can it deal with KV-cache fragmentation?
Models and Hardware: We evaluate three models Yi-6B, Llama-3-8B and Yi-34B, using a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 5). Each GPU has 80GB physical memory. We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B. All three models use GQA which is the most commonly used attention mechanism in recent LLMs.
Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different. Attention kernels used for these two phases are also different and hence we evaluate them separately. The prefill phase requires one time memory allocation potentially spanning multiple pages. In comparison, the decode phase requires incremental memory allocation over the lifetime of a request. We measure the throughput of these phases in terms of tokens processed (or generated) per second.
Figure 8. Prefill tokens processed per second with chunking (chunk size=2048). Using vanilla prefill kernels, vAttention outperforms the paged counterpart of both FlashAttention and FlashInfer. Throughput for longer context lengths drops sharply due to quadratic complexity of prefill attention.
7.1. Portability and Performance for Prefills
To evaluate the prefill phase, we focus on the attention kernels provided by FlashAttention v2.5.6 (Dao, 2023; fag, 2022) and FlashInfer v0.0.3 (fig, 2023; Ye et al., 2024). We do not include vLLM in these experiments because it does not have a prefill kernel of its own but instead uses FlashAttention kernel. We also could not evaluate Yi-34B because FlashInfer kernels do not support Yi-34B’s KV group size of 7 (fip, 2024).
FlashInfer is a library that recently introduced a set of attention kernels optimized for different scenarios e.g., for chunked-prefills – an optimization proposed in Sarathi (Agrawal et al., 2023) and later adopted in various systems (Agrawal et al., 2024; Holmes et al., 2024; Hu et al., 2024). Sarathi splits the input tokens of a prompt into multiple smaller chunks and schedules one chunk at a time, enabling a serving system to add new requests in a batch without pausing ongoing decodes. This helps improve throughput without increasing latency (Agrawal et al., 2024). Both FlashAttention and FlashInfer provide kernels to compute the attention scores of chunked-prefills with and without PagedAttention. We integrate them into vLLM and using chunk size of 2048 tokens, measure time-to-first-token (TTFT) for the following configurations:
FA_Paged: Uses flash_attn_with_kv_cache kernel API of FlashAttention.
FI_Paged: Uses FlashInfer’s PagedAttention kernel, representing state-of-the-art PagedAttention-based kernel for the prefill phase.
FA_vAttention: Uses FlashAttention’s vanilla prefill kernel via the flash_attn_func API.
FI_vAttention: Uses FlashInfer’s vanilla prefill kernel via the single_prefill_with_kv_cache API.
Model Context Length FlashAttention FlashInfer
Paged vAttention Paged vAttention
Yi-6B 16K 2.44 1.38 1.51 1.34
32K 8.07 3.50 3.97 3.4
64K 30.46 9.98 11.85 9.66
128K 119 32.10 39.17 30.9
192K 266 67.81 81.87 65.03
[…]
32K 5.19 2.59 3.06 2.37
Llama-3 64K 18.65 6.84 8.94 6.50
-8B 128K 71.16 20.78 28.98 20.29
192K 157.33 42.31 59.91 41.35
Table 6. Time-to-first-token (TTFT) for a single prompt of varying context lengths.
(a) Decode throughput (measured as the number of output tokens generated per second) with varying batch size.
(b) Mean GPU time spent in computing attention with varying batch sizes.
Figure 9. Decode throughput with varying batch sizes using context length 16K for each request (FA: FlashAttention, bs: block size). We evaluate vLLM and FlashAttention with two different block sizes: 16 and 128. vLLM performs best with block size 16 because its attention kernel is more efficient with smaller block sizes. FlashAttention’s GPU kernel is up to 2.85×2.85$\times$2.85 × faster than the best version of vLLM’s kernel (Yi-6B, 16*16K). However, smaller blocks add CPU overhead e.g., FlashAttention with block size 16 is worse than with block size 128. vAttention provides similar gains that the best version of FlashAttention provides over vLLM, but without user-level physical memory management and without a PagedAttention kernel.
Both vAttention configurations therefore use kernels that support chunked-prefills over a virtually contiguous KV-cache. We add dynamic memory allocation support to them without having to modify their code.
Figure 8 shows the prefill throughput of the four configurations for Yi-6B and Llama-3-8B. In all cases, vAttention delivers consistently higher throughput, outperforming FA_Paged by 1.603.92×1.60-3.92$\times$1.60 - 3.92 × and FI_Paged by 1.031.45×1.03-1.45$\times$1.03 - 1.45 ×. For the same experiment, Table 6 shows the TTFT with different context lengths. Since TTFT directly depends on prefill throughput, compared to using vanilla kernels with vAttention, FA_Paged and FI_Paged increase TTFT by up to 3.92×3.92$\times$3.92 × (Yi-6B, context length 192K) and 1.45×1.45$\times$1.45 × (Llama-3-8B, context length 192K), respectively.
The source of vAttention’s performance gain is twofold in these scenarios. First, the vanilla kernel is faster than the paged kernel in both FlashAttention and FlashInfer. While FlashAttention’s paged kernel is not optimized for prefills (it is optimized for decodes), FlashInfer’s paged kernel is specifically designed to support chunked-prefills. However, the paged kernel is slower than the vanilla kernel as discussed in §3.3. This example illustrates the complexities of transferring performance-critical optimizations between different implementations – even when the implementations are written by the same team. The second source of improvement is vAttention’s less CPU overhead. For example, appending a new K or V tensor to the KV-cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time. Further, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration. vAttention avoids such overheads because it maintains KV-cache’s virtual contiguity and therefore does not require a Block-Table.
7.2. Portability and Performance for Decodes
To evaluate decode performance, we focus on long-context scenarios (16K) because the latency of attention kernel becomes significant only for long contexts [^4] For short contexts, the computation time of the feed-forward-network dominates inference latency (Agrawal et al., 2024). We evaluate the following configurations:
vLLM: We use vLLM v0.2.7 as the primary baseline. vLLM pioneered PagedAttention and uses a custom paged kernel for decodes, derived from FasterTransformer (ftk, [n. d.]).
FA_Paged: For the second baseline, we integrate the FlashAttention kernel into vLLM’s serving stack. This represents a state-of-the-art PagedAttention kernel that includes optimizations such as sequence parallelism and in-place copy of new key and value vectors into the KV-cache. We evaluate the paged kernels of vLLM and FlashAttention with two different block sizes – 16 and 128 – to capture the effect of block size on performance.
FA_vAttention: For vAttention, we integrated the vanilla kernel of FlashAttention into vLLM’s serving stack. The kernel works with a virtually contiguous KV-cache to which we dynamically allocate physical memory using 2MB pages.
9(a) shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch sizes wherein the initial context length of each request is 16K tokens and we generate $256$ tokens for each request. We compute decode throughput based on the mean latency of 256 decode iterations. We summarize the key takeaways below.
First, vAttention outperforms vLLM (both block sizes) and FA_Paged (block size 16), while roughly matching the best configuration of FA_Paged (block size 128). The maximum improvement over vLLM is 1.97×1.97$\times$1.97 × for Yi-6B, 1.3×1.3$\times$1.3 × for Llama-3-8B and 1.6×1.6$\times$1.6 × for Yi-34B. The relative gains over vLLM also increase as the batch size grows. For example, the gain increases from about $\times1$ to 1.97×1.97$\times$1.97 × as batch size increases from 1 to 8 for Yi-6B. This is because the latency of attention computation grows proportional to the total number of tokens in the batch (see 9(b)) whereas the cost of linear operators remains roughly the same (Agrawal et al., 2023; Patel et al., 2023; Agrawal et al., 2024). Therefore, the contribution of attention kernel in the overall latency – and subsequently gain with a more efficient kernel – increases with the batch size. While FA_Paged (block size 128) provides similar gains as vAttention, note that FA_Paged requires a new implementation of the GPU kernel whereas vAttention simply leverages the vanilla kernel of FlashAttention.
Second, 9(b) confirms that performance difference between vLLM and FA_Paged/vAttention is indeed due to the attention kernels. In the worst case, the latency of vLLM’s best PagedAttention kernel (block size 16) is up to 2.85×2.85$\times$2.85 × higher for Yi-6B, up to 1.45×1.45$\times$1.45 × for Llama-3-8B and up to 2.62×2.62$\times$2.62 × for Yi-34B than the FlashAttention kernel.
Finally, throughput can be sensitive to block size even when memory capacity is not a constraint. For example, as discussed in §3.3, vLLM’s attention kernel has a significantly higher latency with block size 128 than with block size 16 (also see 9(b)). In the worst case, block size 128 degrades vLLM’s throughput by 36%. While block size has a smaller impact on FlashAttention, using a small block size can still hurt throughput due to CPU overheads, particularly due to the overhead of creating Block-Tables for every iteration (§3.3). For example, FlashAttention with block size 128 delivers 7% higher throughput than block size 16 for Llama-3-8B (531 vs 494 tokens per second with batch size 32).
Figure 10. Latency of decode iterations with and without overlapping memory allocation with compute (batch size=4,context length=32K). Spikes show the latency impact of synchronous memory allocation.
Config. 64KB 128KB 256KB 2MB
TP-1 7.59 14.56 27.04 35.17
TP-2 15.18 29.12 54.08 70.34
Table 7. Physical memory allocation bandwidth (GB per second) for vAttention with different page sizes.
7.3. Efficacy of Physical Memory Allocation
The PyTorch caching allocator allocates memory objects (e.g., tensors) without requiring a round-trip to the kernel. In contrast, vAttention needs to invoke CUDA’s kernel driver while mapping a new physical page in a request’s KV-cache. In this section, we show that with our optimizations, vAttention can effectively meet the requirements of both the prefill and decodes phases in an LLM serving system.
Figure 11. Time to compute the prefill phase of a single prompt of 16K tokens. vAttention is as good as vLLM since it allocates memory out of the critical path. For illustration purpose, results with different page sizes show the overhead of synchronously allocating physical memory before a request’s prefill phase begins.
Model # Tokens in a physical memory block Max memory waste per request
64KB 128KB 256KB 2MB 64KB 128KB 256KB 2MB
Yi-6B (TP-1) 64 128 256 2048 4MB 8MB 16MB 128MB
Yi-6B (TP-2) 128 256 512 4096 8MB 16MB 32MB 256MB
Llama-3-8B (TP-1) 32 64 128 1024 4MB 8MB 16MB 128MB
Llama-3-8B (TP-2) 64 128 256 2048 8MB 16MB 32MB 256MB
Yi-34B (TP-1) 32 64 128 1024 7.5MB 15MB 30MB 240MB
Yi-34B (TP-2) 64 128 256 2048 15MB 30MB 60MB 480MB
Table 8. Block size (number of tokens in a newly allocated physical memory page) as a function of the page size and TP configuration. Columns on the right show how much memory can be wasted per-request in the worst-case.
Table 7 shows that even with our smallest page size of 64KB, vAttention can allocate as much as 7.6GB per second per GPU. This is more than an order of magnitude higher than the maximum memory allocation rate of 600MB per second of decodes (Figure 6). Larger page sizes and higher TP dimensions increase the memory allocation rate proportionally. This shows that vAttention is more than capable of satisfying the memory allocation bandwidth of decodes.
Further, Figure 10 shows that our optimization of overlapping memory allocation with model execution also hides the latency impact of invoking CUDA APIs. This example shows the latency of consecutive decode iterations for Llama-3-8B running with TP-1 and batch size 4. Without overlapping memory allocation with compute, new pages for the KV-cache are allocated synchronously which increase the latency of some iterations from 25 milliseconds to 41 milliseconds ($\approx$ 4 millisecond latency due to memory allocation to a single request). Note that these latency spikes occur after every 1024 iterations because we used 2MB page size for these experiments and each 2MB page contains 1024 tokens for this model configuration. When memory allocation is overlapped with the model execution of previous decode iteration, the latency effect is completely hidden.
Finally, since a prefill may require more than one page for the KV-cache, we also investigate how different memory allocation schemes impact the latency of a prefill. Figure 11 shows that allocating physical memory synchronously on-demand (when our background thread, deferred reclamation and eager allocation optimizations are all disabled) can add as much as 15% overhead with 64KB page size. Larger page sizes amortize the cost of allocation and reduce the overhead to as low as 3% (with 256KB and 2MB page sizes). vAttention further reduces the cost of allocation with deferred reclamation and eager allocation while overlapping memory allocation with compute. In most cases, these optimizations ensure that a newly arrived request can simply re-use the physical memory pages that were allocated to a previous request. Therefore, vAttention incurs negligible overhead, performing as good as vLLM for prefills.
7.4. Analysis of Memory Fragmentation
Table 8 shows the block size (defined as the minimum number of tokens in a page), and how much physical memory can be (theoretically) wasted due to over allocation in the worst-case. The worst-case occurs when a new page is allocated but remains completely unused. Further, we show each model under two TP configurations – TP-1 and TP-2 – to highlight the effect of TP dimension on block size.
vAttention allocates physical memory equivalent to the page size on each TP worker whereas the per-token physical memory requirement of a worker goes down as TP dimension increases (because KV heads get split across TP workers). Therefore, block size increases proportionally with the TP dimension. Table 8 shows that this results in the smallest block sizes of 32 (Yi-34B TP-1) to 128 (Yi-6B TP-2). In terms of the amount of physical memory, 64KB page size results in a maximum theoretical waste of only 4-15MB per request which increases to 16-60MB for 256KB page size. Overall, the important point to note is that by controlling the granularity of physical memory allocation, vAttention makes memory fragmentation insignificant. Recall that serving throughput saturates at about 200 batch size for all of our models (Figure 6). Hence, even at such large batch sizes, the maximum theoretical memory waste is at most a few GBs. Therefore, similar to vLLM, vAttention is highly effective in reducing fragmentation and allows serving using large batch sizes. However, if required, page size can be reduced further to as low as 4KB which is the minimum page size supported in almost all architectures today, including NVIDIA GPUs (B et al., 2023).
Figure 12. Illustration of code changes required to replace the prefill attention kernel of FlashAttention with that of FlashInfer. vAttention supports dynamic memory allocation transparently, enabling easy switching between kernels.
Implementation Effort
The primary advantage of vAttention is portability: it enables one to seamlessly integrate new attention kernels without having to having to write a paged version of it or change the serving framework. For example, switching between the prefill or decode kernels of FlashAttention and FlashInfer requires only a few lines of code changes as shown in Figure 12. In contrast, in PagedAttention, the developers first need to write a paged attention kernel and then make significant changes in the serving framework. For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes spread over 15 files (vll, 2024b, d, c). Implementing the initial paging support in FlashAttention GPU kernel also required about 280 lines of code changes (fap, 2024) and additional efforts to enable support for smaller block sizes (fas, 2024a). Given the rapid pace of innovation in LLMs, we believe it is important to reduce programming burden: production systems should be able to leverage new optimizations of the attention operator without re-writing code – similar to how optimized implementations of GEMMs are leveraged by deep learning frameworks without programmer intervention.
8. Related Work
In a recent work, GMLake (Guo et al., 2024) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size. In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors for large object allocations. In contrast, vAttention is focused on avoiding fragmentation for LLM inference. Different from training, LLM inference is latency sensitive and requires smaller granularity allocations. We proposed various LLM inference specific optimizations to meet these requirements.
Optimizing LLM inference is an active area of research. Various scheduling systems have been proposed to improve different aspects of LLM serving. For example, Orca (Yu et al., 2022) and vLLM (Kwon et al., 2023) are aimed at improving serving throughput with efficient batching. Sarathi (Agrawal et al., 2023) and SplitFuse (Holmes et al., 2024) split a long prefill into multiple smaller chunks and combine decode tokens with each chunk to improve GPU compute utilization. Based on similar techniques, Sarathi-Serve (Agrawal et al., 2024) proposes stall-free batching to minimize the impact of long-running prefill iterations on decode latency. Splitwise (Patel et al., 2023), DistServe (Zhong et al., 2024) and TetriInfer (Hu et al., 2024) disaggregate the prefill and decode phases, executing them on different replicas so as to avoid interference between the prefill and decode requests. For offline inference on resource-constrained devices, FlexGen (Sheng et al., 2023) proposed a scheduling and offloading strategy to improve throughput. FastServe (Wu et al., 2023) minimizes job completion times in LLM inference using preemptive scheduling.
For all the above systems to work effectively, efficient use of GPU physical memory is essential. Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trt, 2023), LightLLM (lig, 2023) and kernel implementations e.g., in FlashAttention (fag, 2022) and FlashInfer (fig, 2023). In contrast, vAttention offers an alternate approach for dynamic KV-cache memory management. We show that using system support for demand paging can easily add dynamic memory management support to existing kernel implementations.
9. Conclusion
PagedAttention has emerged as the de facto standard for dynamic memory allocation in LLM inference. PagedAttention eliminates the need to reserve GPU memory ahead-of-time and therefore boosts serving throughput by fitting a larger batch size in GPU memory. While PagedAttention effectively tackles memory fragmentation, we argue that its approach of storing KV-cache in non-contiguous virtual memory introduces software complexity as well as portability and efficiency challenges. Instead, we show that using low-level system support for demand paging can avoid the pitfalls of PagedAttention. Our proposed system vAttention adds support for dynamic physical memory allocation to existing attention kernels, eliminating the need to re-write GPU code or write a memory manager in the serving framework. We show that vAttention reduces software complexity while improving portability and performance.

### v2

@@
 Abstract.
-Efficient use of GPU memory is essential for high throughput LLM inference.
-Prior systems reserved memory for the KV-cache ahead-of-time, resulting in wasted capacity due to internal fragmentation.
-Inspired by OS-based virtual memory systems, vLLM proposed PagedAttention to enable dynamic memory allocation for KV-cache.
-This approach eliminates fragmentation, enabling high-throughput LLM serving with larger batch sizes.
+Efficient management of GPU memory is essential for high throughput LLM inference.
+Prior systems used to reserve KV-cache memory ahead-of-time that resulted in wasted capacity due to internal fragmentation.
+Inspired by demand paging, vLLM proposed PagedAttention to enable dynamic memory allocation for KV-cache.
+This approach eliminates fragmentation and improves serving throughout.
 However, to be able to allocate physical memory dynamically, PagedAttention changes the layout of KV-cache from contiguous virtual memory to non-contiguous virtual memory.
-This change requires attention kernels to be rewritten to support paging, and serving framework to implement a memory manager.
-Thus, the PagedAttention model leads to software complexity, portability issues, redundancy and inefficiency.
-In this paper, we propose vAttention for dynamic KV-cache memory management.
-In contrast to PagedAttention, vAttention retains KV-cache in contiguous virtual memory and leverages low-level system support for demand paging, that already exists, to enable on-demand physical memory allocation.
-Thus, vAttention unburdens the attention kernel developer from having to explicitly support paging and avoids re-implementation of memory management in the serving framework.
-We show that vAttention enables seamless dynamic memory management for unchanged implementations of various attention kernels.
-vAttention also generates tokens up to 1.97×1.97$\times$1.97 × faster than vLLM, while processing input prompts up to 3.92×3.92$\times$3.92 × and 1.45×1.45$\times$1.45 × faster than the PagedAttention variants of FlashAttention and FlashInfer.
-copyright: none
+As a consequence, one needs to rewrite the attention kernels to support paging, and implement a memory manager in the serving framework.
+This results in both performance and programming overheads, as well as portability challenges in adopting state-of-the-art attention kernels.
+In this paper, we propose vAttention, a new approach for dynamic KV-cache memory management.
+In contrast to PagedAttention, vAttention stores KV-cache in contiguous virtual memory and leverages OS support for on-demand allocation of physical memory.
+vAttention thus enables one to use state-of-the art attention kernels out-of-the-box by adding support for dynamic allocation of physical memory without having to re-write their code.
+We implement vAttention in the vLLM serving stack to show that it also helps improve decode throughput by up to 1.99×1.$9$ over vLLM, and the end-to-end serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
 1. Introduction
-Large Language Models (LLMs) are being deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (OpenAI, 2023; Chowdhery et al., 2022; cla, [n.
-d.]; bin, [n.
-d.]; bar, [n.
-d.]; git, [n.
-d.]; ama, [n.
-d.]; rep, [n.
-d.]).
-Optimizing LLM inference has thus become important (Agrawal et al., 2023; Holmes et al., 2024; Patel et al., 2023; Zhong et al., 2024; Hu et al., 2024; Kwon et al., 2023; Yu et al., 2022).
-One of the key techniques used to improve LLM serving throughput is batching (Yu et al., 2022; Kwon et al., 2023; Patel et al., 2023; Agrawal et al., 2024).
-Out of the two phases of LLM inference – namely prefill and decode – the decode phase is memory-bound because it processes a single token at-a-time per request.
-Batching amortizes the cost of fetching model weights from GPU memory and boosts throughput by improving memory bandwidth utilization (Agrawal et al., 2023).
+Large Language Models (LLMs) are deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (openai2022gpt4techreport, ; chowdhery2022Palm, ; bingai, ; bard, ; githubcopilot, ; amazoncodewhisperer, ; replitghostwriter, ).
+Given the size and scale of modern LLM deployments, optimizing inference has become extremely important  (sarathi2023, ; splitfuse2024, ; patel2023splitwise, ; distserve2024, ; tetriinfer, ; vllmsosp, ; orca, ).
+LLM inference happens in two phases: a compute-bound prefill phase, followed by several iterations of memory-bound decode phase.
+Batching is a powerful technique to boost LLM serving throughput (orca, ; vllmsosp, ; patel2023splitwise, ; sarathiserve2024, ), as it amortizes the cost of fetching model weights from HBM to the on-device SRAM during decodes.
 System/library and issues related to PagedAttention
 vLLM: Pioneered PagedAttention.
-Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.85×2.85$\times$2.85 × slower than FlashAttention (9(b)).
-Further, block size has as much as 3×3$\times$3 × effect on the execution time of the PagedAttention kernel (Figure 5).
-FlashAttention: PagedAttention version of the kernel is up to 12% slower than the vanilla version. Initial attempts to add paging support failed unit tests (fa-, 2023).
-TensorRT-LLM: Serving throughput dropped by more than 10% in a Python front-end (ten, 2023).
-Recommends using the C++ frontend.
-Even with C++, we observe up to 5% higher latency in some cases with PagedAttention.
-FlashInfer: Using PagedAttention-based prefill kernel increases time-to-first-token (TTFT) by up to 45% (Table 6).
+Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.8×2.8$\times$2.8 × slower than FlashAttention (Table 6).
+Further, changing block size changes the execute latency of the kernel by as much as 3×3$\times$3 × (Figure 3).
+FlashAttention: PagedAttention based prefill kernel is up to $28\%$ slower than the non-paged kernel (Figure 2) while the decode kernel is up to $12\%$ slower.
+Initial attempts to add paging support failed unit tests (fa-paged-crash, ).
+TensorRT-LLM: Serving throughput dropped by more than $10\%$ in a Python front-end (tensorrt-perf-decay, ).
+Recommends using the C++ front-end.
+Even with C++, we observe up to $5\%$ higher latency in some cases with PagedAttention.
+FlashInfer: PagedAttention based prefill kernel is up to $24\%$ slower than the non-paged kernel (Figure 2).
 Table 1.
-The PagedAttention model requires an application to explicitly manage dynamically allocated physical memory.
+The PagedAttention approach requires an application to explicitly manage dynamically allocated physical memory, including re-writing of attention kernels.
 These examples highlight the complexity, performance and maintenance challenges associated with this approach.
-Efficient inference also requires careful allocation of GPU memory.
-For every request, an LLM maintains an in-memory state known as the KV-cache and re-uses it in every iteration for the lifetime of the request (Yu et al., 2022; Agrawal et al., 2023; Patel et al., 2023).
-Achieving high memory capacity utilization during inference is challenging for two reasons: 1) per-request KV-cache grows slowly i.e., one token per iteration (few 10s of milliseconds) and 2) the number of generated tokens, and hence the total size of a request’s KV-cache, is typically not known ahead-of-time.
-Prior systems like Orca (Yu et al., 2022) and FasterTransformer (fas, 2024b) allocate a contiguous chunk of virtual memory (backed by pre-allocated physical memory) for the KV-cache.
-The allocated size corresponded to model’s maximum supported sequence length, e.g., 32K.
-Since models often generate far fewer tokens than the maximum limit, significant GPU memory was wasted due to internal fragmentation.
-Consequently, these systems exhibit poor throughput as they are unable to support large batch sizes.
-System Memory management in user code KV-cache fragmentation KV-cache allocation Portable wrt GPU code
-Programming Effort
-GPU CPU
-vLLM (Kwon et al., 2023) Block-Table(s) Low Fast No High High
-FlashInfer (Ye et al., 2024) Compressed Block-Table(s) Low Fast No High High
-vAttention None Low Slow Yes None Low
-Table 2.
-Dynamic memory management in various systems.
-TensorRT-LLM (trt, 2023), LightLLM (lig, 2023) and FlashAttention (Dao, 2023) also adopt vLLM-style PagedAttention model.
-FlashInfer uses a compressed representation and prefetching to limit the overhead of paging.
-We introduce various LLM-specific optimizations to hide the impact of high allocation latency in vAttention.
-Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (Kwon et al., 2023) to mitigate KV-cache related memory fragmentation.
-Instead of reserving the maximum sequence length of KV-cache memory ahead-of-time, vLLM allocates small blocks of virtual memory (backed by physical memory) on-demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens.
-However, dynamically allocated blocks are not guaranteed to be contiguous in virtual memory (the system may have allocated those to other requests).
-Thus, PagedAttention accepts that KV-cache allocations will be non-contiguous and implements a block-table to stitch together these non-contiguous allocations (§3.2).
-Today, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems e.g., in TensorRT-LLM (trt, 2023), HuggingFace TGI (hft, [n.
-d.]), FlashInfer (Ye et al., 2024), LightLLM (lig, 2023) etc.
-The most notable aspect of the PagedAttention approach is that it stores KV-cache in non-contiguous virtual memory to be able to allocate physical memory dynamically.
-While this approach provides an adequate solution for KV-cache fragmentation, we argue that it has several pitfalls (see Table 1 for empirical evidence and real-world experiences):
+Careful allocation of GPU memory is key to enabling larger batch sizes.
+For each request, the serving framework stores the activations of all the tokens processed so far in GPU memory and reuses them for the subsequent token generation.
+This is called the KV-cache (orca, ; sarathi2023, ; patel2023splitwise, ) which accounts for a majority of GPU memory usage during inference.
+Efficiently allocating GPU memory for the KV-cache of a batch of requests is challenging for two reasons.
+First, the per-request KV-cache grows slowly (one token per iteration), and second, a request’s decode length (or its total KV-cache size) is not known ahead of time.
+Prior systems like Orca (orca, ) and FasterTransformer (fastertransformer, ) allocate GPU memory for each request based on the maximum context length supported by the model (e.g., Yi-34B model supports context length of up to 200K (yi-34b-200k-hf, )).
+However, in practice, the number of decode tokens generated are far less; the average decode length for the chat-based sharegpt dataset is 415 tokens (sarathiserve2024, ).
+Therefore, static memory allocation results in severe internal fragmentation, limiting batch size and serving throughput.
+Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (vllmsosp, ) to mitigate KV-cache related memory fragmentation.
+Instead of reserving the maximum sequence length of KV-cache memory ahead-of-time, vLLM allocates small blocks of GPU memory on demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens.
+This approach provides a near-perfect solution for mitigating fragmentation and hence, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems e.g., in TensorRT-LLM (trtllmgithub, ), HuggingFace TGI (hftgi, ), LightLLM (lightllm:github, ) etc.
+However, the PagedAttention approach faces a fundamental consequence of dynamic memory allocation: dynamically allocated memory blocks are not guaranteed to be contiguous.
+Note that user-level objects are allocated in virtual memory.
+Therefore, in trying to enable dynamic allocation of physical memory, PagedAttention ends up changing the layout of KV-cache from contiguous to non-contiguous virtual memory.
+We argue that this approach has several pitfalls (see Table 1 for empirical evidence and real-world experiences):
 1.
 Requires re-writing the attention kernel (GPU code).
-The elements of a virtually contiguous object can be accessed using index-based lookup which is both simple and efficient [^1] A popular example of this is arrays vs.
-linked lists..
-By storing KV-cache in non-contiguous virtual memory, PagedAttention mandates re-writing GPU code so that the attention kernel can de-reference all the elements of KV-cache.
+The elements of a virtually contiguous object can be accessed using index-based lookup which is both simple and efficient.
+Many implementations of deep learning operators also expect virtual contiguity of input tensors.
+However, by storing KV-cache in non-contiguous virtual memory, PagedAttention mandates re-writing GPU code so that the attention kernel can de-reference all the elements of KV-cache.
 The need to re-write code is a major barrier to using new attention optimizations in production settings.
@@
 Adds software complexity and redundancy (CPU code).
-PagedAttention also forces the developers to implement a memory manager inside the serving framework, making it responsible for (de)allocating KV-cache and tracking the location of dynamically allocated KV-cache blocks.
-This approach essentially translates to re-implementing demand paging – which is an OS functionality – in user code.
+To stitch together dynamically allocated virtual memory blocks, PagedAttention forces developers to implement a memory manager in the serving framework.
+Typically, this requires a Block-Table manager in user space which is essentially a re-implementation of demand paging (an OS functionality) in user code.
 3.
 Introduces performance overhead.
-PagedAttention can add runtime overhead in the critical path of execution in two ways.
-First, it requires GPU kernels to execute extra code related to fetching KV-cache from non-contiguous memory blocks.
-We show that this can slow down attention computation by more than 10% in many cases.
-Second, the user space memory manager can add CPU overhead, contributing up to another 10% cost (§3.3).
-Figure 1.
-Illustration of memory waste due to internal fragmentation.
-Orca (top) reserves memory for the KV-cache ahead-of-time.
-vLLM (bottom) mitigates fragmentation with dynamic memory allocation.
-Shaded boxes represent memory occupied by the KV tokens whereas the white boxes represent allocated but unused memory.
-In this paper, we argue that retaining the virtual memory contiguity of KV-cache is critical for reducing software complexity and redundancy in LLM deployments.
-Instead of re-implementing paging at the user-level, we contend that existing virtual memory abstractions in the OS can be re-purposed for dynamic KV-cache memory management, resulting in simplified deployments as well as higher performance.
-To demonstrate this, we propose vAttention – a system that stores KV-cache in contiguous virtual memory without committing physical memory ahead-of-time.
-We achieve this by leveraging CUDA support of low-level virtual memory APIs which expose distinct interfaces for allocating virtual and physical memory (§5).
-vAttention exposes a set of simple APIs using which a serving framework reserves contiguous virtual space for the KV-cache and allocates physical memory on-demand.
-This approach lends several benefits as listed in Table 2.
-vAttention also improves portability by enabling a seamless re-use of readily available GPU kernels while eliminating the need to implement a memory manager in a serving system.
-Challenges and Optimizations: vAttention solves two key challenges in enabling efficient dynamic memory management without PagedAttention (§6).
-First, the minimum physical memory allocation granularity supported by CUDA APIs is 2MB.
-This size can result in significant wasted capacity depending on the model and workload characteristics (Table 8).
-To address this, we modify the open-source CUDA unified virtual memory driver to add support for finer-grained physical memory allocations of 64KB to 256KB.
-Second, memory allocation using CUDA APIs incurs high latency because each allocation involves a round-trip to the kernel.
-To hide the latency of memory allocation from end-users, we introduce several LLM-specific optimizations such as overlapping memory allocation with compute, executing some operations ahead of time and deferring memory reclamation.
-We show that our optimizations make vAttention an efficient KV-cache memory manager.
+PagedAttention adds runtime overhead in the critical path of execution in two ways.
+First, GPU kernels that compute attention over a non-contiguous KV-cache need to execute additional instructions.
+In many cases, this slows down an attention kernel by more than 20%, compared to when the KV-cache is virtually contiguous.
+Second, the user space memory manager can add CPU overhead contributing up to another 10% cost (§3.3).
+In this paper, we instead advocate for leveraging the OS support of virtual memory and demand paging for dynamic KV-cache memory management.
+This approach alleviates the aforementioned issues that appear with PagedAttention.
+To support our claims, we present the design and implementation of vAttention.
+Leveraging OS support, vAttention stores KV-cache in contiguous virtual memory without committing physical memory ahead-of-time, thereby enabling the use of state-of-the art high-performant attention kernels out-of-the box.
+We achieve this by using CUDA support of low-level virtual memory APIs which expose distinct interfaces for allocating virtual and physical memory (§5).
+Leveraging OS support for memory allocation in an LLM serving system poses two key efficiency challenges (§5.4).
+First, the minimum physical memory allocation granularity supported by CUDA is 2MB.
+This can result in significant wasted capacity and fragmentation.
+We address this challenge by modifying the open-source CUDA unified virtual memory driver, adding support for finer-grained physical memory allocations in multiples of 64KB.
+Our evaluation shows that using 64KB pages as the granularity of virtual-to-physical address translation does not affect the execution latency of attention kernels i.e., we do not find any evidence of TLB thrashing.
+Second, memory allocation using CUDA APIs incurs high latency because each allocation involves a round-trip to the OS kernel.
+We introduce several LLM-specific optimizations such as overlapping memory allocation with compute, opportunistically allocating pages ahead of time, and deferring memory reclamation.
+These optimizations hide the latency cost of memory allocation from end-users, making vAttention an efficient KV-cache memory manager.
 Overall, we make the following contributions in this paper:
-  • We present vAttention, a system that retains the virtual contiguity of KV-cache while enabling dynamic allocation of physical memory.
-  • We show that vAttention is able to seamlessly add dynamic memory management support to unmodified attention kernels of FlashAttention (fag, 2022) and FlashInfer (fig, 2023) while also being performant.
-  • We evaluate Yi-6B, Llama-3-8B and Yi-34B on 1-2 A100 GPUs and show that using FlashAttention’s original kernel, vAttention outperforms vLLM by up to 1.97×1.97$\times$1.97 ×, while reducing the time-to-first-token (TTFT) by up to 1.45×1.45$\times$1.45 × over the PagedAttention variant of FlashInfer.
+  • We present vAttention – a memory management approach that retains the virtual contiguity of KV-cache while enabling dynamic physical memory allocation.
+  • We implement vAttention in vLLM serving stack to show that it seamlessly adds dynamic memory management support to unmodified attention kernels of FlashAttention (fagithub, ) and FlashInfer (figithub, ).
+  • We compare vAttention against PagedAttention based alternatives of vLLM, FlashAttention and FlashInfer by evaluating Yi-6B, Llama-3-8B and Yi-34B on 1-2 A100 GPUs.
+Using FlashAttention’s non-paged kernel, vAttention outperforms vLLM by up to 1.99×1.$9$ in decode throughput, while improving the end-to-end LLM serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
 2. Background
 2.1. Large Language Models
-Given an input sequence, an LLM predicts the probability of an output sequence wherein a sequence is a set of tokens (Kwon et al., 2023).
+Given an input sequence, an LLM predicts the probability of an output sequence wherein a sequence is a set of tokens (vllmsosp, ).
 Each inference request begins with a prefill phase that processes all its prompt tokens in parallel.
 The prefill phase produces the first output token of a request.
-Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (Agrawal et al., 2023).
-LLMs are built atop one of variants of the transformer architecture (Vaswani et al., 2017).
+Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (sarathi2023, ).
+LLMs are built atop one of variants of the transformer architecture (attentionpaper, ).
 A transformer block contains two types of operators: position-wise and sequence-wise.
@@
 The attention score is computed separately for each request in the batch.
-Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention [^2] An exception to this is sliding window attention that fixes an upper bound on how many of the prior tokens are used in computing attention..
+Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention.
 Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations: we refer to this state as KV-cache.
@@
 In deep learning frameworks, the K-cache (or V-cache) at each layer is typically represented as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request.
-We refer to a kernel implementation that computes attention scores over contiguously stored K and V as a vanilla kernel.
+We refer to a kernel implementation that computes attention scores over contiguously stored K and V as a non-paged kernel.
 2.2. Fragmentation and PagedAttention
@@
 This is challenging because the total context length of a request is not known in advance.
-Serving systems worked around this challenge by pre-reserving KV-cache space assuming that each context is as long as the maximum length supported by the model (e.g.,200K for Yi-6B-200K).
+Serving systems worked around this challenge by pre-reserving KV-cache space assuming that each context is as long as the maximum length supported by the model (e.g., 200K for Yi-34B-200K).
 vLLM shows that this strategy is prone to severe internal fragmentation.
-In fact, vLLM showed that prior reservation wastes memory even if the context lengths are known in advance.
+In fact, vLLM showed that prior reservation is suboptimal even if the context lengths are known in advance.
 This is because the per-request KV-cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
@@
 This way, vLLM allocates only as much memory as a request needs, and only when required – not ahead-of-time.
-Figure 1 shows an example of how reservation-based systems such as Orca (Yu et al., 2022) can waste significant memory due to fragmentation and how vLLM avoids it with dynamic memory allocation.
-Figure 2.
-PagedAttention involves two layers of address translation.
-First, the framework tracks the virtual memory addresses of KV-cache blocks via Block-Tables.
-Second, the GPU performs virtual-to-physical address translation via OS-managed Page-Tables.
+Figure 1. PagedAttention involves two layers of memory management: one in user space and one in OS kernel space.
 3. Issues with the PagedAttention Model
-Despite being inspired by demand paging, PagedAttention adopts an approach that is different from conventional demand paging: it requires an application’s code to be modified to adapt to dynamically allocated physical memory whereas conventional demand paging is transparent to applications.
-This section elaborates on some of the issues that arise with such an approach.
+Despite being inspired by demand paging, the PagedAttention approach is different from it: PagedAttention mandates an implementation of paging in user space whereas conventional demand paging is transparent to applications.
+This section elaborates on issues that arise with such an approach.
 3.1. Requires Re-writing the Attention Kernel
@@
 By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV-cache blocks.
-Writing correct and performant GPU kernels can be challenging for most programmers (fa-, 2023).
-Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (Dao et al., 2022; Dao, 2023; fla, 2023; Hong et al., 2024; Ye et al., 2024; Zhang et al., 2023; Bikshandi and Shah, 2023; Ainslie et al., 2023; Shazeer, 2019; Child et al., 2019; Beltagy et al., 2020), and this trend is likely to continue.
+Writing correct and performant GPU kernels can be challenging for most programmers (fa-paged-crash, ).
+Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (flashattention, ; flashattention2, ; flashdecoding, ; flashdecoding++, ; flashinfer, ; h2o2023, ; flashattnhopper, ; groupedqueryattention, ; multiqueryattention, ; generatinglongsequence2019, ; longformer, ), and this trend is likely to continue.
 In the PagedAttention model, keeping up with new research requires continued efforts in porting new optimizations to a PagedAttention-aware implementation.
 Production systems can therefore easily fall behind research, potentially losing performance and competitive advantage.
-To provide an example, 9(b) shows that the paged kernel of vLLM is already up to 2.85×2.85$\times$2.85 × slower than the FlashAttention counterpart for grouped-query attention (Ainslie et al., 2023).
-Figure 3. Overhead of PagedAttention in FlashAttention’s decode kernel (model: Yi-6B). Legend shows the batch size.
+To provide an example, Table 6 shows that the paged kernel of vLLM is already up to 2.8×2.8$\times$2.8 × slower than the FlashAttention kernel (groupedqueryattention, ).
 3.2. Adds Redundancy in the Serving Framework
@@
 To facilitate this, the serving system needs to track the virtual memory addresses of KV-cache blocks and pass them to the attention kernel at runtime.
-This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half Figure 2).
+This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half in Figure 1).
+Figure 2.
+Overhead of PagedAttention in prefill kernels (model: Llama-3-8B, one A100 GPU).
+Numbers on top show overhead over the corresponding non-paged implementation of FlashAttention (FA) and FlashInfer (FI).
 3.3. Performance Overhead
-PagedAttention also leads to potential performance issues both on GPU and CPU. We investigate them separately.
 3.3.1. Runtime overhead on the GPU
 PagedAttention slows down attention computation by adding extra code in the critical path.
-For example, vLLM acknowledges that their PagedAttention-based implementation was $20-26\%$ slower than the original FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (Kwon et al., 2023).
-Figure 3 shows that the paged decode kernel in FlashAttention is also slower than the vanilla kernel.
-Our further analysis reveals that the number of instructions executed in PagedAttention is up to $13\%$ higher than the vanilla kernel.
-We also find that the overhead of paging reduces at high batch sizes or long context lengths.
-This is because computing attention for decodes is memory bound and when the KV-cache size is large, memory stalls hide the instruction overhead.
-Figure 4. Overhead of PagedAttention in FlashInfer’s prefill kernel (model: Yi-6B). Legend shows the batch size.
-However, hiding instruction overhead is more challenging in a prefill attention kernel which is compute-bound with $N^{2}$ complexity.
-For example, Figure 4 shows that the paged version of FlashInfer (Ye et al., 2024) prefill kernel is up to 14% slower than the vanilla kernel.
-The paged kernel is also missing some well-known optimizations (fip, 2024).
-To highlight another example of difficulty involved in writing an efficient attention kernel, Figure 5 shows that the performance of vLLM’s paged decode kernel is significantly worse with larger block sizes of 64 and 128.
+For example, vLLM acknowledges that their PagedAttention-based implementation was $20-26\%$ slower than the original FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (see Figure 18a in  (vllmsosp, )).
+In addition, Figure 2 shows that incorporating PagedAttention has also added a significant performance overhead in other state-of-the-art kernel libraries.
+For example, PagedAttention based prefill kernels of FlashAttention and FlashInfer are up to 28% and 24% slower than the non-paged kernels in the corresponding libraries.
+Our analysis reveals that the number of instructions executed in PagedAttention kernels is $7-13\%$ higher than the non-paged kernels.
+To highlight another example of difficulty involved in writing an efficient attention kernel with paging support, Figure 3 shows that the performance of vLLM’s paged decode kernel is significantly worse with larger block sizes of 64 and 128.
 Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in L1 cache.
@@
 In our earlier experiments, we observed that Block-Table preparation in vLLM was contributing 30% latency in decode iterations.
-While a recent fix (vll, 2024a) has mitigated some of this overhead, we find that it can still be as high as $10\%$.
-High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11%, from 412 tokens/sec to 365 tokens/sec (trt, 2023).
+While a recent fix (vllm-blocktable-issue, ) has mitigated some of this overhead, we find that it can still be as high as $10\%$.
+High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11%, from 412 tokens/sec to 365 tokens/sec (tensorrt-perf-decay, ).
 This issue was attributed to the Python runtime of TensorRT-LLM and moving to a C++ runtime can mitigate the CPU overhead.
+However, doing so would be an unwelcome change for most programmers.
+Figure 3. Latency of vLLM’s paged decode kernel is sensitive to block size (model: Llama-3-8B, one A100 GPU).
 Overall, this section shows that the PagedAttention model adds a significant programming burden while also being inefficient.
 vAttention introduces a more systematic approach to dynamic KV-cache memory management by leveraging the existing system support for demand paging.
-However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads in terms of memory management.
-Figure 5. Latency of vLLM’s paged decode kernel is sensitive to block size (model: Yi-6B). Legend shows the block size.
-(a) Output tokens per second.
-(b) Decode iteration latency.
-(c) Memory allocation bandwidth.
-Figure 6.
-LLM inference throughput plateaus at large batch sizes (a) beyond which latency suffers (b) without improving throughput.
-This also means that the rate at which physical memory needs to be allocated also saturates (c).
+However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads from a memory management perspective.
 4. Insights into LLM Serving Systems
-To highlight the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single NVIDIA A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism.
-We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure latency, throughput and memory requirement of the decode phase (see §6 for our discussion and optimizations for the prefill phase).
+To highlight the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism.
+We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure the throughput and memory requirement of the decode phase (see §5.4 for our discussion and optimizations for the prefill phase).
 Observation-1: On a per-iteration basis, KV-cache memory requirement is known in advance.
-This is due to auto-regressive decoding that generates one token at a time, per-request.
-Therefore, with every iteration, the KV-cache memory footprint of a request grows uniformly by one token until the request completes.
+This is a dream property for a memory allocator, thanks to the auto-regressive decoding wherein every forward pass of the model generates one token per request.
+Therefore, with every iteration, the KV-cache memory footprint of a request grows uniformly by one token as long as the request is active.
 Observation-2: KV-cache does not require high memory allocation bandwidth.
@@
 For example, the per-token memory footprint of Yi-6B, Llama-3-8B and Yi-34B is 64KB, 128KB and 240KB, respectively.
-Further, each iteration runs for 10s-100s of milliseconds (6(b)) implying that a request requires at most a few megabytes of memory per second.
-While batching improves system throughput, the number of tokens generated per second plateaus beyond a certain batch size (6(a)).
+Further, each iteration runs for 10s-100s of milliseconds implying that a request requires at most a few megabytes of memory per second.
+While batching improves system throughput (sarathi2023, ; sarathiserve2024, ; patel2023splitwise, ; distserve2024, ), the number of tokens generated per second plateaus beyond a certain batch size (Figure 4, left).
 This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 128 for Yi-34B).
-For all the models we studied, we observe that the highest memory allocation rate is at most 600MB per second (6(c)).
+For all the models we studied, we observe that the highest memory allocation rate is at most 600MB per second (Figure 4, right).
+Figure 4. Decode throughput (left) and the rate of physical memory allocation (right) saturate at large batch sizes.
 In vAttention, we leverage these observations to implement an efficient dynamic memory management system for KV-cache.
-In the next section, we begin with a high-level design description of vAttention (§5.1), then discuss how vAttention is used for serving LLMs (§5.3) and finally describe our optimizations (§6).
-5. vAttention: System Design
-Our goal is to improve efficiency and portability by adding dynamic memory allocation support to existing kernels.
-To achieve this goal, vAttention leverages system support for dynamic memory allocation instead of implementing paging in user space.
+In the next section, we begin with a high-level overview of vAttention (§5.1), then discuss how vAttention is used for serving LLMs (§5.3) and finally describe our optimizations (§5.4).
+5. vAttention: Design and Implementation
+To facilitate dynamic allocation of physical memory to unmodified attention kernels, vAttention leverages system support for demand paging instead of implementing it in user space.
 5.1. Design Overview
 vAttention builds on the ability to allocate virtual memory and physical memory separately.
-Specifically, we allocate a large contiguous buffer for the KV-cache in virtual memory ahead-of-time (similar to reservation-based allocators) while deferring the allocation of physical memory to runtime i.e., allocate physical memory only when required (similar to PagedAttention).
+Specifically, we allocate a large contiguous buffer for the KV-cache in virtual memory ahead-of-time (similar to reservation-based allocators) while deferring the allocation of physical memory to runtime (similar to PagedAttention).
 This way, vAttention preserves virtual contiguity of KV-cache without wasting physical memory.
-This approach is feasible because memory capacity and fragmentation are limiting factors only for physical memory whereas virtual memory is abundant e.g., modern 64-bits systems provide a 128TB user-managed virtual address space to each process33364-bits systems typically utilize 48 bits for virtual addresses, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space..
+This approach is feasible because memory capacity and fragmentation are limiting factors only for physical memory whereas virtual memory is abundant e.g., modern 64-bits systems provide a 128TB user-managed virtual memory per process11164-bits systems use only 48 bits for virtual addresses today, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space..
 Figure 5.
@@
 (e) when R3 arrives, vAttention assigns R1’s tensor to it which is already backed by physical memory.
-5.1.1. Pre-reserving virtual memory
-Since virtual memory is abundant, we pre-allocate enough virtual memory space that is large enough to hold the KV-cache of the maximum batch size (configurable) that needs to be supported.
-Number of virtual memory buffers: Each layer in an LLM maintains its own K and V tensors: we refer to them individually as K-cache and V-cache.
+Pre-reserving virtual memory
+Since virtual memory is abundant, we pre-allocate it in size that is large enough to hold the KV-cache of the maximum batch size (configurable) that needs to be supported.
+In doing so, we assume that each request would generate as many tokens as the maximum sequence length supported by the model.
+Number of virtual memory buffers
+A serving framework typically maintains separate K and V tensors for each layer of the model: we refer to them individually as K-cache and V-cache.
 We allocate separate virtual memory buffers for K-cache and V-cache.
@@
 In a multi-GPU job, each worker reserves $2\times N^{\prime}$ buffers where $N^{\prime}$ is the number of layers managed by that worker ($N^{\prime}=N$ with tensor-parallelism whereas $N^{\prime}<N$ with pipeline-parallelism).
-Size of a virtual memory buffer: The maximum size of a buffer is $BS=B\times L\times S$ where B is the maximum batch size, L is the maximum context length supported by the model, and $S$ is the size of a single token’s per-layer K-cache (or V-cache) on a worker.
-Further, $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16).
-Note that $S$ is constant for a given model configuration.
-Consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2).
-In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$.
-For this model, the maximum size of K-cache (or V-cache) per-worker per-layer is $S=200MB$ $(200K*4*128*2)$.
-Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$).
-Therefore, the total virtual memory requirement for 60 layers of Yi-34B is 120 buffers of 100GB each (12TB total).
-Note that the amount of virtual address space available grows with the number of GPUs e.g., with two TP workers, the amount of virtual address space available is 256TB.
-Therefore, virtual memory allocations can be satisfied easily.
-5.1.2. On-demand physical memory allocation
-vAttention preferentially allocates physical memory one page at a time and only when a request has used all of its previously allocated physical memory pages.
-To show how it works, we refer to a simple example in Figure 7.
-The example shows how vAttention manages the K-cache (or V-cache) at one layer of the model assuming maximum batch size of two.
-Rest of the K-cache and V-cache buffers are managed similarly at all layers.
-5.2. Leveraging Low-level CUDA Support
-The standard GPU memory allocation interface cudaMalloc does not support demand paging i.e., it allocates virtual memory and physical memory at the same time.
-However, recent CUDA versions provide programmers a fine-grained control over virtual and physical memory (cud, 2024; Guo et al., 2024).
-We leverage these low-level APIs in vAttention.
-5.2.1. CUDA virtual memory APIs
-Table 3 provides a high-level overview of CUDA APIs that allow separating the allocation of virtual memory from physical memory (see the leftmost column).
-The allocation granularity depends on the page size used by the GPU and the size of virtual memory buffer or a physical memory handle must be a multiple of the allocation granularity.
-Different sub-regions of a virtual memory buffer can be backed by physical memory independently of other sub-regions in that buffer (see Figure 7c for an example).
-For simplicity, we refer to the granularity at which physical memory is allocated as page size.
 Latency (microseconds)
-CUDA APIs vAttention APIs Description 64KB 128KB 256KB 2MB
+CUDA VM APIs vAttention VM APIs Description 64KB 128KB 256KB 2MB
 cuMemAddressReserve * vMemReserve * Allocate a buffer in virtual memory 18 17 16 2
@@
 Table 2.
-Low-level APIs for virtual memory management and their latency with different allocation sizes.
+CUDA VM (virtual memory) APIs.
 * represents APIs that we use once while instantiating or terminating the serving framework.
@@
 CUDA APIs (prefixed with cu) support only 2MB allocation sizes, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
-5.2.2. Extending PyTorch caching allocator:
+Size of a virtual memory buffer
+The maximum size of a buffer is $BS=B\times L\times S$ where B is the maximum batch size, L is the maximum context length supported by the model, and $S$ is the size of a single token’s per-layer K-cache (or V-cache) on a worker.
+Further, $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16).
+Note that $S$ is constant for a given model configuration.
+Consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2).
+In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$.
+For this model, the maximum size of K-cache (or V-cache) per-worker per-layer is $S=200MB$ $(200K*4*128*2)$.
+Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$).
+Therefore, the total virtual memory requirement for 60 layers is 120 buffers of 100GB each (12TB total).
+Note that the amount of virtual address space available grows with the number of GPUs e.g., with two TP workers, the amount of virtual address space available is 256TB.
+Therefore, virtual memory allocations can be satisfied easily.
+Figure 5 shows how vAttention allocates physical memory pages dynamically.
+5.2. Leveraging Low-level CUDA Support
+The standard GPU memory allocation interface cudaMalloc does not support demand paging i.e., it allocates virtual memory and physical memory at the same time.
+However, recent CUDA versions provide programmers a fine-grained control over virtual and physical memory (cudavirtualmemory, ; gmlakeasplos24, ).
+We leverage these low-level APIs in vAttention.
+5.2.1. CUDA virtual memory APIs
+Table 2 provides an overview of CUDA APIs that allow separating the allocation of virtual memory from physical memory.
+The allocation granularity depends on the page size used by the GPU.
+Further, the size of a virtual memory buffer or a physical memory handle must be a multiple of the physical memory allocation granularity.
+It is also important to highlight that physical memory pages can be allocated (or de-allocated) to sub-regions in a virtual memory buffer independently of other sub-regions.
+For simplicity, we refer to the granularity of physical memory allocation as page size.
+5.2.2. Extending PyTorch caching allocator
 KV-cache is a collection of tensors.
@@
 We refer to tensors allocated via these APIs as virtual tensors.
-5.2.3. Request-level KV-cache indexing:
+5.2.3. Request-level KV-cache indexing
 Note that each virtual tensor represents the K-cache (or V-cache) of a layer for the maximum batch size B.
@@
 We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers.
-Our library exposes a set of simple APIs listed in Table 4 to the serving framework.
-5.3.1. Initial setup:
+Our library exposes a set of simple APIs to the serving framework (see Table 3 and Algorithm 1).
+5.3.1. Initial setup
 When the serving framework starts, each model worker loads the vAttention library and configures it with model parameters $N^{\prime},H,D,P$, $B$ and a preferred page size via the init API.
@@
 However, these pages are not mapped into the KV-cache yet.
+5.3.2. Scheduling a new request
+When a new request is scheduled for the first time, the serving framework obtains a new reqId from vAttention via alloc_reqid.
+All subsequent memory management operations of the request are tagged with this reqId.
 APIs Description
@@
 return value: 0 (success), -1 (failure).
-Table 4. APIs exposed to a serving framework for dynamic KV-cache memory management with vAttention.
-5.3.2. Scheduling a new request:
-When a new request is scheduled for the first time, the serving framework obtains a new reqId from vAttention via alloc_reqid.
-All subsequent memory management operations of the request are tagged with this reqId.
-5.3.3. Model execution:
+Table 3. Key APIs that vAttention exposes to a serving framework for dynamic KV-cache memory management.
+5.3.3. Model execution
 Before scheduling a batch for execution, the framework needs to ensure that the KV-cache sub-tensors of each active request are backed by physical memory.
@@
 vAttention internally tracks the number of pages mapped for each request and maps a new page only when the last page allocated to that request is fully utilized.
-5.3.4. Request completion.
+5.3.4. Request completion
 A request terminates when a user specified context length or the maximum context length supported by the model is reached, or when the model produces a special end-of-sequence token.
@@
 Internally, vAttention may unmap the pages of a completed request or defer them to be freed later.
-6. vAttention: Optimizations
-There are two primary challenges in using CUDA’s virtual memory support for serving LLMs.
+Algorithm 1 Using vAttention in a serving framework.
+1:max_batch_size $\leftarrow$ B
+2:cache_seq_len $\leftarrow$ [0]*B
+3:req_batch_idx $\leftarrow$ dict()
+4:vattention.init_cache(config_params)
+5:while !request_pool.is_empty() do
+6:     for $R_{i}$ in new_requests do
+7:         if can_schedule($R_{i}$then
+8:              idx $\leftarrow$ vattention.alloc_reqid()
+9:              req_batch_idx[$R_{i}$] $\leftarrow$ idx
+10:              cache_seq_len[idx] $\leftarrow$ prompt_len($R_{i}$)
+11:         end if
+12:     end for
+13:     vattention.step(cache_seq_len)
+14:     model.forward()
+15:     for $R_{i}$ in active_requests do
+16:         idx $\leftarrow$ req_batch_idx[$R_{i}$]
+17:         if is_complete($R_{i}$then
+18:              cache_seq_len[idx] $\leftarrow$ 0
+19:              vattention.free_reqid(idx)
+20:         else
+21:              cache_seq_len[idx] +=+=+ = 1
+22:         end if
+23:     end for
+24:end while
+5.4. Optimizations
+There are two primary challenges in using CUDA virtual memory support for serving LLMs.
 First, cuMemCreate currently allocates a minimum of 2MB physical memory page.
@@
 This section details a set of simple-yet-effective optimizations that we introduce to overcome these limitations.
-6.1. Mitigating internal fragmentation
+5.4.1. Mitigating internal fragmentation
 We mitigate internal fragmentation by reducing the granularity of physical memory allocation.
-NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB.
+NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB (mcmgpus, ; gpu-mismanaged, ; gpu-reverse-engg-tlb1, ; pascalmmu, ).
 Therefore, in principal, physical memory can be allocated in any multiple of 4KB sizes.
@@
 In contrast to CUDA APIs, our APIs can allocate memory in 64KB, 128KB and 256KB page sizes.
-A serving framework can configure a desired page size in vAttention while initializing it: we recommend using 256KB pages by default.
+A serving framework can configure a desired page size in vAttention while initializing it.
 The last set of columns in Table 2 shows the latency of each API with different page sizes.
-6.2. Hiding memory allocation latency
+5.4.2. Hiding memory allocation latency
 The serving framework invokes the step API in every iteration.
@@
 Consider, for example, that the KV-cache of one request needs to be extended for Yi-34B which has 60 layers.
-This requires 120 calls to vMemMap each of which takes about 9 microseconds.
-Therefore, growing the KV-cache of one request by one page will add about 1 millisecond latency to the corresponding iteration and would grow proportional to amount of physical memory that needs to be mapped.
+This requires 120 calls to cuMemMap + cuMemSetAccess each of which takes about 40 microseconds.
+Therefore, growing the KV-cache of one request by new pages (at all layers) adds about 5 millisecond latency to the corresponding iteration.
+Further, the latency overhead grows proportional to the number of requests that need new pages in a given iteration.
 We propose the following optimizations to hide the latency of allocation:
-6.2.1. Overlapping memory allocation with compute
+Overlapping memory allocation with compute
 We leverage the predictability of memory demand to overlap memory allocation with computation.
@@
 If not, required pages are mapped synchronously.
-6.2.2. Deferred reclamation + eager allocation
+Deferred reclamation + eager allocation
 We observe that allocating physical memory for a new request can be avoided in many cases.
@@
 To avoid allocating new pages to R2 from scratch, vAttention simply defers the reclamation of R1’s pages and assigns R1’s reqId to R2.
-This way, R2 uses the same tensors for its KV-cache that R1 was using which are already backed by physical pages.
-Therefore, new pages for R2 are required only if its context length is bigger than that of R1.
-We further optimize memory allocation by proactively mapping physical pages before they are needed.
-We do so by using one of the inactive reqId’s KV-cache.
-When a new request arrives, we can allocate this reqId without mapping any physical pages.
-We then select a new reqId that would be allocated next and map physical pages for it.
-In most cases, these eager optimizations obviate the need to allocate new physical pages even for the prefill phase of new requests.
-Finally, we trigger memory reclamation only when the number of physical memory pages cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory).
+This way, R2 uses the same tensors for its KV-cache that R1 was using – which are already backed by physical pages (for security reasons, we still zero-fill such pages before allowing a new request to use them).
+Therefore, new pages are required only if the context length of R2 is bigger than that of R1.
+We further optimize memory allocation by proactively allocating a small number of physical pages ahead of time.
+For this purpose, we try to keep a certain number of pages (configurable based on expected workload) mapped into the virtual tensors of one of the inactive reqId.
+When a new request arrives, we allocate this reqId and identify a new reqId to be allocated next and map physical pages for it.
+In most cases, these eager allocations obviate the need to allocate physical pages in the critical path of prefill execution.
+Finally, we trigger memory reclamation only when the number of physical pages cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory).
 We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
+6. Evaluation
 Model Hardware ##\## Q Heads ##\## KV Heads ##\## Layers
@@
 Table 4. Models and hardware used for evaluation.
-7. Evaluation
+Figure 6.
+Prefill throughput.
+vAttention backed systems outperform the paged counterparts of both FlashAttention and FlashInfer.
+Throughput for longer contexts is lower due to the quadratic complexity of prefill attention.
 Our evaluation seeks to answer the following questions:
-  • How does vAttention perform for prefill and decode phases in LLM inference? What are the portability and performance advantages of vAttention.
-  • How efficiently can vAttention allocate GPU memory for LLM serving workloads, and how effectively can it deal with KV-cache fragmentation?
-Models and Hardware: We evaluate three models Yi-6B, Llama-3-8B and Yi-34B, using a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 5).
+  • How does vAttention perform for prefill and decode phases (§6.1, §6.2)?
+  • How vAttention impacts LLM serving throughput (§6.3)?
+  • What is the effect of each of our optimizations (§6.4)?
+Models and hardware: We evaluate three models Yi-6B (yi-6b-200k-hf, ), Llama-3-8B (llama-8b-hf, ) and Yi-34B (yi-34b-200k-hf, ), using a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 4).
 Each GPU has 80GB physical memory.
 We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B.
-All three models use GQA which is the most commonly used attention mechanism in recent LLMs.
-Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different.
-Attention kernels used for these two phases are also different and hence we evaluate them separately.
+Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different (sarathiserve2024, ; splitfuse2024, ; distserve2024, ).
+Attention kernels used for these two phases are also different and hence we also evaluate them separately.
 The prefill phase requires one time memory allocation potentially spanning multiple pages.
-In comparison, the decode phase requires incremental memory allocation over the lifetime of a request.
-We measure the throughput of these phases in terms of tokens processed (or generated) per second.
-Figure 8.
-Prefill tokens processed per second with chunking (chunk size=2048).
-Using vanilla prefill kernels, vAttention outperforms the paged counterpart of both FlashAttention and FlashInfer.
-Throughput for longer context lengths drops sharply due to quadratic complexity of prefill attention.
-7.1. Portability and Performance for Prefills
-To evaluate the prefill phase, we focus on the attention kernels provided by FlashAttention v2.5.6 (Dao, 2023; fag, 2022) and FlashInfer v0.0.3 (fig, 2023; Ye et al., 2024).
-We do not include vLLM in these experiments because it does not have a prefill kernel of its own but instead uses FlashAttention kernel.
-We also could not evaluate Yi-34B because FlashInfer kernels do not support Yi-34B’s KV group size of 7 (fip, 2024).
-FlashInfer is a library that recently introduced a set of attention kernels optimized for different scenarios e.g., for chunked-prefills – an optimization proposed in Sarathi (Agrawal et al., 2023) and later adopted in various systems (Agrawal et al., 2024; Holmes et al., 2024; Hu et al., 2024).
-Sarathi splits the input tokens of a prompt into multiple smaller chunks and schedules one chunk at a time, enabling a serving system to add new requests in a batch without pausing ongoing decodes.
-This helps improve throughput without increasing latency (Agrawal et al., 2024).
-Both FlashAttention and FlashInfer provide kernels to compute the attention scores of chunked-prefills with and without PagedAttention.
-We integrate them into vLLM and using chunk size of 2048 tokens, measure time-to-first-token (TTFT) for the following configurations:
-FA_Paged: Uses flash_attn_with_kv_cache kernel API of FlashAttention.
-FI_Paged: Uses FlashInfer’s PagedAttention kernel, representing state-of-the-art PagedAttention-based kernel for the prefill phase.
-FA_vAttention: Uses FlashAttention’s vanilla prefill kernel via the flash_attn_func API.
-FI_vAttention: Uses FlashInfer’s vanilla prefill kernel via the single_prefill_with_kv_cache API.
-Model Context Length FlashAttention FlashInfer
-Paged vAttention Paged vAttention
-Yi-6B 16K 2.44 1.38 1.51 1.34
-32K 8.07 3.50 3.97 3.4
-64K 30.46 9.98 11.85 9.66
-128K 119 32.10 39.17 30.9
-192K 266 67.81 81.87 65.03
+In comparison, the decode phase requires incremental memory allocation over the lifetime of a request (vllmsosp, ).
+We define throughput as the number of tokens processed (or generated) per second.
+Serving framework: For a fair comparison, we use vLLM v0.2.7 as a common serving framework in all our experiments.
+We integrated state-of-the-art kernel libraries of both FlashAttention v2.5.9 (flashattention, ; flashattention2, ; fagithub, ) and FlashInfer v0.4.0 (figithub, ) as attention back-ends into vLLM, and further added support for dynamic memory allocation via vAttention to their non-paged kernels.
+While FlashAttention and FlashInfer are both based on the same underlying techniques (e.g., FlashDecoding (flashdecoding, )), they use different Block-Table formats; the former use a simple lookup table whereas the latter uses a compressed Block-Table to aid fetching KV-cache block locations.
+Baselines: We compare performance obtained by using the non-paged attention kernels (backed by vAttention for dynamic memory allocation), and their paged counterparts.
+In addition, we also compare against vLLM’s decode kernel (note that vLLM does not have a paged prefill kernel).
+For a fair comparison, we also profiled each system to find its best performing configuration.
+Accordingly, we set KV-cache block size to 16 for both vLLM and FlashInfer, and 256 for FlashAttention.
+Using a higher block size for vLLM increases its kernel latency by up to 3×3$\times$3 × as shown in Figure 3, and using a smaller block size for FlashAttention paged kernel increases its latency by up to $9\%$ (Table 9).
+We find that the choice of page size does not affect vAttention as long as fragmentation is not a concern (Figure 11).
+Model Context Length FA_Paged FA_vAttention FI_Paged FI_vAttention
+Yi-6B 64K 10.6 (7.0) 9.1 (5.5) 10.9 (6.0) 9.1 (5.4)
+128K 37.9 (30.3) 30.5 (23.1) 35.4 (25.4) 30.7 (23.3)
+192K 81.5 (70.0) 64.6 (53.6) 73.0 (58.3) 65.1 (53.6)
 […]
-32K 5.19 2.59 3.06 2.37
-Llama-3 64K 18.65 6.84 8.94 6.50
--8B 128K 71.16 20.78 28.98 20.29
-192K 157.33 42.31 59.91 41.35
-Table 6. Time-to-first-token (TTFT) for a single prompt of varying context lengths.
-(a) Decode throughput (measured as the number of output tokens generated per second) with varying batch size.
-(b) Mean GPU time spent in computing attention with varying batch sizes.
-Figure 9.
-Decode throughput with varying batch sizes using context length 16K for each request (FA: FlashAttention, bs: block size).
-We evaluate vLLM and FlashAttention with two different block sizes: 16 and 128.
-vLLM performs best with block size 16 because its attention kernel is more efficient with smaller block sizes.
-FlashAttention’s GPU kernel is up to 2.85×2.85$\times$2.85 × faster than the best version of vLLM’s kernel (Yi-6B, 16*16K).
-However, smaller blocks add CPU overhead e.g., FlashAttention with block size 16 is worse than with block size 128.
-vAttention provides similar gains that the best version of FlashAttention provides over vLLM, but without user-level physical memory management and without a PagedAttention kernel.
-Both vAttention configurations therefore use kernels that support chunked-prefills over a virtually contiguous KV-cache.
-We add dynamic memory allocation support to them without having to modify their code.
-Figure 8 shows the prefill throughput of the four configurations for Yi-6B and Llama-3-8B.
-In all cases, vAttention delivers consistently higher throughput, outperforming FA_Paged by 1.603.92×1.60-3.92$\times$1.60 - 3.92 × and FI_Paged by 1.031.45×1.03-1.45$\times$1.03 - 1.45 ×.
-For the same experiment, Table 6 shows the TTFT with different context lengths.
-Since TTFT directly depends on prefill throughput, compared to using vanilla kernels with vAttention, FA_Paged and FI_Paged increase TTFT by up to 3.92×3.92$\times$3.92 × (Yi-6B, context length 192K) and 1.45×1.45$\times$1.45 × (Llama-3-8B, context length 192K), respectively.
-The source of vAttention’s performance gain is twofold in these scenarios.
-First, the vanilla kernel is faster than the paged kernel in both FlashAttention and FlashInfer.
-While FlashAttention’s paged kernel is not optimized for prefills (it is optimized for decodes), FlashInfer’s paged kernel is specifically designed to support chunked-prefills.
-However, the paged kernel is slower than the vanilla kernel as discussed in §3.3.
-This example illustrates the complexities of transferring performance-critical optimizations between different implementations – even when the implementations are written by the same team.
-The second source of improvement is vAttention’s less CPU overhead.
-For example, appending a new K or V tensor to the KV-cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time.
-Further, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration.
-vAttention avoids such overheads because it maintains KV-cache’s virtual contiguity and therefore does not require a Block-Table.
-7.2. Portability and Performance for Decodes
-To evaluate decode performance, we focus on long-context scenarios (16K) because the latency of attention kernel becomes significant only for long contexts [^4] For short contexts, the computation time of the feed-forward-network dominates inference latency (Agrawal et al., 2024).
-We evaluate the following configurations:
-vLLM: We use vLLM v0.2.7 as the primary baseline. vLLM pioneered PagedAttention and uses a custom paged kernel for decodes, derived from FasterTransformer (ftk, [n. d.]).
-FA_Paged: For the second baseline, we integrate the FlashAttention kernel into vLLM’s serving stack.
-This represents a state-of-the-art PagedAttention kernel that includes optimizations such as sequence parallelism and in-place copy of new key and value vectors into the KV-cache.
-We evaluate the paged kernels of vLLM and FlashAttention with two different block sizes – 16 and 128 – to capture the effect of block size on performance.
-FA_vAttention: For vAttention, we integrated the vanilla kernel of FlashAttention into vLLM’s serving stack.
-The kernel works with a virtually contiguous KV-cache to which we dynamically allocate physical memory using 2MB pages.
-9(a) shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch sizes wherein the initial context length of each request is 16K tokens and we generate $256$ tokens for each request.
-We compute decode throughput based on the mean latency of 256 decode iterations.
-We summarize the key takeaways below.
-First, vAttention outperforms vLLM (both block sizes) and FA_Paged (block size 16), while roughly matching the best configuration of FA_Paged (block size 128).
-The maximum improvement over vLLM is 1.97×1.97$\times$1.97 × for Yi-6B, 1.3×1.3$\times$1.3 × for Llama-3-8B and 1.6×1.6$\times$1.6 × for Yi-34B.
-The relative gains over vLLM also increase as the batch size grows.
-For example, the gain increases from about $\times1$ to 1.97×1.97$\times$1.97 × as batch size increases from 1 to 8 for Yi-6B.
-This is because the latency of attention computation grows proportional to the total number of tokens in the batch (see 9(b)) whereas the cost of linear operators remains roughly the same (Agrawal et al., 2023; Patel et al., 2023; Agrawal et al., 2024).
-Therefore, the contribution of attention kernel in the overall latency – and subsequently gain with a more efficient kernel – increases with the batch size.
-While FA_Paged (block size 128) provides similar gains as vAttention, note that FA_Paged requires a new implementation of the GPU kernel whereas vAttention simply leverages the vanilla kernel of FlashAttention.
-Second, 9(b) confirms that performance difference between vLLM and FA_Paged/vAttention is indeed due to the attention kernels.
-In the worst case, the latency of vLLM’s best PagedAttention kernel (block size 16) is up to 2.85×2.85$\times$2.85 × higher for Yi-6B, up to 1.45×1.45$\times$1.45 × for Llama-3-8B and up to 2.62×2.62$\times$2.62 × for Yi-34B than the FlashAttention kernel.
-Finally, throughput can be sensitive to block size even when memory capacity is not a constraint.
-For example, as discussed in §3.3, vLLM’s attention kernel has a significantly higher latency with block size 128 than with block size 16 (also see 9(b)).
-In the worst case, block size 128 degrades vLLM’s throughput by 36%.
-While block size has a smaller impact on FlashAttention, using a small block size can still hurt throughput due to CPU overheads, particularly due to the overhead of creating Block-Tables for every iteration (§3.3).
-For example, FlashAttention with block size 128 delivers 7% higher throughput than block size 16 for Llama-3-8B (531 vs 494 tokens per second with batch size 32).
-Figure 10. Latency of decode iterations with and without overlapping memory allocation with compute (batch size=4,context length=32K). Spikes show the latency impact of synchronous memory allocation.
+-8B 128K 20.4 (15.4) 16.8 (11.6) 20.3 (12.8) 16.8 (11.4)
+192K 43.3 (35.6) 34.8 (26.9) 40.9 (29.7) 34.7 (26.7)
+[…]
+Yi-34B 128K 82.8 (56.9) 68.4 (43.2) 76.0 (46.7) 67.4 (42.5)
+192K 170.7 (131.8) 136.9 (98.8) 148.8 (104.7) 134.6 (96.5)
+Table 5. Prefill completion and attention (in parenthesis) time with different attention back-ends (unit: seconds).
+Figure 7. Decode throughput. FA_vAttention is on par with FA_Paged (note the overlapping lines) which is the best among all PagedAttention based alternatives, while outperforming FI_Paged and vLLM.
+6.1. Prefill Evaluation
+We evaluate four configs for prefill: FA_Paged, FI_Paged, FA_vAttention and FI_vAttention.
+Configurations with the “_Paged” suffix represent the PagedAttention-based kernel and those with “_vAttention” use the non-paged kernels of the respective library.
+Figure 6 shows the prefill throughput for Yi-6B, Llama-3-8B and Yi-34B.
+We summarize our key findings below.
+Small contexts: For small contexts, prefill cost is dominated by the linear operators i.e., attention’s contribution is relatively low (sarathi2023, ).
+Hence, even though vAttention helps speed up attention computation, the throughput of both paged and vAttention back-ends is nearly identical in FlashAttention.
+However, for FlashInfer, we find that using vAttention helps improve prefill throughput even for small contexts.
+This is because FlashInfer incurs various other sources of overhead in the paged version.
+First, appending a new K or V tensor to the KV-cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time (the copy operation has been optimized for FlashAttention by vLLM (vllm-copy-kernels, )).
+Second, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration.
+vAttention avoids such overheads because it maintains KV-cache’s virtual contiguity, eliminating the need for a Block-Table.
+Long contexts: The contribution of attention computation becomes significant at 16K and higher context lengths in our experiments.
+Therefore, even for FlashAttention back-end, vAttention outperforms the paged counterpart.
+For example, at context length 192K, FA_vAttention outperforms FA_Paged by 1.24×1.24$\times$1.24 ×, 1.26×1.26$\times$1.26 × and 1.24×1.24$\times$1.24 × for Yi-6B, Llama-3-8B and Yi-34B, respectively.
+Similarly, for FlashInfer back-ends, FI_vAttention improves prefill throughput by up to 1.25×1.25$\times$1.25 × and 1.36×1.36$\times$1.36 × for Yi-6B and Llama-3-8B (context length 16K), and 1.17×1.17$\times$1.17 × for Yi-34B (context length 32K).
+Attention time: vAttention’s improvement in prefill throughput is primarily due to faster attention kernels enabled by a (virtually) contiguous KV-cache.
+This is because the prefill phase of a long prompt is primarily dominated by attention computation, as can be observed by comparing the numbers inside parenthesis with the total prefill completion time in Table 5.
+For FlashAttention, nearly all the gains of vAttention are due to faster attention kernels e.g., prefill gains of 1.5 seconds, 7.4 seconds and 16.9 seconds for Yi-6B are almost entirely due to gains in attention computation.
+vAttention enabled kernels also help with the FlashInfer back-end.
+In addition, FI_Paged also has other sources of overheads e.g., in the 14 seconds of total savings (Yi-34B, 192K context), only 7 seconds is due to attention and the rest is due to other sources.
+Figure 8. Makespan of serving a set of 50 requests with varying context lengths and the ratio of prefill to decode tokens.
+6.2. Decode Evaluation
+For decodes, in addition to FA_Paged and FI_Paged, we also evaluate the throughput obtained with vLLM’s decode kernel (the first ever kernel to support PagedAttention).
+For vAttention, we use FlashAttention’s non-paged kernel.
+Unfortunately FlashInfer’s non-paged decode kernel has significantly higher latency compared to all the other kernels we used in these experiments (up to 14.6×14.6$\times$14.6 ×, Table 10).
+Hence, while vAttention supports dynamic memory allocation for FlashInfer’s non-paged decode kernel, we omit it for evaluation in this section (note that vAttention is responsible only for allocating memory – it cannot make a slow kernel fast).
+Figure 7 shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch size up to 32 (except for Yi-34B which runs out of memory for batch size 32).
+We set the initial context length of each request to 16K tokens and calculate decode throughput based on the mean latency of 400 decode iterations (see Figure 7).
+First, vAttention is on par with the best of PagedAttention as shown by FA_Paged and FA_vAttention in Figure 7.
+In comparison, FI_Paged has somewhat lower throughput and vLLM is the worst for all models and configurations.
+For example, FA_Paged and FA_vAttention outperform vLLM by up to 1.99×1.$9$, 1.58×1.58$\times$1.58 × and 1.53×1.53$\times$1.53 × for Yi-6B, Llama-3-8B and Yi-34B, respectively.
+The primary reason is that vLLM’s decode kernel has significantly higher latency than the FlashAttention based kernels; while FlashAttention has continuously adopted new optimizations (e.g., FlashDecoding (flashdecoding, )), vLLM has lagged behind.
+For example, Table 6 shows that vLLM’s PagedAttention kernel incurs up to 2.8×2.8$\times$2.8 ×, 1.5×1.5$\times$1.5 ×, and 2.5×2.5$\times$2.5 × higher latency than FlashAttention kernels for Yi-6B, Llama-3-8B and Yi-34B.
+This is despite vLLM being in an actively maintained open-source serving stack, as well as being used by various companies for serving LLMs.
+This is an important result that underlines the importance of adopting new optimizations.
+Model BS vLLM FA_Paged FI_Paged FA_vAttention
+Yi-6B 16 32.3 11.5 15.2 11.3
+32 64.1 25.5 25.4 25.3
+[…]
+32 35.3 25.4 23.23 25.3
+[…]
+16 55.1 21.7 28.8 21.8
+Table 6. Total latency of attention kernel (sum of all layers) per decode iteration (in milliseconds, BS = batch size).
+Second, relative gains of FA_vAttention and FA_Paged increase over vLLM with the batch size e.g., as batch size increases from 4 to 32 for Llama-3-8B, relative gains increase from 1.05×1.05$\times$1.05 × to 1.58×1.58$\times$1.58 ×.
+This is because the latency of a decode attention kernel is proportional to the total number of tokens in the batch (vidur, ).
+Therefore, the contribution of attention kernel in the overall latency – and consequently gains with a more efficient kernel – increase with the batch size.
+Further, for the same reasons as discussed in §6.1 (faster attention and lower CPU overhead), FA_vAttention delivers up to 1.23×1.23$\times$1.23 × higher throughput than FI_Paged (Yi-6B, batch size 12).
+Finally, note that vAttention is only as good as the state-of-the-art PagedAttention for decode throughput, as compared to prefills where it outperforms PagedAttention.
+We believe this is due to memory bound nature of decode attention i.e., memory stalls make it possible to hide the effect of additional compute that paging support requires.
+This is evident from Table 6 wherein FA_Paged and FA_vAttention have similar latency.
+However, hiding compute overhead in a prefill attention kernel is hard because it is already compute bound (see Table 5 for latency gains in the prefill attention kernel).
+6.3. End-to-end Throughput Evaluation
+For end-to-end performance, we measure the makespan of serving a static trace of 50 requests for long context scenarios.
+To understand system performance for different workloads, we vary the initial context length of requests from 32K to 128K, and the ratio of prefill to decode tokens (referred as P:D) from 500 to 50.
+Note that a higher P:D as well as a longer context indicates that the workload is more prefill bound; we expect vAttention to improve performance in such cases.
+Figure 8 shows the makespan for all the three models.
+For Yi-6B, FA_vAttention is up to 1.22×1.$2$, 1.16×1.16$\times$1.16 × and 1.13×1.13$\times$1.13 × faster than FA_Paged for P:D ratio of 500, 100 and 50.
+FA_vAttention is also up to 1.19×1.19$\times$1.19 × faster for Llama-3-8B and up to 1.16×1.16$\times$1.16 × faster for Yi-34B, compared to FA_Paged.
+Our gains over FA_Paged are higher at higher P:D ratios, and for a fixed P:D, our gains increase as the context length grows.
+These results are consistent with our expectation since a higher P:D and context lengths both mean that the system spends more time processing prefills where use of non-paged attention kernels outperform paged kernels.
+Interestingly, we observe that despite being faster than FA_Paged in prefills, FI_Paged under-performs FA_Paged in end-to-end throughput in most cases (due to inefficient decodes as shown in Figure 7).
+Hence, our gains over FI_Paged are relatively higher (up to 1.29×1.29$\times$1.29 ×), except for some cases where FI_Paged outperforms FA_Paged (e.g., Yi-6B and Yi-34B, context length=128K, P:D=500).
+Figure 9. Latency of decode iterations with and without overlapping memory allocation with compute (batch size=32, context length=4K–8K, model: Llama-3-8B)
+6.4. Ablation Studies
+6.4.1. Hiding allocation latency
+Figure 9 shows the latency of more than 2500 decode iterations of Llama-3-8B.
+For this experiment, we used a batch of 32 requests and initialized the prefill context length of each request to be in between 4K–8K (chosen randomly).
+It is evident that overlapping memory allocation with compute effectively hides the latency of allocating memory via CUDA APIs.
+We used 2MB pages for this experiment to show that even the worst case latency can be hidden (Table 2 shows that 2MB pages incur highest latency).
+In contrast, allocating memory synchronously via CUDA APIs leads to frequent latency spikes of 5ms–20ms, depending on how many requests need new physical memory pages.
+6.4.2. Deferred reclamation
+Figure 10 shows that synchronous memory allocation using CUDA APIs for prefills incurs overhead of up to 1.15×1.15$\times$1.15 × with 64KB pages and up to 1.03×1.03$\times$1.03 × with 2MB pages.
+In most cases, deferred reclamation eliminates the need to invoke CUDA APIs for prefills because a newly arrived request can simply re-use physical memory pages that were allocated to a prior request.
+This way, deferred reclamation ensures that prefill latency is not affected by memory allocation.
+6.4.3. Memory allocation bandwidth
+Table 7 shows that even with 64KB (our smallest) page size, vAttention can allocate as much as 7.6GB per second per GPU.
+This is more than an order of magnitude higher than the maximum memory allocation rate of 600MB per second of decodes (Figure 4).
+Larger page sizes and higher TP dimensions increase the memory allocation rate proportionally.
+Therefore, memory allocation bandwidth of CUDA APIs is more than enough for LLM inference.
+Figure 10. Prefill completion time of a single prompt of 16K tokens with different memory allocation strategies.
 Config. 64KB 128KB 256KB 2MB
@@
 TP-2 15.18 29.12 54.08 70.34
-Table 7. Physical memory allocation bandwidth (GB per second) for vAttention with different page sizes.
-7.3. Efficacy of Physical Memory Allocation
-The PyTorch caching allocator allocates memory objects (e.g., tensors) without requiring a round-trip to the kernel.
-In contrast, vAttention needs to invoke CUDA’s kernel driver while mapping a new physical page in a request’s KV-cache.
-In this section, we show that with our optimizations, vAttention can effectively meet the requirements of both the prefill and decodes phases in an LLM serving system.
-Figure 11.
-Time to compute the prefill phase of a single prompt of 16K tokens.
-vAttention is as good as vLLM since it allocates memory out of the critical path.
-For illustration purpose, results with different page sizes show the overhead of synchronously allocating physical memory before a request’s prefill phase begins.
+Table 7. Physical memory allocation bandwidth (GB per second) with varying allocation granularity.
+6.4.4. Effect of page size.
+In many applications, use of smaller pages can potentially degrade performance due to TLB thrashing (neighbourhoodawareat, ; ingens, ; hawkeye, ; architecturalsupportforgpuat, ).
+We find that this is not the case for LLM inference.
+For example, Figure 11 shows that the execution latency of an attention kernel remains largely unaffected when the page size for KV-cache is reduced to 64KB, from the original 2MB.
+In a separate experiment, we find that these results are also consistent with very large models e.g., Llama-3-70B and GPT-3-175B.
+We attribute this to the regular computation pattern of the attention operator as well as the hand-tuned implementations that explicitly try to avoid irregular memory accesses.
+Further, we find that 2MB pages are good enough for many online serving scenarios where latency constraints limit the maximum batch size.
+However, for throughput-oriented scenarios, smaller pages may be preferred.
+For example, using 2MB pages could serve batch sizes of up to 187, 203 and 56 for Yi-6B, Llama-3-8B and Yi-34B on a dynamic trace (dataset: OpenChat (wang2023openchat, ), load: 7 queries per seconds).
+In contrast, 64KB pages helped serve batch sizes of up to 240, 258 and 68 for our models (see Figure 12).
+Figure 11. Effect of page size on the performance of FlashAttention’s prefill (left) and decode (right) attention kernels (model: Llama-3-8B).
+Figure 12. Maximum batch size obtained with different page sizes for a dynamic workload. Large pages limit batch size due to internal fragmentation.
+Figure 13. Illustration of code changes needed to replace the prefill attention kernel of FlashAttention by FlashInfer when using vAttention for dynamic physical memory allocation.
 Model # Tokens in a physical memory block Max memory waste per request
@@
 Yi-6B (TP-2) 128 256 512 4096 8MB 16MB 32MB 256MB
-Llama-3-8B (TP-1) 32 64 128 1024 4MB 8MB 16MB 128MB
+[…]
 Llama-3-8B (TP-2) 64 128 256 2048 8MB 16MB 32MB 256MB
-Yi-34B (TP-1) 32 64 128 1024 7.5MB 15MB 30MB 240MB
+[…]
 Yi-34B (TP-2) 64 128 256 2048 15MB 30MB 60MB 480MB
 Table 8.
-Block size (number of tokens in a newly allocated physical memory page) as a function of the page size and TP configuration.
-Columns on the right show how much memory can be wasted per-request in the worst-case.
-Table 7 shows that even with our smallest page size of 64KB, vAttention can allocate as much as 7.6GB per second per GPU.
-This is more than an order of magnitude higher than the maximum memory allocation rate of 600MB per second of decodes (Figure 6).
-Larger page sizes and higher TP dimensions increase the memory allocation rate proportionally.
-This shows that vAttention is more than capable of satisfying the memory allocation bandwidth of decodes.
-Further, Figure 10 shows that our optimization of overlapping memory allocation with model execution also hides the latency impact of invoking CUDA APIs.
-This example shows the latency of consecutive decode iterations for Llama-3-8B running with TP-1 and batch size 4.
-Without overlapping memory allocation with compute, new pages for the KV-cache are allocated synchronously which increase the latency of some iterations from 25 milliseconds to 41 milliseconds ($\approx$ 4 millisecond latency due to memory allocation to a single request).
-Note that these latency spikes occur after every 1024 iterations because we used 2MB page size for these experiments and each 2MB page contains 1024 tokens for this model configuration.
-When memory allocation is overlapped with the model execution of previous decode iteration, the latency effect is completely hidden.
-Finally, since a prefill may require more than one page for the KV-cache, we also investigate how different memory allocation schemes impact the latency of a prefill.
-Figure 11 shows that allocating physical memory synchronously on-demand (when our background thread, deferred reclamation and eager allocation optimizations are all disabled) can add as much as 15% overhead with 64KB page size.
-Larger page sizes amortize the cost of allocation and reduce the overhead to as low as 3% (with 256KB and 2MB page sizes).
-vAttention further reduces the cost of allocation with deferred reclamation and eager allocation while overlapping memory allocation with compute.
-In most cases, these optimizations ensure that a newly arrived request can simply re-use the physical memory pages that were allocated to a previous request.
-Therefore, vAttention incurs negligible overhead, performing as good as vLLM for prefills.
-7.4. Analysis of Memory Fragmentation
-Table 8 shows the block size (defined as the minimum number of tokens in a page), and how much physical memory can be (theoretically) wasted due to over allocation in the worst-case.
+Block size (number of tokens in a physical memory page) as a function of the page size and degree of tensor parallelism.
+Columns on the right show how much physical memory can be wasted per-request in the worst-case.
+6.5. Programming Effort
+vAttention makes it feasible to replace one attention kernel with another with only a few lines of code changes in the serving framework (Figure 13).
+In contrast, in PagedAttention, a developer first needs to write a paged kernel and then make significant changes in the serving framework.
+For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes in over 15 files (vllmfi1, ; vllmfi2, ; vllmfi3, ).
+Implementing the initial paging support in FlashAttention kernel also required $\approx 280$ lines of code changes (fapaging, ) and additional efforts to enable support for smaller block sizes (fasmall, ).
+7. Discussion
+Batch Size FA_Paged (Block Size = 256) FA_Paged (Block Size = 16)
+Model Ratio
+Yi-6B 1 0.085 0.085 1.00×1.$0$
+2 0.093 0.094 1.01×1.01$\times$1.01 ×
+4 0.109 0.115 1.05×1.05$\times$1.05 ×
+[…]
+2 0.097 0.105 1.07×1.07$\times$1.07 ×
+4 0.136 0.143 1.05×1.05$\times$1.05 ×
+[…]
+2 0.088 0.095 1.08×1.08$\times$1.08 ×
+4 0.116 0.117 1.01×1.01$\times$1.01 ×
+Table 9. Impact of varying block size on the execution latency (in milliseconds) of FlashAttention decode kernel.
+7.1. Further Analysis of Internal Fragmentation
+vAttention uses page size as the unit of physical memory allocation as opposed to the PagedAttention’s unit of KV-cache block size.
+Therefore, it is natural to ask what is the relationship between page size and block size.
+To compute this relationship, let $S$ be the size of a single token’s per-layer K-cache (or V-cache) on a worker (see §5.1).
+Then, page size t translates to KV-cache block size $=t/S$.
+Note that $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16).
+Therefore, block size = $t/(H\times D\times P)$.
+Table 8 shows the block size for different model configurations and physical memory allocation sizes; we show each model under two TP configurations – TP-1 and TP-2 – to highlight the effect of TP dimension on block size.
+In addition, the table also shows how much physical memory can be (theoretically) wasted by vAttention due to internal fragmentation in the worst-case.
 The worst-case occurs when a new page is allocated but remains completely unused.
-Further, we show each model under two TP configurations – TP-1 and TP-2 – to highlight the effect of TP dimension on block size.
+Context Length Batch Size FI_NonPaged FI_Paged Ratio
+8K 4 0.902 0.071 12.7×12.7$\times$12.7 ×
+8K 8 0.931 0.128 7.3×7.3$\times$7.3 ×
+8K 16 0.943 0.229 4.2×4.2$\times$4.2 ×
+[…]
+16K 8 1.821 0.233 7.8×7.8$\times$7.8 ×
+16K 16 1.875 0.450 4.2×4.2$\times$4.2 ×
+[…]
+32K 8 3.291 0.458 7.2×7.2$\times$7.2 ×
+32K 16 3.365 0.891 3.8×3.8$\times$3.8 ×
+Table 10.
+Latency (in milliseconds) of paged and non-paged decode attention kernels in FlashInfer.
+Ratio denotes the latency of non-paged kernel divided by the latency of paged kernel (model: Llama-3-8B, 2 A100 GPUs).
 vAttention allocates physical memory equivalent to the page size on each TP worker whereas the per-token physical memory requirement of a worker goes down as TP dimension increases (because KV heads get split across TP workers).
 Therefore, block size increases proportionally with the TP dimension.
-Table 8 shows that this results in the smallest block sizes of 32 (Yi-34B TP-1) to 128 (Yi-6B TP-2).
-In terms of the amount of physical memory, 64KB page size results in a maximum theoretical waste of only 4-15MB per request which increases to 16-60MB for 256KB page size.
-Overall, the important point to note is that by controlling the granularity of physical memory allocation, vAttention makes memory fragmentation insignificant.
-Recall that serving throughput saturates at about 200 batch size for all of our models (Figure 6).
-Hence, even at such large batch sizes, the maximum theoretical memory waste is at most a few GBs.
-Therefore, similar to vLLM, vAttention is highly effective in reducing fragmentation and allows serving using large batch sizes.
-However, if required, page size can be reduced further to as low as 4KB which is the minimum page size supported in almost all architectures today, including NVIDIA GPUs (B et al., 2023).
-Figure 12.
-Illustration of code changes required to replace the prefill attention kernel of FlashAttention with that of FlashInfer.
-vAttention supports dynamic memory allocation transparently, enabling easy switching between kernels.
-Implementation Effort
-The primary advantage of vAttention is portability: it enables one to seamlessly integrate new attention kernels without having to having to write a paged version of it or change the serving framework.
-For example, switching between the prefill or decode kernels of FlashAttention and FlashInfer requires only a few lines of code changes as shown in Figure 12.
-In contrast, in PagedAttention, the developers first need to write a paged attention kernel and then make significant changes in the serving framework.
-For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes spread over 15 files (vll, 2024b, d, c).
-Implementing the initial paging support in FlashAttention GPU kernel also required about 280 lines of code changes (fap, 2024) and additional efforts to enable support for smaller block sizes (fas, 2024a).
-Given the rapid pace of innovation in LLMs, we believe it is important to reduce programming burden: production systems should be able to leverage new optimizations of the attention operator without re-writing code – similar to how optimized implementations of GEMMs are leveraged by deep learning frameworks without programmer intervention.
+Table 8 shows that using 2MB pages leads to large KV-cache block sizes of 1024 (Llama-3-8B and Yi-34B, TP-1) to 4096 (Yi-6B, TP-2) which could waste significant amount of physical memory (100s of MBs, per-request).
+Use of smaller (e.g., 64KB) pages reduces internal fragmentation proportionally, resulting in KV-cache block size of 32 (Yi-34B and Llama-3-8B, TP-1) to 128 (Yi-6B, TP-2), which in turn limit the maximum theoretical waste of only 4-15MB physical memory, per request.
+This shows that controlling the granularity of physical memory allocation is essential for reducing fragmentation.
+If required, page size can be reduced further to as low as 4KB which is the minimum page size supported in almost all architectures today, including NVIDIA GPUs.
+7.2. Effect of Block Size on FlashAttention Kernel
+We find that varying the block size of KV-cache also affects the execution latency of FlashAttention’s decode kernel (in addition to vLLM wherein using a larger block size increase the latency of decode kernel by up to 3×3$\times$3 × as shown in Figure 3).
+For example, Table 9 shows that using a smaller block size of 16 increases the latency of FlashAttention decode kernel by up to $5\%$ for Yi-6B, up to $8\%$ for Llama-3-8B and up to $9\%$ for Yi-34B.
+This is likely because work distribution and shared memory usage in a paged kernel needs to explicitly account for block size.
+Hence, writing an implementation that is equally efficient for all block sizes could be non-trivial.
+In fact, it is interesting to see that vLLM decode kernel performs better with small block sizes whereas FlashAttention decode kernel performs better with large block sizes.
+While the impact of block size on kernel latency is much less pronounced in FlashAttention compared to vLLM, a degradation in performance is still undesirable for an operator as important as attention.
+vAttention is free of such implementation challenges because it does not require a block table.
+7.3. FlashInfer Non-paged Decode Attention Kernel
+Note that we did not include FlashInfer non-paged decode kernel in our evaluation (§6).
+There are two reasons for this.
+First, the non-paged kernel has a much higher latency than the paged counterpart of FlashInfer e.g., Table 10 shows that the latency of non-paged kernel is up to 14.6×14.6$\times$14.6 × higher than the paged kernel.
+Further, the latency of the paged kernel follows an expected trend i.e., increasing the context length or batch size leads to a proportional increase in the execution latency.
+In contrast, increasing the batch size does not affect the latency of the non-paged kernel (e.g., context lengths of 8K and 32K).
+This behavior of the non-paged kernel is counter-intuitive and likely due to un-optimized implementation.
+Note that the job of a memory allocator is to simply allocate memory – it cannot accelerate a kernel.
+Second, the non-paged decode kernel of FlashInfer does not support variable sequence lengths – a fundamental requirement for LLM inference.
 8. Related Work
-In a recent work, GMLake (Guo et al., 2024) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size.
+Optimizing LLM inference is an active area of research.
+Various techniques have been proposed to improve different aspects of LLM serving like batching (orca, ; vllmsosp, ; sarathiserve2024, ), disaggregation (patel2023splitwise, ; distserve2024, ; tetriinfer, ), scheduling (flexgen, ; fastserve, ).
+However, central to all these techniques is the need for efficient KV-cache memory management.
+Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trtllmgithub, ), LightLLM (lightllm:github, ), and libraries e.g., FlashAttention (fagithub, ) and FlashInfer (figithub, ).
+In contrast, vAttention offers an alternate – which we believe is also a more principled – approach to dynamic KV-cache memory management.
+In a recent work, GMLake (gmlakeasplos24, ) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size.
 In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors for large object allocations.
@@
 We proposed various LLM inference specific optimizations to meet these requirements.
-Optimizing LLM inference is an active area of research.
-Various scheduling systems have been proposed to improve different aspects of LLM serving.
-For example, Orca (Yu et al., 2022) and vLLM (Kwon et al., 2023) are aimed at improving serving throughput with efficient batching.
-Sarathi (Agrawal et al., 2023) and SplitFuse (Holmes et al., 2024) split a long prefill into multiple smaller chunks and combine decode tokens with each chunk to improve GPU compute utilization.
-Based on similar techniques, Sarathi-Serve (Agrawal et al., 2024) proposes stall-free batching to minimize the impact of long-running prefill iterations on decode latency.
-Splitwise (Patel et al., 2023), DistServe (Zhong et al., 2024) and TetriInfer (Hu et al., 2024) disaggregate the prefill and decode phases, executing them on different replicas so as to avoid interference between the prefill and decode requests.
-For offline inference on resource-constrained devices, FlexGen (Sheng et al., 2023) proposed a scheduling and offloading strategy to improve throughput.
-FastServe (Wu et al., 2023) minimizes job completion times in LLM inference using preemptive scheduling.
-For all the above systems to work effectively, efficient use of GPU physical memory is essential.
-Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trt, 2023), LightLLM (lig, 2023) and kernel implementations e.g., in FlashAttention (fag, 2022) and FlashInfer (fig, 2023).
-In contrast, vAttention offers an alternate approach for dynamic KV-cache memory management.
-We show that using system support for demand paging can easily add dynamic memory management support to existing kernel implementations.
+Similar to our approach, PyTorch also recently added support for expandable segments (expandable-segments, ) that dynamically attach physical memory pages to a pre-reserved virtual memory buffer, if enabled.
+However, PyTorch uses expandable segments opportunistically, using synchronous memory allocation, and allocates physical memory pages only in multiples of 2MB granularity.
+In contrast, vAttention caters to the specific requirements of the KV-cache.
+Some other concurrent works are also motivated to optimize attention kernels (flashattnhopper, ; mirage-tensor-compiler, ; thunderkittens, ; heavyhitters, ; flashdecoding++, ; leanattention, ; flash-attention-3, ).
+As one would expect, new research ideas do not start with paging support in mind.
+vAttention would make it easier to deploy them.
 9. Conclusion
-PagedAttention has emerged as the de facto standard for dynamic memory allocation in LLM inference.
-PagedAttention eliminates the need to reserve GPU memory ahead-of-time and therefore boosts serving throughput by fitting a larger batch size in GPU memory.
-While PagedAttention effectively tackles memory fragmentation, we argue that its approach of storing KV-cache in non-contiguous virtual memory introduces software complexity as well as portability and efficiency challenges.
-Instead, we show that using low-level system support for demand paging can avoid the pitfalls of PagedAttention.
-Our proposed system vAttention adds support for dynamic physical memory allocation to existing attention kernels, eliminating the need to re-write GPU code or write a memory manager in the serving framework.
-We show that vAttention reduces software complexity while improving portability and performance.
+In this paper, we propose vAttention for dynamic memory management in LLM serving systems.
+The key highlight of vAttention is that it leverages system support for demand paging instead of implementing it in user space.
+We present various examples to show that the vAttention approach reduces programming burden while improving portability and performance compared to the popular PagedAttention approach that many LLM serving systems use today.

### v3

@@
 Abstract.
-Efficient management of GPU memory is essential for high throughput LLM inference.
-Prior systems used to reserve KV-cache memory ahead-of-time that resulted in wasted capacity due to internal fragmentation.
-Inspired by demand paging, vLLM proposed PagedAttention to enable dynamic memory allocation for KV-cache.
-This approach eliminates fragmentation and improves serving throughout.
-However, to be able to allocate physical memory dynamically, PagedAttention changes the layout of KV-cache from contiguous virtual memory to non-contiguous virtual memory.
-As a consequence, one needs to rewrite the attention kernels to support paging, and implement a memory manager in the serving framework.
-This results in both performance and programming overheads, as well as portability challenges in adopting state-of-the-art attention kernels.
-In this paper, we propose vAttention, a new approach for dynamic KV-cache memory management.
-In contrast to PagedAttention, vAttention stores KV-cache in contiguous virtual memory and leverages OS support for on-demand allocation of physical memory.
-vAttention thus enables one to use state-of-the art attention kernels out-of-the-box by adding support for dynamic allocation of physical memory without having to re-write their code.
-We implement vAttention in the vLLM serving stack to show that it also helps improve decode throughput by up to 1.99×1.$9$ over vLLM, and the end-to-end serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
+PagedAttention is a popular approach for dynamic memory allocation in LLM serving systems.
+It enables on-demand allocation of GPU memory to mitigate KV cache fragmentation — a phenomenon that crippled the batch size (and consequently throughput) in prior systems.
+However, in trying to allocate physical memory at runtime, PagedAttention ends up changing the virtual memory layout of the KV cache from contiguous to non-contiguous.
+Such a design leads to non-trivial programming and performance overheads.
+We present vAttention — an approach that mitigates fragmentation in physical memory while retaining the virtual memory contiguity of the KV cache.
+We achieve this by decoupling the allocation of virtual and physical memory using CUDA virtual memory management APIs.
+We also introduce various LLM-specific optimizations to address the limitations of CUDA virtual memory support.
+Overall, vAttention is a simpler, portable, and performant alternative to PagedAttention: it supports various attention kernels out-of-the-box and improves LLM serving throughput by up to 1.23×1.23$\times$1.23 × compared to the use of PagedAttention-based kernels of FlashAttention-2 and FlashInfer.
+Large language models; KV cache; fragmentation; memory management
+journalyear: 2025copyright: rightsretainedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3669940.3707256isbn: 979-8-4007-0698-1/25/03ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Virtual memoryccs: Software and its engineering Software design techniquesccs: General and reference Performanceccs: Computing methodologies Neural networks
 1. Introduction
-Large Language Models (LLMs) are deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (openai2022gpt4techreport, ; chowdhery2022Palm, ; bingai, ; bard, ; githubcopilot, ; amazoncodewhisperer, ; replitghostwriter, ).
-Given the size and scale of modern LLM deployments, optimizing inference has become extremely important  (sarathi2023, ; splitfuse2024, ; patel2023splitwise, ; distserve2024, ; tetriinfer, ; vllmsosp, ; orca, ).
-LLM inference happens in two phases: a compute-bound prefill phase, followed by several iterations of memory-bound decode phase.
-Batching is a powerful technique to boost LLM serving throughput (orca, ; vllmsosp, ; patel2023splitwise, ; sarathiserve2024, ), as it amortizes the cost of fetching model weights from HBM to the on-device SRAM during decodes.
 System/library and issues related to PagedAttention
-vLLM: Pioneered PagedAttention.
-Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.8×2.8$\times$2.8 × slower than FlashAttention (Table 6).
-Further, changing block size changes the execute latency of the kernel by as much as 3×3$\times$3 × (Figure 3).
-FlashAttention: PagedAttention based prefill kernel is up to $28\%$ slower than the non-paged kernel (Figure 2) while the decode kernel is up to $12\%$ slower.
-Initial attempts to add paging support failed unit tests (fa-paged-crash, ).
-TensorRT-LLM: Serving throughput dropped by more than $10\%$ in a Python front-end (tensorrt-perf-decay, ).
+vLLM (Kwon et al., 2023): Pioneered PagedAttention.
+Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.8×2.8$\times$2.8 × slower than FlashAttention-2 (Table 7).
+Furthermore, changing block size changes the execution time of the kernel by as much as 1.9×1.9$\times$1.9 × (Figure 3).
+FlashAttention-2 (Dao, 2023): PagedAttention-based prefill kernel is up to $37\%$ slower than the non-paged kernel (Figure 2) while the decode kernel is up to $12\%$ slower.
+Initial attempts to add paging support failed unit tests (fa-, 2023).
+FlashAttention-3 (Shah et al., 2024) / SDPA in cuDNN-9 (Nicely and NVIDIA, 2024): State-of-the-art attention kernels for NVIDIA Hopper architecture did not support PagedAttention when released.
+TensorRT-LLM (trt, 2023): Serving throughput dropped by more than $10\%$ in a Python front-end (ten, 2023).
 Recommends using the C++ front-end.
 Even with C++, we observe up to $5\%$ higher latency in some cases with PagedAttention.
-FlashInfer: PagedAttention based prefill kernel is up to $24\%$ slower than the non-paged kernel (Figure 2).
+FlashInfer (Ye et al., 2024): PagedAttention-based prefill kernel is up to $42\%$ slower than the non-paged kernel (Figure 2).
 Table 1.
@@
 These examples highlight the complexity, performance and maintenance challenges associated with this approach.
-Careful allocation of GPU memory is key to enabling larger batch sizes.
-For each request, the serving framework stores the activations of all the tokens processed so far in GPU memory and reuses them for the subsequent token generation.
-This is called the KV-cache (orca, ; sarathi2023, ; patel2023splitwise, ) which accounts for a majority of GPU memory usage during inference.
-Efficiently allocating GPU memory for the KV-cache of a batch of requests is challenging for two reasons.
-First, the per-request KV-cache grows slowly (one token per iteration), and second, a request’s decode length (or its total KV-cache size) is not known ahead of time.
-Prior systems like Orca (orca, ) and FasterTransformer (fastertransformer, ) allocate GPU memory for each request based on the maximum context length supported by the model (e.g., Yi-34B model supports context length of up to 200K (yi-34b-200k-hf, )).
-However, in practice, the number of decode tokens generated are far less; the average decode length for the chat-based sharegpt dataset is 415 tokens (sarathiserve2024, ).
-Therefore, static memory allocation results in severe internal fragmentation, limiting batch size and serving throughput.
-Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (vllmsosp, ) to mitigate KV-cache related memory fragmentation.
-Instead of reserving the maximum sequence length of KV-cache memory ahead-of-time, vLLM allocates small blocks of GPU memory on demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens.
-This approach provides a near-perfect solution for mitigating fragmentation and hence, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems e.g., in TensorRT-LLM (trtllmgithub, ), HuggingFace TGI (hftgi, ), LightLLM (lightllm:github, ) etc.
-However, the PagedAttention approach faces a fundamental consequence of dynamic memory allocation: dynamically allocated memory blocks are not guaranteed to be contiguous.
+Large Language Models (LLMs) are being deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (OpenAI, 2023; Chowdhery et al., 2022; bin, 2024; bar, 2024; git, 2024; ama, 2024; rep, 2024).
+Given the size and scale of modern LLM deployments, optimizing inference has become extremely important  (Agrawal et al., 2023; Holmes et al., 2024; Patel et al., 2024; Zhong et al., 2024; Hu et al., 2024; Kwon et al., 2023; Yu et al., 2022; Kamath et al., 2024).
+Batching is a powerful technique to boost LLM serving throughput (Yu et al., 2022; Kwon et al., 2023; Patel et al., 2024; Agrawal et al., 2024b).
+However, achieving a large batch size requires careful allocation of GPU memory.
+For each request, the serving framework stores the activations of all the tokens processed so far in GPU memory and reuses them for generating subsequent tokens.
+This is called the KV cache (Yu et al., 2022; Agrawal et al., 2023; Patel et al., 2024) which accounts for a majority of GPU memory usage during inference.
+Efficiently allocating GPU memory for the KV cache is challenging for two reasons.
+First, the per-request KV cache grows slowly (one token per iteration), and second, a request’s decode length (or its total KV cache size) is not known ahead of time.
+Prior systems like Orca (Yu et al., 2022) and FasterTransformer (fas, 2024b) allocate memory for each request based on the maximum context length supported by the model (e.g., Yi-34B model supports context length of up to 200K (yi-, 2024a)).
+However, the number of decode tokens generated are far less in practice, e.g., the average decode length for the chat-based sharegpt dataset is 415 tokens (Agrawal et al., 2024b).
+Therefore, static memory allocation could create severe internal fragmentation, limiting batch size and serving throughput.
+Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (Kwon et al., 2023) that allocates small blocks of GPU memory on demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens.
+This approach provides a near-perfect solution for mitigating fragmentation and hence, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems, e.g., TensorRT-LLM, HuggingFace TGI, LightLLM (trt, 2023; hft, 2024; lig, 2023) etc.
+However, we show that PagedAttention faces a fundamental consequence of dynamic memory allocation: dynamically allocated objects are not guaranteed to be contiguous.
 Note that user-level objects are allocated in virtual memory.
-Therefore, in trying to enable dynamic allocation of physical memory, PagedAttention ends up changing the layout of KV-cache from contiguous to non-contiguous virtual memory.
-We argue that this approach has several pitfalls (see Table 1 for empirical evidence and real-world experiences):
-1.
-Requires re-writing the attention kernel (GPU code).
-The elements of a virtually contiguous object can be accessed using index-based lookup which is both simple and efficient.
-Many implementations of deep learning operators also expect virtual contiguity of input tensors.
-However, by storing KV-cache in non-contiguous virtual memory, PagedAttention mandates re-writing GPU code so that the attention kernel can de-reference all the elements of KV-cache.
-The need to re-write code is a major barrier to using new attention optimizations in production settings.
-2.
-Adds software complexity and redundancy (CPU code).
-To stitch together dynamically allocated virtual memory blocks, PagedAttention forces developers to implement a memory manager in the serving framework.
-Typically, this requires a Block-Table manager in user space which is essentially a re-implementation of demand paging (an OS functionality) in user code.
-3.
-Introduces performance overhead.
-PagedAttention adds runtime overhead in the critical path of execution in two ways.
-First, GPU kernels that compute attention over a non-contiguous KV-cache need to execute additional instructions.
-In many cases, this slows down an attention kernel by more than 20%, compared to when the KV-cache is virtually contiguous.
-Second, the user space memory manager can add CPU overhead contributing up to another 10% cost (§3.3).
-In this paper, we instead advocate for leveraging the OS support of virtual memory and demand paging for dynamic KV-cache memory management.
-This approach alleviates the aforementioned issues that appear with PagedAttention.
-To support our claims, we present the design and implementation of vAttention.
-Leveraging OS support, vAttention stores KV-cache in contiguous virtual memory without committing physical memory ahead-of-time, thereby enabling the use of state-of-the art high-performant attention kernels out-of-the box.
-We achieve this by using CUDA support of low-level virtual memory APIs which expose distinct interfaces for allocating virtual and physical memory (§5).
-Leveraging OS support for memory allocation in an LLM serving system poses two key efficiency challenges (§5.4).
-First, the minimum physical memory allocation granularity supported by CUDA is 2MB.
-This can result in significant wasted capacity and fragmentation.
-We address this challenge by modifying the open-source CUDA unified virtual memory driver, adding support for finer-grained physical memory allocations in multiples of 64KB.
-Our evaluation shows that using 64KB pages as the granularity of virtual-to-physical address translation does not affect the execution latency of attention kernels i.e., we do not find any evidence of TLB thrashing.
-Second, memory allocation using CUDA APIs incurs high latency because each allocation involves a round-trip to the OS kernel.
-We introduce several LLM-specific optimizations such as overlapping memory allocation with compute, opportunistically allocating pages ahead of time, and deferring memory reclamation.
-These optimizations hide the latency cost of memory allocation from end-users, making vAttention an efficient KV-cache memory manager.
-Overall, we make the following contributions in this paper:
-  • We present vAttention – a memory management approach that retains the virtual contiguity of KV-cache while enabling dynamic physical memory allocation.
-  • We implement vAttention in vLLM serving stack to show that it seamlessly adds dynamic memory management support to unmodified attention kernels of FlashAttention (fagithub, ) and FlashInfer (figithub, ).
-  • We compare vAttention against PagedAttention based alternatives of vLLM, FlashAttention and FlashInfer by evaluating Yi-6B, Llama-3-8B and Yi-34B on 1-2 A100 GPUs.
-Using FlashAttention’s non-paged kernel, vAttention outperforms vLLM by up to 1.99×1.$9$ in decode throughput, while improving the end-to-end LLM serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
+Therefore, in trying to enable dynamic allocation of physical memory, PagedAttention ends up changing the virtual memory layout of KV cache from contiguous to non-contiguous.
+We argue that this approach has several pitfalls (§3).
+First, it requires rewriting attention kernels, i.e., to enable de-referencing all tokens of the non-contiguous KV cache.
+Second, it forces developers to implement a memory manager in the serving framework, i.e., to stitch together dynamically allocated virtual memory blocks.
+Third, it adds runtime overhead in the critical path of both CPU and GPU execution.
+Table 1 provides empirical evidence and real-world experiences to support these arguments.
+The fundamental issue with PagedAttention and prior systems is that they rely on the reservation-based memory allocation method exposed by the GPU runtime.
+In this method (used by cudaMalloc), the runtime allocates both virtual and physical memory on the GPU meaning that physical memory is allocated even if the corresponding virtual memory is not accessed.
+This is in stark contrast to OS-based demand paging (Kwon et al., 2016; Panwar et al., 2019).
+We show that separating the allocation of virtual and physical memory allows for more effective KV cache memory management.
+To support our claim, we introduce vAttention (§5) — an approach that stores KV cache in contiguous virtual memory without committing physical memory ahead-of-time.
+vAttention decouples the allocation of virtual and physical memory using the CUDA virtual memory management (VMM) APIs (cud, 2024).
+In building vAttention, we find that using CUDA VMM support for KV cache management poses two key efficiency challenges for an LLM serving system (§6).
+First, memory allocation using CUDA VMM APIs incurs high latency because each allocation involves a round-trip to the OS kernel.
+We tackle latency issues with several LLM-specific optimizations such as overlapping memory allocation with compute, opportunistically allocating pages ahead of time, and deferring memory reclamation.
+Second, CUDA supports memory allocation only at the granularity of large pages, i.e., in multiples of 2MB.
+Use of large pages can create significant fragmentation.
+We address this challenge by modifying the open-source CUDA unified virtual memory driver, adding support for smaller 64KB pages.
+Our evaluation shows that use of 64KB pages has no negative impact on the performance of attention kernels, i.e., we do not find any evidence of TLB thrashing.
+Together, these optimizations mitigate fragmentation while hiding the latency cost of on demand memory allocation, making vAttention a simpler, portable and performant alternative to PagedAttention.
+Overall, we make the following contributions:
+  • We present vAttention – a memory management approach that retains the virtual contiguity of KV cache while enabling dynamic allocation of physical memory.
+Our implementation of vAttention in vLLM seamlessly adds dynamic memory allocation support to various unmodified attention kernels.
+  • We compare vAttention against PagedAttention-based alternatives of vLLM, FlashAttention-2 and FlashInfer on Yi-6B, Llama-3-8B and Yi-34B with 1-2 A100 GPUs.
+Using FlashAttention-2’s non-paged attention kernel, vAttention outperforms vLLM by up to 1.99×1.$9$ in decode throughput.
+In long-context scenarios, it also improves the end-to-end LLM serving throughput by up to 1.18×1.18$\times$1.18 × and 1.23×1.23$\times$1.23 × over PagedAttention based kernels of FlashAttention-2 and FlashInfer
+  • We demonstrate the portability benefit of vAttention with the recently launched FlashAttention-3 kernel (FA3 (Shah et al., 2024)).
+FA3 is optimized for the NVIDIA Hopper architecture and was not released with PagedAttention support.
+vAttention supports FA3 out-of-the-box, leading to 1.261.5×1.26-1.5$\times$1.26 - 1.5 × higher throughout over PagedAttention based FlashAttention-2.
 2. Background
 2.1. Large Language Models
-Given an input sequence, an LLM predicts the probability of an output sequence wherein a sequence is a set of tokens (vllmsosp, ).
-Each inference request begins with a prefill phase that processes all its prompt tokens in parallel.
-The prefill phase produces the first output token of a request.
-Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (sarathi2023, ).
-LLMs are built atop one of variants of the transformer architecture (attentionpaper, ).
-A transformer block contains two types of operators: position-wise and sequence-wise.
+Given an input sequence, an LLM predicts the probability of an output sequence: a sequence is a set of tokens (Kwon et al., 2023).
+Each inference request begins with a prefill phase that processes prompt tokens in parallel and produces the first output token of the request.
+Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (Agrawal et al., 2023).
+LLMs are built atop one of the variants of the transformer architecture (Vaswani et al., 2017); an LLM consists of multiple transformer blocks.
+Internally, a transformer block contains two types of operators: position-wise and sequence-wise (Yu et al., 2022).
 The former category includes feed-forward network, layer normalization, activation, embedding layer, output sampling layer, and residual connections whereas attention is a sequence-level operator.
-In this paper, we primarily focus on attention since it is the primary consumer of GPU memory in LLM inference.
-For the attention operator, the model first computes the query, key and value vectors from a given sequence of tokens $(x_{1},x_{2},....,x_{K})\in\mathbb{R}^{K\times E}$ where E represents the embedding size of the model.
+We primarily focus on attention since it is the primary consumer of GPU memory in LLM inference.
+Table 2 summarizes the notations used in the paper.
+For the attention operator, the model computes the query, key and value vectors from a given sequence of tokens $(x_{1},x_{2},....,x_{K})\in\mathbb{R}^{K\times E}$ where E represents the embedding size of the model.
 For each $x_{i}$, query, key and value vectors are computed as follows:
 (1) $q_{i}=W_{q}x_{i},\;\;\;k_{i}=W_{k}x_{i},\;\;\;v_{i}=W_{v}x_{i}$
-The resulting $k_{i}$ and $v_{i}$ are appended to the key and value vectors of the prior tokens of the corresponding request, producing two matrices $K,V\in\mathbb{R}^{L^{\prime}\times(H\times D)}$ where $L^{\prime}$ represents the context length of the request seen so far, H is the number of KV heads and D is the dimension of each KV head.
+The resulting $k_{i}$ and $v_{i}$ are appended to the key and value vectors of the prior tokens of the corresponding request, producing two matrices $K,V\in\mathbb{R}^{L^{\prime}\times(H\times D)}$ where $L^{\prime}$ represents the context length of the request seen so far, H is the number of KV heads of the model on a worker and D is the dimension of each KV head.
 Then, attention is computed as follows:
 (2) $Attention(q_{i},K,V)=softmax(\frac{q_{i}K^{T}}{scale})V$
+Symbol Definition
+$N$ Number of layers on a particular worker
+$H$ Number of KV heads on a particular worker
+$D$ Dimension of each attention head
+$P$ Number of bytes needed to store one element
+$B$ Maximum batch size
+$L$ Maximum context length supported by the model
+$L^{\prime}$ Context length of a request seen thus far
+Table 2. Notations used in the paper.
 The attention score is computed separately for each request in the batch.
+A request executes until the model generates a special end-of-sequence token or reaches the maximum context length for the request.
 Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention.
-Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations: we refer to this state as KV-cache.
-A request is executed until the model generates a special end-of-sequence token or reaches the maximum context length for the request.
-Structure of the KV-cache and terminology: An LLM consists of multiple layers of the transformer block and each layer maintains its own cache of keys and values.
-In this paper, we refer to the cache of all transformer blocks collectively as KV-cache while using the term K-cache or V-cache for keys and values, respectively.
-In deep learning frameworks, the K-cache (or V-cache) at each layer is typically represented as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request.
-We refer to a kernel implementation that computes attention scores over contiguously stored K and V as a non-paged kernel.
+Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations.
+We refer to this state of all layers collectively as KV cache.
+In systems prior to PagedAttention, the K cache (or V cache) at each layer of a worker is typically allocated as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request.
 2.2. Fragmentation and PagedAttention
-To improve serving throughput, production systems rely on batching which requires careful allocation of GPU memory.
+Serving LLMs with high throughput requires careful allocation of GPU memory.
 This is challenging because the total context length of a request is not known in advance.
-Serving systems worked around this challenge by pre-reserving KV-cache space assuming that each context is as long as the maximum length supported by the model (e.g., 200K for Yi-34B-200K).
+Serving systems worked around this challenge by pre-reserving KV cache space assuming that each context is as long as the maximum length supported by the model (e.g., 200K for Yi-34B-200K).
 vLLM shows that this strategy is prone to severe internal fragmentation.
 In fact, vLLM showed that prior reservation is suboptimal even if the context lengths are known in advance.
-This is because the per-request KV-cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
-Inspired by the OS-based virtual memory systems, vLLM proposed PagedAttention to mitigate fragmentation by dynamically allocating memory for the KV-cache.
-PagedAttention splits KV-cache into fixed-sized blocks and allocates memory for one block at a time.
+This is because the per-request KV cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
+Inspired by the OS-based virtual memory systems, vLLM proposed PagedAttention to mitigate fragmentation by dynamically allocating memory for the KV cache.
+PagedAttention splits KV cache into fixed-sized blocks and allocates memory for one block at a time.
 This way, vLLM allocates only as much memory as a request needs, and only when required – not ahead-of-time.
-Figure 1. PagedAttention involves two layers of memory management: one in user space and one in OS kernel space.
-3. Issues with the PagedAttention Model
-Despite being inspired by demand paging, the PagedAttention approach is different from it: PagedAttention mandates an implementation of paging in user space whereas conventional demand paging is transparent to applications.
+GPUs and page sizes: NVIDIA GPUs support multiple page sizes in the hardware (Pratheek et al., 2022; Nayak et al., 2021; Zhang et al., 2023a; pas, 2024).
+A single call to a CUDA VMM API can allocate one or more physical pages, which we refer to as a page-group.
+We use page-groups to support multiple allocation granularities for the KV cache, similar to how multiple page sizes are commonly used in conventional OS-based virtual memory systems (Kwon et al., 2016; Panwar et al., 2019).
+3. Issues with the PagedAttention Approach
+Despite being inspired by demand paging, the PagedAttention approach is different from it: PagedAttention implements demand paging in user space whereas conventional demand paging is transparent to applications.
 This section elaborates on issues that arise with such an approach.
 3.1. Requires Re-writing the Attention Kernel
-PagedAttention necessitates re-writing the attention kernel.
-This is because conventional implementations of the attention operator assume that the two input tensors K and V (Equation 2) are stored in contiguous memory.
-By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV-cache blocks.
-Writing correct and performant GPU kernels can be challenging for most programmers (fa-paged-crash, ).
-Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (flashattention, ; flashattention2, ; flashdecoding, ; flashdecoding++, ; flashinfer, ; h2o2023, ; flashattnhopper, ; groupedqueryattention, ; multiqueryattention, ; generatinglongsequence2019, ; longformer, ), and this trend is likely to continue.
+Conventional implementations of the attention operator assume that the two input tensors K and V (Equation 2) are stored in contiguous memory.
+By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV cache blocks.
+Writing correct and performant GPU kernels can be challenging for most programmers (fa-, 2023).
+Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (Dao et al., 2024; Dao, 2023; fla, 2023; Hong et al., 2024; Ye et al., 2024; Zhang et al., 2023b; Bikshandi and Shah, 2023; Ainslie et al., 2023; Shazeer, 2019; Child et al., 2019; Beltagy et al., 2020; Kamath et al., 2024), and this trend is likely to continue.
 In the PagedAttention model, keeping up with new research requires continued efforts in porting new optimizations to a PagedAttention-aware implementation.
 Production systems can therefore easily fall behind research, potentially losing performance and competitive advantage.
-To provide an example, Table 6 shows that the paged kernel of vLLM is already up to 2.8×2.8$\times$2.8 × slower than the FlashAttention kernel (groupedqueryattention, ).
+To provide an example, Table 7 shows that the paged kernel of vLLM is already up to 2.8×2.8$\times$2.8 × slower than the FlashAttention-2 kernel (Ainslie et al., 2023).
+Figure 1. PagedAttention involves two layers of memory management: one in user space and one in OS kernel space.
 3.2. Adds Redundancy in the Serving Framework
-PagedAttention makes an LLM serving system responsible for managing the mappings between KV-cache and dynamically allocated memory blocks.
-For example, consider a request that allocates four KV-cache blocks over time (left half of Figure 1).
+PagedAttention makes an LLM serving system responsible for managing the mappings between KV cache and dynamically allocated memory blocks.
+For example, consider a request that allocates four KV cache blocks over time (left half of Figure 1).
 These blocks are usually non-contiguous in virtual memory.
-During the computation of Equation 2, PagedAttention kernel needs to access all the elements of the four KV-cache blocks.
-To facilitate this, the serving system needs to track the virtual memory addresses of KV-cache blocks and pass them to the attention kernel at runtime.
+During the computation of Equation 2, PagedAttention kernel needs to access all the elements of the four KV cache blocks.
+To facilitate this, the serving system needs to track the virtual memory addresses of KV cache blocks and pass them to the attention kernel at runtime.
 This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half in Figure 1).
+3.3. Performance Overhead
+3.3.1. Runtime overhead on the GPU
+PagedAttention slows down attention computation by adding extra code in the critical path of execution.
+For example, the vLLM paper acknowledges that the PagedAttention-based implementation was $20-26\%$ slower than the corresponding none-paged FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (see Figure 18a in  (Kwon et al., 2023)).
+In addition, Figure 2 shows that incorporating PagedAttention has also added a significant performance overhead in other state-of-the-art kernel libraries.
+For example, PagedAttention based prefill kernels of FlashAttention-2 and FlashInfer are up to 37% and 42% slower than the non-paged kernels in the corresponding libraries.
+Our analysis reveals that the number of instructions executed in PagedAttention kernels is $7-13\%$ higher than the non-paged kernels.
+Caching page indices also increases register pressure, causing register spilling.
+To highlight another example of difficulty involved in writing an efficient paged kernel, Figure 3 shows that the performance of vLLM’s paged decode kernel is significantly worse with large block sizes of 64 and 128.
+Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in the L1 cache.
 Figure 2.
 Overhead of PagedAttention in prefill kernels (model: Llama-3-8B, one A100 GPU).
-Numbers on top show overhead over the corresponding non-paged implementation of FlashAttention (FA) and FlashInfer (FI).
-3.3. Performance Overhead
-3.3.1. Runtime overhead on the GPU
-PagedAttention slows down attention computation by adding extra code in the critical path.
-For example, vLLM acknowledges that their PagedAttention-based implementation was $20-26\%$ slower than the original FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (see Figure 18a in  (vllmsosp, )).
-In addition, Figure 2 shows that incorporating PagedAttention has also added a significant performance overhead in other state-of-the-art kernel libraries.
-For example, PagedAttention based prefill kernels of FlashAttention and FlashInfer are up to 28% and 24% slower than the non-paged kernels in the corresponding libraries.
-Our analysis reveals that the number of instructions executed in PagedAttention kernels is $7-13\%$ higher than the non-paged kernels.
-To highlight another example of difficulty involved in writing an efficient attention kernel with paging support, Figure 3 shows that the performance of vLLM’s paged decode kernel is significantly worse with larger block sizes of 64 and 128.
-Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in L1 cache.
+Numbers on top show overhead over the corresponding non-paged implementation of FlashAttention-2 (FA2) and FlashInfer (FI).
+Figure 3. Latency of vLLM’s paged decode kernel is sensitive to block size (model: Llama-3-8B, one A100 GPU).
 3.3.2. Runtime overhead on the CPU
@@
 To enable PagedAttention, a serving system needs to supply Block-Tables to the attention kernel.
-In vLLM, the latency of preparing a Block-Table depends on batch composition and grows proportional to max_num_blocks $\times$ batch_size where max_num_blocks refers to the number of KV-cache blocks in the longest request of the batch.
-This is because vLLM manages a Block-Table as a 2D tensor and aligns the number of KV-cache blocks in each request by padding unoccupied slots with zeros.
+In vLLM, the latency of preparing a Block-Table depends on batch composition and grows proportional to max_num_blocks $\times$ batch_size where max_num_blocks refers to the number of KV cache blocks in the longest request of the batch.
+This is because vLLM manages a Block-Table as a 2D tensor and aligns the number of KV cache blocks in each request by padding unoccupied slots with zeros.
 If a batch contains a few long and many short requests, such padding results in a significant overhead.
 In our earlier experiments, we observed that Block-Table preparation in vLLM was contributing 30% latency in decode iterations.
-While a recent fix (vllm-blocktable-issue, ) has mitigated some of this overhead, we find that it can still be as high as $10\%$.
-High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11%, from 412 tokens/sec to 365 tokens/sec (tensorrt-perf-decay, ).
+While a recent fix (vll, 2024b) has mitigated some of this overhead, we find that it can still be as high as $10\%$.
+High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11% (ten, 2023).
 This issue was attributed to the Python runtime of TensorRT-LLM and moving to a C++ runtime can mitigate the CPU overhead.
-However, doing so would be an unwelcome change for most programmers.
-Figure 3. Latency of vLLM’s paged decode kernel is sensitive to block size (model: Llama-3-8B, one A100 GPU).
-Overall, this section shows that the PagedAttention model adds a significant programming burden while also being inefficient.
-vAttention introduces a more systematic approach to dynamic KV-cache memory management by leveraging the existing system support for demand paging.
+However, doing so requires non-trivial programming effort.
+Overall, this section shows that PagedAttention adds a significant programming burden while also being inefficient.
+In vAttention, we propose a more principled approach to dynamic KV cache memory management.
 However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads from a memory management perspective.
 4. Insights into LLM Serving Systems
-To highlight the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism.
-We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure the throughput and memory requirement of the decode phase (see §5.4 for our discussion and optimizations for the prefill phase).
-Observation-1: On a per-iteration basis, KV-cache memory requirement is known in advance.
-This is a dream property for a memory allocator, thanks to the auto-regressive decoding wherein every forward pass of the model generates one token per request.
-Therefore, with every iteration, the KV-cache memory footprint of a request grows uniformly by one token as long as the request is active.
-Observation-2: KV-cache does not require high memory allocation bandwidth.
+To understand the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism (TP).
+We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure the throughput and memory requirement of the decode phase (see §6 for our discussion and optimizations for the prefill phase).
+Observation-1: KV cache memory requirement is predictable on a per-iteration basis.
+Due to auto-regressive decoding, once a request enters the decode phase, its KV cache size increases uniformly by one token per iteration.
+This allows the serving system to determine in advance if additional memory will be required during the iteration’s execution.
+[^1] Note that this prediction is possible only at the granularity of individual iterations; the total KV cache requirement of a request remains unknown as it depends on the total number of output tokens in the request.
+Observation-2: KV cache does not require high memory allocation bandwidth.
 The memory footprint of a single token across all layers is typically few 10s-100s of kilobytes of memory.
@@
 Further, each iteration runs for 10s-100s of milliseconds implying that a request requires at most a few megabytes of memory per second.
-While batching improves system throughput (sarathi2023, ; sarathiserve2024, ; patel2023splitwise, ; distserve2024, ), the number of tokens generated per second plateaus beyond a certain batch size (Figure 4, left).
-This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 128 for Yi-34B).
-For all the models we studied, we observe that the highest memory allocation rate is at most 600MB per second (Figure 4, right).
-Figure 4. Decode throughput (left) and the rate of physical memory allocation (right) saturate at large batch sizes.
-In vAttention, we leverage these observations to implement an efficient dynamic memory management system for KV-cache.
-In the next section, we begin with a high-level overview of vAttention (§5.1), then discuss how vAttention is used for serving LLMs (§5.3) and finally describe our optimizations (§5.4).
-5. vAttention: Design and Implementation
-To facilitate dynamic allocation of physical memory to unmodified attention kernels, vAttention leverages system support for demand paging instead of implementing it in user space.
-5.1. Design Overview
-vAttention builds on the ability to allocate virtual memory and physical memory separately.
-Specifically, we allocate a large contiguous buffer for the KV-cache in virtual memory ahead-of-time (similar to reservation-based allocators) while deferring the allocation of physical memory to runtime (similar to PagedAttention).
-This way, vAttention preserves virtual contiguity of KV-cache without wasting physical memory.
-This approach is feasible because memory capacity and fragmentation are limiting factors only for physical memory whereas virtual memory is abundant e.g., modern 64-bits systems provide a 128TB user-managed virtual memory per process11164-bits systems use only 48 bits for virtual addresses today, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space..
+While batching improves system throughput (Agrawal et al., 2023, 2024b; Patel et al., 2024; Zhong et al., 2024), the number of tokens generated per second plateaus beyond a certain batch size (4(a)).
+This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 256 for Yi-34B).
+For all the models we studied, we observe that the highest memory allocation rate is at most 750MB per second (4(b)).
+vAttention leverages these observations to optimize KV cache memory management.
+(a) Decode throughput.
+(b) Rate of memory allocation.
+Figure 4. Decode throughput (top) and the rate of physical memory allocation (bottom) saturate at large batch sizes.
 Figure 5.
-Dynamic memory management in vAttention for a single K-cache (or V-cache) tensor.
+Dynamic memory management in vAttention for a single K cache (or V cache) tensor.
 (a) shows a virtual tensor for a batch of two requests with no physical memory allocation yet.
@@
 (e) when R3 arrives, vAttention assigns R1’s tensor to it which is already backed by physical memory.
-Pre-reserving virtual memory
-Since virtual memory is abundant, we pre-allocate it in size that is large enough to hold the KV-cache of the maximum batch size (configurable) that needs to be supported.
-In doing so, we assume that each request would generate as many tokens as the maximum sequence length supported by the model.
-Number of virtual memory buffers
-A serving framework typically maintains separate K and V tensors for each layer of the model: we refer to them individually as K-cache and V-cache.
-We allocate separate virtual memory buffers for K-cache and V-cache.
-For a single GPU job, this requires pre-reserving $2\times N$ buffers where $N$ is the number of layers in the model.
-In a multi-GPU job, each worker reserves $2\times N^{\prime}$ buffers where $N^{\prime}$ is the number of layers managed by that worker ($N^{\prime}=N$ with tensor-parallelism whereas $N^{\prime}<N$ with pipeline-parallelism).
+5. vAttention: Design and Implementation
+Our primary observation is that physical memory fragmentation can be avoided without making KV cache non-contiguous in virtual memory.
+To realize this, vAttention decouples the allocation of virtual memory from physical memory by leveraging system support for demand paging (instead of implementing demand paging in user space, as in PagedAttention).
+5.1. Design Overview
+vAttention employs distinct allocation policies for virtual and physical memory.
+Specifically, we allocate a large contiguous buffer for the KV cache in virtual memory ahead-of-time (similar to systems prior to PagedAttention) while deferring the allocation of physical memory to runtime (similar to PagedAttention).
+This design preserves virtual contiguity of KV cache without fragmenting physical memory.
+Note that this approach could fragment and waste virtual memory.
+However, this is not an issue since virtual memory is abundant, e.g., modern 64-bit systems provide a 128TB user-addressable virtual memory per process.22264-bit systems use only 48 bits for virtual addresses today, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space.
+5.1.1. Pre-reserving virtual memory
+Since virtual memory is abundant, we pre-allocate it in size that is large enough to hold the KV cache of the maximum batch size (configurable) that needs to be supported.
+In doing so, we assume that each request’s context length is same as the maximum supported by the model.
+5.1.2. Number of virtual memory buffers
+A serving framework maintains separate K and V tensors for each layer of the model. Therefore, we reserve $2\times N$ buffers on a worker where $N$ is the number of layers managed by that worker.
 Latency (microseconds)
@@
 Table 3.
-CUDA VM (virtual memory) APIs.
+CUDA VMM APIs.
 * represents APIs that we use once while instantiating or terminating the serving framework.
 Rest of the APIs are used for (un)mapping physical memory pages at runtime.
-CUDA APIs (prefixed with cu) support only 2MB allocation sizes, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
-Size of a virtual memory buffer
-The maximum size of a buffer is $BS=B\times L\times S$ where B is the maximum batch size, L is the maximum context length supported by the model, and $S$ is the size of a single token’s per-layer K-cache (or V-cache) on a worker.
-Further, $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16).
-Note that $S$ is constant for a given model configuration.
-Consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2).
+CUDA APIs (prefixed with cu) support only 2MB pages, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
+5.1.3. Size of a virtual memory buffer
+The maximum size of a buffer is $BS=B\times S$ where B is the maximum batch size and $S$ is the maximum size of a single request’s per-layer K cache (or V cache) on a worker.
+Further, $S=L\times H\times D\times P$, where $L$ is the maximum context length supported by the model, $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16).
+As an example, consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2).
 In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$.
-For this model, the maximum size of K-cache (or V-cache) per-worker per-layer is $S=200MB$ $(200K*4*128*2)$.
+For this configuration, $S=200MB$ $(200K*4*128*2)$.
 Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$).
 Therefore, the total virtual memory requirement for 60 layers is 120 buffers of 100GB each (12TB total).
-Note that the amount of virtual address space available grows with the number of GPUs e.g., with two TP workers, the amount of virtual address space available is 256TB.
-Therefore, virtual memory allocations can be satisfied easily.
-Figure 5 shows how vAttention allocates physical memory pages dynamically.
-5.2. Leveraging Low-level CUDA Support
-The standard GPU memory allocation interface cudaMalloc does not support demand paging i.e., it allocates virtual memory and physical memory at the same time.
-However, recent CUDA versions provide programmers a fine-grained control over virtual and physical memory (cudavirtualmemory, ; gmlakeasplos24, ).
-We leverage these low-level APIs in vAttention.
+Note that the size of virtual address space available grows with the number of workers, e.g., with two TP workers, the total size of user-addressable virtual address space is 256TB.
+Therefore, virtual memory is always plentiful to satisfy large allocations.
+Figure 5 shows an example of how vAttention allocates physical memory pages dynamically.
+5.2. Leveraging CUDA Virtual Memory Support
+The standard GPU memory allocation interface cudaMalloc does not support demand paging, i.e., it allocates virtual memory and physical memory at the same time.
+However, recent CUDA versions provide programmers a fine-grained control over managing virtual and physical memory, including support for decoupling their allocations (cud, 2024; Guo et al., 2024).
+We leverage these low-level APIs.
 5.2.1. CUDA virtual memory APIs
-Table 2 provides an overview of CUDA APIs that allow separating the allocation of virtual memory from physical memory.
+Table 3 provides an overview of CUDA VMM APIs that allow decoupling the allocation of virtual memory from physical memory.
 The allocation granularity depends on the page size used by the GPU.
 Further, the size of a virtual memory buffer or a physical memory handle must be a multiple of the physical memory allocation granularity.
-It is also important to highlight that physical memory pages can be allocated (or de-allocated) to sub-regions in a virtual memory buffer independently of other sub-regions.
-For simplicity, we refer to the granularity of physical memory allocation as page size.
+Physical memory pages can be allocated to (or de-allocated from) sub-regions in a virtual memory buffer independently of other sub-regions.
+Figure 6.
+An LLM serving system interacts with vAttention for KV cache management with a set of simple APIs listed in Table 4.
+All other memory objects (e.g., activations) are allocated by the standard PyTorch caching allocator.
 5.2.2. Extending PyTorch caching allocator
-KV-cache is a collection of tensors.
+KV cache is a collection of tensors.
 In current deep learning frameworks such as PyTorch, a tensor allocated via APIs such as torch.empty comes with pre-allocated physical memory.
-This is because the PyTorch caching allocator uses the cudaMalloc interface to allocate GPU memory (both virtual and physical).
+This is because the PyTorch caching allocator relies on the cudaMalloc interface (Figure 6).
 Relying on the low-level API support from CUDA, we extend the PyTorch caching allocator to allow an application to reserve a virtual memory buffer for a tensor without committing physical memory ahead-of-time.
 We refer to tensors allocated via these APIs as virtual tensors.
-5.2.3. Request-level KV-cache indexing
-Note that each virtual tensor represents the K-cache (or V-cache) of a layer for the maximum batch size B.
+5.2.3. Request-level KV cache indexing
+A virtual tensor represents the K cache (or V cache) of a layer for the maximum batch size B.
 In these tensors, different requests occupy different non-overlapping sub-regions (say sub-tensors).
 We locate the sub-tensor of a request with a unique integer identifier reqId that lies in the range of 00 to $B-1$ (note that at most $B$ requests run simultaneously).
-The K-cache (or V-cache) offset of a request’s sub-tensor in the virtual tensor of the entire batch is $\texttt{reqId}\times S$ where $S$ is the maximum K-cache (or V-cache) size of a request on a worker.
+The K cache (or V cache) offset of a request’s sub-tensor in the virtual tensor of the entire batch is $\texttt{reqId}\times S$ where $S$ is the maximum size of per-layer K cache (or V cache) of a request on a worker.
 The request identifier reqId is allocated by vAttention.
-5.3. Serving LLMs with vAttention
-We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers.
-Our library exposes a set of simple APIs to the serving framework (see Table 3 and Algorithm 1).
-5.3.1. Initial setup
-When the serving framework starts, each model worker loads the vAttention library and configures it with model parameters $N^{\prime},H,D,P$, $B$ and a preferred page size via the init API.
-Internally, vAttention reserves $2\times N^{\prime}$ virtual tensors (using our modified PyTorch caching allocator) for the KV-cache at each worker.
-These virtual tensors are reserved for the lifetime of the serving application.
-In addition, vAttention also pre-allocates physical memory pages during initialization.
-However, these pages are not mapped into the KV-cache yet.
-5.3.2. Scheduling a new request
-When a new request is scheduled for the first time, the serving framework obtains a new reqId from vAttention via alloc_reqid.
-All subsequent memory management operations of the request are tagged with this reqId.
 APIs Description
 init Initializes vAttention with model parameters.
-arguments: $N^{\prime},B,L,H^{\prime},P$, page_size.
-return value: a list of KV-cache tensors.
+arguments: $N,B,L,H,D,P$, page-group-size.
+return value: a list of KV cache tensors.
 alloc_reqid Allocates an unused reqId and marks it active
@@
 return value: None
-step Ensures physical memory pages are mapped
+step Ensures KV cache tensors are backed by physical pages up to the current context length of each active request
 arguments: an array of size B containing sequence lengths of each reqId
 return value: 0 (success), -1 (failure).
-Table 3. Key APIs that vAttention exposes to a serving framework for dynamic KV-cache memory management.
-5.3.3. Model execution
-Before scheduling a batch for execution, the framework needs to ensure that the KV-cache sub-tensors of each active request are backed by physical memory.
-For this purpose, before dispatching the first kernel of an iteration to the GPU, the framework invokes the step API, specifying the current context length of each request (context length is set to 0 for each inactive reqId).
-Internally, vAttention ensures that enough physical pages are mapped for each active reqId before returning execution back to the framework.
-If vAttention cannot satisfy the memory demand, it returns with a failure in response to which a serving framework can preempt one or more requests to allow forward progress (this is similar to vLLM’s default behavior).
-We leave more sophisticated policies such as swapping out KV-cache to CPU memory as future work.
-Depending on whether a request is in the prefill phase or decode phase, different number of physical memory pages may need to be mapped for a given iteration.
-The prefill phase processes the input tokens of given prompt in parallel and populates one slot in the K-cache (and V-cache) of the request at each layer of the model.
-Therefore, the number of pages needed to be mapped depends on the number of prompt tokens being scheduled.
-If the total K-cache size of all prompt tokens at one layer of the model is $s$ and page size is $t$, then each worker needs to ensure that at least $(s+t-1)/t$ physical memory pages are mapped in each of the $2\times N^{\prime}$ KV-cache sub-tensors of the given reqId.
-For a request in the decode phase, the number of new pages required is at most one per request.
-This is because each iteration produces only one output token for a request.
-vAttention internally tracks the number of pages mapped for each request and maps a new page only when the last page allocated to that request is fully utilized.
-5.3.4. Request completion
-A request terminates when a user specified context length or the maximum context length supported by the model is reached, or when the model produces a special end-of-sequence token.
-The framework notifies vAttention of a request’s completion with free_reqid.
-Internally, vAttention may unmap the pages of a completed request or defer them to be freed later.
+Table 4. Key APIs that vAttention exposes to a serving framework for dynamic KV cache memory management.
+5.3. Serving LLMs with vAttention
+We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers.
+Our library exposes a set of simple APIs to the serving framework (shown in Figure 6Table 4 and Algorithm 1).
+For simplicity, we discuss the use of these APIs from a single worker’s perspective; all workers behave the same.
+5.3.1. Initial setup
+When the serving framework starts, each model worker loads the vAttention library and configures it with model parameters $N,H,D,P$, $B$ and a preferred page-group size (§6.2) via the init API (line 4 in Algorithm 1).
+Internally, vAttention reserves $2\times N$ virtual tensors on the worker, as shown in Figure 5(a), where $N$ is the number of layers hosted by the worker.
+These virtual tensors are reserved for the lifetime of the serving application.
+In addition, vAttention also pre-allocates physical memory pages at each worker during initialization.
+However, these pages are not mapped into the KV cache at this point.
 Algorithm 1 Using vAttention in a serving framework.
@@
 3:req_batch_idx $\leftarrow$ dict()
-4:vattention.init_cache(config_params)
+4:vattention.init(config_params)
 5:while !request_pool.is_empty() do
@@
 24:end while
-5.4. Optimizations
-There are two primary challenges in using CUDA virtual memory support for serving LLMs.
-First, cuMemCreate currently allocates a minimum of 2MB physical memory page.
-Large pages can waste physical memory due to internal fragmentation.
-Second, invoking CUDA APIs incurs high latency.
-This section details a set of simple-yet-effective optimizations that we introduce to overcome these limitations.
-5.4.1. Mitigating internal fragmentation
+5.3.2. Scheduling a new request
+When a new request is scheduled for the first time, the serving framework obtains a new reqId from vAttention via alloc_reqid (line 8).
+All subsequent memory management operations of the request are tagged with this reqId.
+5.3.3. Model execution
+Before scheduling a batch for execution, the framework needs to ensure that the KV cache sub-tensors of each active request are backed by physical memory (Figure 5(b) and (c)).
+For this purpose, before dispatching the first kernel of an iteration to the GPU, the framework invokes the step API (line 13), specifying the current context length of each request (context length is set to 0 for each inactive reqId).
+Internally, vAttention ensures that enough physical pages are mapped for each active reqId before returning execution back to the framework.
+If vAttention cannot satisfy the memory demand, it returns with a failure in response to which a serving framework can preempt one or more requests to allow forward progress (this is similar to vLLM’s default behavior).
+We leave more sophisticated policies such as swapping out KV cache to CPU memory as future work.
+Depending on whether a request is in the prefill phase or decode phase, different amount of physical memory may need to be mapped for a given iteration.
+The prefill phase processes the input tokens of a given prompt in parallel.
+Therefore, the amount of physical memory needed to be mapped depends on the number of prompt tokens being scheduled.
+If the total K cache size of all prompt tokens at one layer of the model is $s$ and page-group size is $t$, then each worker needs to ensure that at least $(s+t-1)/t$ page-groups are mapped in each of the $2\times N$ KV cache sub-tensors of the given reqId.
+For a request in the decode phase, the number of new page-groups required is at most one per virtual tensor.
+This is because each iteration produces only one output token for a request.
+vAttention internally tracks the number of page-groups mapped for each request and maps new page-groups only when prior page-groups are about to be exhausted.
+5.3.4. Request completion
+A request terminates when it reaches user specified or the maximum context length supported by the model, or when the model produces a special end-of-sequence token.
+The framework notifies vAttention of a request’s completion with free_reqid (line 19).
+Internally, vAttention may unmap the physical pages of a completed request or defer them to be freed later (§6.1.2).
+Supporting continuous batching: Continuous batching poses one challenge in computing attention in our design.
+When a request from somewhere in the middle of a batch exits, it creates an unused hole in the virtual tensors of KV cache.
+This layout is not supported by implementations that expect the query (Q) and KV cache to be of the same size in batch ($B$) dimension.
+However, FlashAttention provides rich API support to address this issue; its argument cache_batch_idx allows Q and KV cache to have different batch sizes and also be arranged in arbitrary order (i.e., batch index 0 in Q can be mapped to batch index 1 in KV cache).
+vAttention benefits from this API support in terms of both performance and ease of programming; when the request composition of a batch changes, we update cache_batch_idx of running requests such that their Q tensors map to their respective KV cache based on their reqId.
+6. Optimizations
+There are two challenges in using CUDA virtual memory support for serving LLMs.
+First, invoking CUDA VMM APIs at runtime incurs high latency.
+Second, cuMemCreate currently allocates memory only at the granularity of large pages, i.e., multiples of 2MB.
+Use of large pages can waste physical memory due to internal fragmentation.
+This section details a set of simple-yet-effective optimizations that we introduce to overcome these challenges.
+6.1. Hiding Latency of Memory Allocation
+The serving framework invokes the step API in every iteration.
+The latency of step depends on the number of page-groups that need to be mapped into KV cache.
+Consider, for example, that the KV cache of one request needs to be extended for Yi-34B which has 60 layers.
+This requires 120 calls to cuMemMap + cuMemSetAccess each of which takes about 40 microseconds.
+Therefore, growing the KV cache of one request by new page-groups (two per layer) adds about 5 millisecond latency to the corresponding iteration.
+The latency overhead grows proportional to the number of requests that need new page-groups in a given iteration.
+We propose the following optimizations to hide this latency:
+6.1.1. Overlapping memory allocation with compute (decode phase)
+We leverage the predictability of memory demand to overlap memory allocation with computation.
+In particular, note that each iteration produces a single output token for every decode request.
+Therefore, memory demand for a decode iteration is known ahead-of-time.
+Further, in the decode phase, a request requires at most one new page-group for each of its virtual tensors.
+vAttention keeps track of the current context length and how much physical memory is already mapped for each request.
+Using this information, it determines when a request would need more memory and uses a background thread to allocate new page-groups when the preceding iteration is executing.
+For example, consider that a request R1 would require more physical memory in iteration i.
+When the serving framework invokes step API in iteration i-1, vAttention launches a background thread that maps page-groups for iteration i.
+Since per-iteration latency is typically in the range of 10s-100s of milliseconds, the background thread has enough time to prepare physical memory mappings for an iteration before it starts executing.
+This way, vAttention hides the latency of CUDA APIs by performing memory allocations out of the critical path.
+6.1.2. Deferred reclamation + eager allocation (prefill phase)
+We observe that allocating physical memory for the prefill phase can be avoided in many cases.
+Consider that a request R1 completed in iteration i and a new request R2 joins the running batch in iteration i+1.
+To avoid allocating page-groups to R2 from scratch, vAttention simply defers the reclamation of R1’s page-groups (Figure 5(d)) and assigns R1’s reqId to R2.
+This way, R2 uses the same tensors for its KV cache that R1 was using – which are already backed by physical pages (Figure 5(e)).
+Therefore, new allocations are required only if R2’s context length is higher than R1.
+We further optimize the prefill phase by proactively allocating a small number of page-groups ahead of time.
+For this purpose, we try to keep a certain number of page-groups mapped into the virtual tensors of one of the inactive reqId.
+When a new request arrives, we allocate this reqId.
+At the same time, we identify a new reqId to be allocated next and eagerly map physical page-groups for it.
+In most cases, these eager allocations obviate the need to allocate physical memory in the critical path of prefill execution.
+Finally, we trigger memory reclamation only when the number of page-groups cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory).
+We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
+Figure 7.
+Prefill throughput.
+vAttention backed systems outperform the paged counterparts of both FlashAttention-2 and FlashInfer.
+Throughput for longer contexts is lower due to the quadratic complexity of prefill attention.
+6.2. Mitigating Internal Fragmentation
 We mitigate internal fragmentation by reducing the granularity of physical memory allocation.
-NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB (mcmgpus, ; gpu-mismanaged, ; gpu-reverse-engg-tlb1, ; pascalmmu, ).
+NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB (Pratheek et al., 2022; Nayak et al., 2021; Zhang et al., 2023a; pas, 2024).
 Therefore, in principal, physical memory can be allocated in any multiple of 4KB sizes.
-The simplest way to achieve this would be to extend the existing CUDA virtual memory APIs (listed in Table 2) to also support allocating smaller pages (similar to how mmap in Linux supports multiple page sizes).
-Unfortunately, the CUDA APIs are implemented in the closed-source NVIDIA drivers which makes it impossible for us to modify their implementation.
+The simplest way to achieve this would be to extend the existing CUDA VMM APIs (listed in Table 3) to also support allocating smaller pages (similar to how mmap in Linux supports multiple page sizes (Kwon et al., 2016; Panwar et al., 2019)).
+Unfortunately, the CUDA VMM APIs are implemented in the closed-source NVIDIA drivers which makes it impossible for us to modify their implementation.
 Fortunately, some part of NVIDIA drivers (particularly related to unified memory management) is open-source.
@@
 The second column in Table 3 shows our new APIs: most of our APIs have a one-to-one relationship with existing CUDA APIs except for vMemMap that combines the functionality of cuMemMap and cuMemSetAccess, and vMemRelease that combines the functionality of cuMemUnmap and cuMemRelease for simplicity.
-In contrast to CUDA APIs, our APIs can allocate memory in 64KB, 128KB and 256KB page sizes.
-A serving framework can configure a desired page size in vAttention while initializing it.
-The last set of columns in Table 2 shows the latency of each API with different page sizes.
-5.4.2. Hiding memory allocation latency
-The serving framework invokes the step API in every iteration.
-The latency of step depends on how many new pages need to be mapped in the virtual tensors of KV-cache.
-Consider, for example, that the KV-cache of one request needs to be extended for Yi-34B which has 60 layers.
-This requires 120 calls to cuMemMap + cuMemSetAccess each of which takes about 40 microseconds.
-Therefore, growing the KV-cache of one request by new pages (at all layers) adds about 5 millisecond latency to the corresponding iteration.
-Further, the latency overhead grows proportional to the number of requests that need new pages in a given iteration.
-We propose the following optimizations to hide the latency of allocation:
-Overlapping memory allocation with compute
-We leverage the predictability of memory demand to overlap memory allocation with computation.
-In particular, note that each iteration produces a single output token for every decode request.
-Therefore, memory demand for a decode iteration is known ahead-of-time.
-Further, in the decode phase, a request requires at most one new page.
-vAttention keeps track of the current context length and how many physical memory pages are already mapped for each request.
-Using this information, it determines when a request would need a new page and uses a background thread to allocate a new page when the preceding iteration is executing.
-For example, consider that a request R1 would require a new page in iteration i.
-When the serving framework invokes step API in iteration i-1, vAttention launches a background thread that maps physical memory pages for iteration i.
-Since iteration latency is typically in the range of 10s-100s of milliseconds, the background thread has enough time to prepare physical memory mappings for an iteration before it starts executing.
-This way, vAttention hides the latency of CUDA APIs by mapping physical pages in the KV-cache tensors out of the critical path.
-Note that in every iteration, step API still needs to ensure that physical pages required for the current iteration are actually mapped.
-If not, required pages are mapped synchronously.
-Deferred reclamation + eager allocation
-We observe that allocating physical memory for a new request can be avoided in many cases.
-Consider that a request R1 completed in iteration i and a new request R2 joins the running batch in iteration i+1.
-To avoid allocating new pages to R2 from scratch, vAttention simply defers the reclamation of R1’s pages and assigns R1’s reqId to R2.
-This way, R2 uses the same tensors for its KV-cache that R1 was using – which are already backed by physical pages (for security reasons, we still zero-fill such pages before allowing a new request to use them).
-Therefore, new pages are required only if the context length of R2 is bigger than that of R1.
-We further optimize memory allocation by proactively allocating a small number of physical pages ahead of time.
-For this purpose, we try to keep a certain number of pages (configurable based on expected workload) mapped into the virtual tensors of one of the inactive reqId.
-When a new request arrives, we allocate this reqId and identify a new reqId to be allocated next and map physical pages for it.
-In most cases, these eager allocations obviate the need to allocate physical pages in the critical path of prefill execution.
-Finally, we trigger memory reclamation only when the number of physical pages cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory).
-We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
+In contrast to CUDA VMM APIs, our APIs allocate physical memory in 64KB, 128KB and 256KB sized page-groups.
+A serving framework can configure a desired page-group size in vAttention while initializing it (we use the standard 2MB pages if the configured page-group size is 2MB).
+Table 3 shows the latency of each API with different page-group sizes.
 7. Evaluation
+Our evaluation answers the following questions:
+  • How does vAttention impact the performance of attention kernels and the overall performance of prefill and decode phases (§7.1, §7.2)?
+  • How does vAttention impact the end-to-end LLM serving throughput (§7.3, §7.4, §7.5)?
+  • What is the effect of each of our optimizations (§7.6)?
+Models and hardware: We evaluate three models Yi-6B (yi-, 2024b), Llama-3-8B (lla, 2024) and Yi-34B (yi-, 2024a).
+We conduct most of our evaluation on a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 5).
+Each GPU has 80GB physical memory.
+We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B.
+To demonstrate the portability benefit of vAttention, we use 1–2 H100 GPUs.
+Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different (Agrawal et al., 2024b; Holmes et al., 2024; Zhong et al., 2024).
+Attention kernels used for these two phases are also different and hence we evaluate them separately.
+The prefill phase requires one time memory allocation potentially spanning multiple pages.
+In comparison, the decode phase requires incremental memory allocation over the lifetime of a request (Kwon et al., 2023).
+We define prefill throughput as the number of prompt tokens processed per second, and decode throughput as the number of tokens generated per second.
 Model Hardware ##\## Q Heads ##\## KV Heads ##\## Layers
@@
 Table 5. Models and hardware used for evaluation.
-Figure 6.
-Prefill throughput.
-vAttention backed systems outperform the paged counterparts of both FlashAttention and FlashInfer.
-Throughput for longer contexts is lower due to the quadratic complexity of prefill attention.
-Our evaluation seeks to answer the following questions:
-  • How does vAttention perform for prefill and decode phases (§6.1, §6.2)?
-  • How vAttention impacts LLM serving throughput (§6.3)?
-  • What is the effect of each of our optimizations (§6.4)?
-Models and hardware: We evaluate three models Yi-6B (yi-6b-200k-hf, ), Llama-3-8B (llama-8b-hf, ) and Yi-34B (yi-34b-200k-hf, ), using a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 4).
-Each GPU has 80GB physical memory.
-We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B.
-Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different (sarathiserve2024, ; splitfuse2024, ; distserve2024, ).
-Attention kernels used for these two phases are also different and hence we also evaluate them separately.
-The prefill phase requires one time memory allocation potentially spanning multiple pages.
-In comparison, the decode phase requires incremental memory allocation over the lifetime of a request (vllmsosp, ).
-We define throughput as the number of tokens processed (or generated) per second.
 Serving framework: For a fair comparison, we use vLLM v0.2.7 as a common serving framework in all our experiments.
-We integrated state-of-the-art kernel libraries of both FlashAttention v2.5.9 (flashattention, ; flashattention2, ; fagithub, ) and FlashInfer v0.4.0 (figithub, ) as attention back-ends into vLLM, and further added support for dynamic memory allocation via vAttention to their non-paged kernels.
-While FlashAttention and FlashInfer are both based on the same underlying techniques (e.g., FlashDecoding (flashdecoding, )), they use different Block-Table formats; the former use a simple lookup table whereas the latter uses a compressed Block-Table to aid fetching KV-cache block locations.
+We integrated state-of-the-art kernel libraries of both FlashAttention-2 v2.5.9 (Dao et al., 2024; Dao, 2023; fag, 2022) and FlashInfer v0.4.0 (fig, 2023) as attention back-ends into vLLM, and further added support for dynamic memory allocation via vAttention to their non-paged kernels.
+While FlashAttention-2 and FlashInfer are both based on the same underlying techniques (e.g., FlashDecoding (fla, 2023)), they use different Block-Table formats; the former use a simple lookup table whereas the latter uses a compressed Block-Table to optimize lookups.
 Baselines: We compare performance obtained by using the non-paged attention kernels (backed by vAttention for dynamic memory allocation), and their paged counterparts.
 In addition, we also compare against vLLM’s decode kernel (note that vLLM does not have a paged prefill kernel).
-For a fair comparison, we also profiled each system to find its best performing configuration.
-Accordingly, we set KV-cache block size to 16 for both vLLM and FlashInfer, and 256 for FlashAttention.
-Using a higher block size for vLLM increases its kernel latency by up to 3×3$\times$3 × as shown in Figure 3, and using a smaller block size for FlashAttention paged kernel increases its latency by up to $9\%$ (Table 9).
-We find that the choice of page size does not affect vAttention as long as fragmentation is not a concern (Figure 11).
-Model Context Length FA_Paged FA_vAttention FI_Paged FI_vAttention
+We also profiled each system to find its best performing configuration.
+Accordingly, we set KV cache block size to 16 for both vLLM and FlashInfer, and 256 for FlashAttention-2.
+Using a higher block size for vLLM increases its kernel latency by up to 1.9×1.9$\times$1.9 × as shown in Figure 3, and using a smaller block size for FlashAttention-2 paged kernel increases its latency by up to $9\%$.
+We find that the choice of page size does not affect vAttention as long as fragmentation is not a concern.
+Model Context Length FlashAttention-2 FlashInfer
+Paged vAttention Paged vAttention
 Yi-6B 64K 10.6 (7.0) 9.1 (5.5) 10.9 (6.0) 9.1 (5.4)
@@
 Table 6. Prefill completion and attention (in parenthesis) time with different attention back-ends (unit: seconds).
-Figure 7. Decode throughput. FA_vAttention is on par with FA_Paged (note the overlapping lines) which is the best among all PagedAttention based alternatives, while outperforming FI_Paged and vLLM.
 7.1. Prefill Evaluation
-We evaluate four configs for prefill: FA_Paged, FI_Paged, FA_vAttention and FI_vAttention.
+We evaluate 4 configurations for prefill: FA2_Paged, FI_Paged, FA2_vAttention and FI_vAttention.
 Configurations with the “_Paged” suffix represent the PagedAttention-based kernel and those with “_vAttention” use the non-paged kernels of the respective library.
-Figure 6 shows the prefill throughput for Yi-6B, Llama-3-8B and Yi-34B.
-We summarize our key findings below.
-Small contexts: For small contexts, prefill cost is dominated by the linear operators i.e., attention’s contribution is relatively low (sarathi2023, ).
-Hence, even though vAttention helps speed up attention computation, the throughput of both paged and vAttention back-ends is nearly identical in FlashAttention.
+Figure 7 shows the prefill throughput for our models.
+We summarize key findings below.
+Small contexts: For small contexts, prefill cost is dominated by the linear operators, i.e., attention’s contribution is relatively low (Agrawal et al., 2023).
+Hence, even though vAttention helps speed up attention computation, the throughput of both paged and vAttention back-ends is nearly identical in FlashAttention-2.
 However, for FlashInfer, we find that using vAttention helps improve prefill throughput even for small contexts.
 This is because FlashInfer incurs various other sources of overhead in the paged version.
-First, appending a new K or V tensor to the KV-cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time (the copy operation has been optimized for FlashAttention by vLLM (vllm-copy-kernels, )).
+First, appending a new K or V tensor to the KV cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time (the copy operation has been optimized for FlashAttention-2 by vLLM (vll, 2024a)).
 Second, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration.
-vAttention avoids such overheads because it maintains KV-cache’s virtual contiguity, eliminating the need for a Block-Table.
+vAttention avoids such overheads because it maintains KV cache’s virtual contiguity, eliminating the need for a Block-Table.
+Figure 8. Decode throughput. FA2_vAttention is on par with FA2_Paged (note the overlapping lines) which is the best among all PagedAttention based alternatives, while outperforming FI_Paged and vLLM.
 Long contexts: The contribution of attention computation becomes significant at 16K and higher context lengths in our experiments.
-Therefore, even for FlashAttention back-end, vAttention outperforms the paged counterpart.
-For example, at context length 192K, FA_vAttention outperforms FA_Paged by 1.24×1.24$\times$1.24 ×, 1.26×1.26$\times$1.26 × and 1.24×1.24$\times$1.24 × for Yi-6B, Llama-3-8B and Yi-34B, respectively.
+Therefore, even for FlashAttention-2 back-end, vAttention outperforms the paged counterpart.
+For example, at context length 192K, FA2_vAttention outperforms FA2_Paged by 1.24×1.24$\times$1.24 ×, 1.26×1.26$\times$1.26 × and 1.24×1.24$\times$1.24 × for Yi-6B, Llama-3-8B and Yi-34B, respectively.
 Similarly, for FlashInfer back-ends, FI_vAttention improves prefill throughput by up to 1.25×1.25$\times$1.25 × and 1.36×1.36$\times$1.36 × for Yi-6B and Llama-3-8B (context length 16K), and 1.17×1.17$\times$1.17 × for Yi-34B (context length 32K).
-Attention time: vAttention’s improvement in prefill throughput is primarily due to faster attention kernels enabled by a (virtually) contiguous KV-cache.
-This is because the prefill phase of a long prompt is primarily dominated by attention computation, as can be observed by comparing the numbers inside parenthesis with the total prefill completion time in Table 5.
-For FlashAttention, nearly all the gains of vAttention are due to faster attention kernels e.g., prefill gains of 1.5 seconds, 7.4 seconds and 16.9 seconds for Yi-6B are almost entirely due to gains in attention computation.
+Attention time: vAttention’s improvement in prefill throughput is primarily due to faster attention kernels enabled by a (virtually) contiguous KV cache.
+This is because the prefill phase of a long prompt is primarily dominated by attention computation, as can be observed by comparing the numbers inside parenthesis with total prefill completion time in Table 6.
+For FlashAttention-2, nearly all the gains of vAttention are due to faster attention kernels, e.g., prefill gains of 1.5 seconds, 7.4 seconds and 16.9 seconds for Yi-6B are all due to gains in attention computation.
 vAttention enabled kernels also help with the FlashInfer back-end.
-In addition, FI_Paged also has other sources of overheads e.g., in the 14 seconds of total savings (Yi-34B, 192K context), only 7 seconds is due to attention and the rest is due to other sources.
-Figure 8. Makespan of serving a set of 50 requests with varying context lengths and the ratio of prefill to decode tokens.
-6.2. Decode Evaluation
-For decodes, in addition to FA_Paged and FI_Paged, we also evaluate the throughput obtained with vLLM’s decode kernel (the first ever kernel to support PagedAttention).
-For vAttention, we use FlashAttention’s non-paged kernel.
-Unfortunately FlashInfer’s non-paged decode kernel has significantly higher latency compared to all the other kernels we used in these experiments (up to 14.6×14.6$\times$14.6 ×, Table 10).
-Hence, while vAttention supports dynamic memory allocation for FlashInfer’s non-paged decode kernel, we omit it for evaluation in this section (note that vAttention is responsible only for allocating memory – it cannot make a slow kernel fast).
-Figure 7 shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch size up to 32 (except for Yi-34B which runs out of memory for batch size 32).
-We set the initial context length of each request to 16K tokens and calculate decode throughput based on the mean latency of 400 decode iterations (see Figure 7).
-First, vAttention is on par with the best of PagedAttention as shown by FA_Paged and FA_vAttention in Figure 7.
-In comparison, FI_Paged has somewhat lower throughput and vLLM is the worst for all models and configurations.
-For example, FA_Paged and FA_vAttention outperform vLLM by up to 1.99×1.$9$, 1.58×1.58$\times$1.58 × and 1.53×1.53$\times$1.53 × for Yi-6B, Llama-3-8B and Yi-34B, respectively.
-The primary reason is that vLLM’s decode kernel has significantly higher latency than the FlashAttention based kernels; while FlashAttention has continuously adopted new optimizations (e.g., FlashDecoding (flashdecoding, )), vLLM has lagged behind.
-For example, Table 6 shows that vLLM’s PagedAttention kernel incurs up to 2.8×2.8$\times$2.8 ×, 1.5×1.5$\times$1.5 ×, and 2.5×2.5$\times$2.5 × higher latency than FlashAttention kernels for Yi-6B, Llama-3-8B and Yi-34B.
-This is despite vLLM being in an actively maintained open-source serving stack, as well as being used by various companies for serving LLMs.
-This is an important result that underlines the importance of adopting new optimizations.
-Model BS vLLM FA_Paged FI_Paged FA_vAttention
+In addition, FI_Paged also has other sources of overheads, e.g., in the 14 seconds of total savings (Yi-34B, 192K context), only 7 seconds is due to attention and the rest is due to other sources.
+Model BS vLLM FA2_Paged FI_Paged FA2_vAttention
 Yi-6B 16 32.3 11.5 15.2 11.3
@@
 Table 7. Total latency of attention kernel (sum of all layers) per decode iteration (in milliseconds, BS = batch size).
-Second, relative gains of FA_vAttention and FA_Paged increase over vLLM with the batch size e.g., as batch size increases from 4 to 32 for Llama-3-8B, relative gains increase from 1.05×1.05$\times$1.05 × to 1.58×1.58$\times$1.58 ×.
-This is because the latency of a decode attention kernel is proportional to the total number of tokens in the batch (vidur, ).
-Therefore, the contribution of attention kernel in the overall latency – and consequently gains with a more efficient kernel – increase with the batch size.
-Further, for the same reasons as discussed in §6.1 (faster attention and lower CPU overhead), FA_vAttention delivers up to 1.23×1.23$\times$1.23 × higher throughput than FI_Paged (Yi-6B, batch size 12).
+7.2. Decode Evaluation
+For decodes, in addition to FA2_Paged and FI_Paged, we also evaluate the throughput obtained with vLLM’s decode kernel (the first ever kernel to support PagedAttention).
+For vAttention, we use FlashAttention-2’s non-paged kernel.
+Unfortunately FlashInfer’s non-paged decode kernel has significantly higher latency (up to 14.6×14.6$\times$14.6 ×) compared to all these other kernels.
+Hence, while vAttention can support dynamic memory allocation for FlashInfer’s non-paged decode kernel, we omit it for evaluation in this section.
+Figure 8 shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch size up to 32 (except for Yi-34B which runs out of memory for batch size 32).
+We set the initial context length of each request to 16K tokens and calculate decode throughput based on the mean latency of 400 decode iterations.
+First, vAttention is on par with the best of PagedAttention as shown by FA2_Paged and FA2_vAttention in Figure 8.
+In comparison, FI_Paged has somewhat lower throughput and vLLM is the worst for all models and configurations.
+For example, FA2_Paged and FA2_vAttention outperform vLLM by up to 1.99×1.$9$, 1.58×1.58$\times$1.58 × and 1.53×1.53$\times$1.53 × for Yi-6B, Llama-3-8B and Yi-34B, respectively.
+The primary reason is that vLLM’s decode kernel has significantly higher latency than the FlashAttention-2 based kernels; while FlashAttention-2 has continuously adopted new optimizations (e.g., FlashDecoding (fla, 2023)), vLLM has lagged behind.
+For example, Table 7 shows that vLLM’s PagedAttention kernel incurs up to 2.8×2.8$\times$2.8 ×, 1.5×1.5$\times$1.5 ×, and 2.5×2.5$\times$2.5 × higher latency than FlashAttention-2 kernels for Yi-6B, Llama-3-8B and Yi-34B.
+This is despite vLLM being in an actively maintained open-source serving stack, as well as being used by various companies for serving LLMs.
+This is an important result that underlines the importance of low software complexity and portability.
+Second, relative gains of FA2_vAttention and FA2_Paged increase over vLLM with the batch size, e.g., as batch size increases from 4 to 32 for Llama-3-8B, relative gains increase from 1.05×1.05$\times$1.05 × to 1.58×1.58$\times$1.58 ×.
+This is because the latency of a decode attention kernel is proportional to the total number of tokens in the batch (Agrawal et al., 2024a).
+Therefore, the contribution of attention kernel in the overall latency – hence gains with a more efficient kernel – increase with the batch size.
+Further, for the same reasons as discussed in §7.1 (faster attention and lower CPU overhead), FA2_vAttention delivers up to 1.23×1.23$\times$1.23 × higher throughput than FI_Paged (Yi-6B, batch size 12).
 Finally, note that vAttention is only as good as the state-of-the-art PagedAttention for decode throughput, as compared to prefills where it outperforms PagedAttention.
-We believe this is due to memory bound nature of decode attention i.e., memory stalls make it possible to hide the effect of additional compute that paging support requires.
-This is evident from Table 6 wherein FA_Paged and FA_vAttention have similar latency.
-However, hiding compute overhead in a prefill attention kernel is hard because it is already compute bound (see Table 5 for latency gains in the prefill attention kernel).
-6.3. End-to-end Throughput Evaluation
-For end-to-end performance, we measure the makespan of serving a static trace of 50 requests for long context scenarios.
-To understand system performance for different workloads, we vary the initial context length of requests from 32K to 128K, and the ratio of prefill to decode tokens (referred as P:D) from 500 to 50.
-Note that a higher P:D as well as a longer context indicates that the workload is more prefill bound; we expect vAttention to improve performance in such cases.
-Figure 8 shows the makespan for all the three models.
-For Yi-6B, FA_vAttention is up to 1.22×1.$2$, 1.16×1.16$\times$1.16 × and 1.13×1.13$\times$1.13 × faster than FA_Paged for P:D ratio of 500, 100 and 50.
-FA_vAttention is also up to 1.19×1.19$\times$1.19 × faster for Llama-3-8B and up to 1.16×1.16$\times$1.16 × faster for Yi-34B, compared to FA_Paged.
-Our gains over FA_Paged are higher at higher P:D ratios, and for a fixed P:D, our gains increase as the context length grows.
-These results are consistent with our expectation since a higher P:D and context lengths both mean that the system spends more time processing prefills where use of non-paged attention kernels outperform paged kernels.
-Interestingly, we observe that despite being faster than FA_Paged in prefills, FI_Paged under-performs FA_Paged in end-to-end throughput in most cases (due to inefficient decodes as shown in Figure 7).
-Hence, our gains over FI_Paged are relatively higher (up to 1.29×1.29$\times$1.29 ×), except for some cases where FI_Paged outperforms FA_Paged (e.g., Yi-6B and Yi-34B, context length=128K, P:D=500).
+This is because in case of decode attention, paged and non-paged kernels have similar latency as shown in Table 7.
+We believe this is due to memory bound nature of decode attention, i.e., memory stalls make it possible to hide the effect of additional compute that paging support requires.
+However, hiding compute overhead in a prefill attention kernel is hard because it is already compute bound (see Table 6 for vAttention’s gains in the prefill kernel).
+Figure 9. Offline inference throughput while serving long-context requests from the arXiv-Summarization dataset (arx, 2024).
+7.3. End-to-end Performance: Offline Scenarios
+We measure the end-to-end system throughput in an offline scenario as the number of requests completed per minute.
+We evaluate a total of 427 long-context requests from the arXiv-Summarization workload trace wherein the total context length per-request varies from 64K tokens to 192K tokens and the number of output tokens per-request varies from 17 to 5153.
+The mean prefill to decode token ratio is 356.
+Figure 9 shows throughput for different attention backends for all three models.
+FA2_vAttention outperforms FA2_Paged by 1.18×1.18$\times$1.18 ×, 1.15×1.15$\times$1.15 × and 1.13×1.13$\times$1.13 ×, and FI_Paged by 1.19×1.19$\times$1.19 ×, 1.23×1.23$\times$1.23 ×, and 1.14×1.14$\times$1.14 × for Yi-6B, Llama-3-8B and Yi-34B, respectively.
+In general, vAttention’s performance gains are proportional to the context length and the prefill to decode token ratio (P:D ratio); these factors determine how much prefill attention contributes to the total runtime.
+A higher P:D ratio as well as longer context lengths indicate that the workload is more prefill bound.
+vAttention provides higher gains in such cases.
+7.4. End-to-end Performance: Online Scenarios
+We evaluate an online inference scenario serving long-context requests from arXiv-Summarization (arx, 2024).
+In total, we run 512 requests wherein the per-request input context length varies from 22K to 45K tokens (mean 29K), the number of decode tokens varies from 6 to 3250 (mean 348) and the mean P:D ratio is 129.
+We evaluate performance near the serving capacity of the system which denotes the maximum load a system can handle without incurring high queuing delays (Kwon et al., 2023; Agrawal et al., 2024b).
+We vary input load (queries-per-second or QPS) based on Poisson distribution and schedule requests in first-come-first-serve order.
+Figure 10 shows the CDF of the end-to-end request execution latency for this experiment.
+We see that vAttention consistently outperforms both baselines.
+For example, FA2_vAttention reduces the median request execution latency over FA2_Paged by up to 42% (QPS 0.25) for Yi-6B, by up to 28% (QPS 0.3) for Llama-3-8B and by up to 29% for Yi-34B (QPS 0.1).
+The primary reason for vAttention’s performance gain is that it can compute the prefill phase of new requests faster than the PagedAttention-based systems which substantially reduces queuing delays.
+Consistent with our prior results, FI_Paged is slower than FA2_Paged and hence our gains our FI_Paged are relatively higher compared to our gains over FA2_Paged.
+Figure 10. CDF of end-to-end request execution latency in online inference under varying load.
+Figure 11. Offline inference throughput on H100 GPUs.
+7.5. Exemplifying Portability: FlashAttention-3
+We demonstrate the portability benefit of vAttention with the recently released FA3 kernel (Shah et al., 2024).
+FA3 is optimized for the NVIDIA Hopper architecture and did not support PagedAttention when released.
+Therefore, dynamic memory allocation via PagedAttention is not feasible for FA3 at the time of writing this paper (integration of FA3 into vLLM is also a work in progress (fa3, 2024a, b)).
+vAttention not only enables dynamic memory allocation with FA3, it also requires no code changes to deploy FA3.
+Figure 11 shows the offline inference throughput of our models with 1–2 H100 GPUs (Yi-6B deployed on a single GPU and the other models on two GPUs each) on the same arXiv-Summarization-based workload as evaluated in §7.3.
+FA3 with vAttention provides an additional speedup of up to 1.35×1.35$\times$1.35 × (Yi-6B) over FA2_vAttention which is already up to 1.15×1.15$\times$1.15 × faster than FA2_Paged.
+7.6. Ablation Studies
 Figure 12. Latency of decode iterations with and without overlapping memory allocation with compute (batch size=32, context length=4K–8K, model: Llama-3-8B)
-6.4. Ablation Studies
 7.6.1. Hiding allocation latency
-Figure 9 shows the latency of more than 2500 decode iterations of Llama-3-8B.
-For this experiment, we used a batch of 32 requests and initialized the prefill context length of each request to be in between 4K–8K (chosen randomly).
-It is evident that overlapping memory allocation with compute effectively hides the latency of allocating memory via CUDA APIs.
-We used 2MB pages for this experiment to show that even the worst case latency can be hidden (Table 2 shows that 2MB pages incur highest latency).
-In contrast, allocating memory synchronously via CUDA APIs leads to frequent latency spikes of 5ms–20ms, depending on how many requests need new physical memory pages.
+Figure 12 shows the latency of more than 500 decode iterations of Llama-3-8B.
+For this experiment, we used a batch of 32 requests with per request prefill context length varying between 2K to 3K tokens.
+It is evident that overlapping memory allocation with compute effectively hides the latency of CUDA VMM APIs.
+We used 2MB pages for this experiment to show that even the worst case memory allocation latency can be hidden by overlapping it with compute (Table 3 shows that 2MB pages incur highest latency).
+In contrast, allocating memory synchronously via CUDA APIs leads to frequent latency spikes of 5 to 15 milliseconds, depending on how many requests require physical memory allocation in a given iteration.
 7.6.2. Deferred reclamation
-Figure 10 shows that synchronous memory allocation using CUDA APIs for prefills incurs overhead of up to 1.15×1.15$\times$1.15 × with 64KB pages and up to 1.03×1.03$\times$1.03 × with 2MB pages.
-In most cases, deferred reclamation eliminates the need to invoke CUDA APIs for prefills because a newly arrived request can simply re-use physical memory pages that were allocated to a prior request.
+Figure 13 shows that synchronous memory allocation incurs prefill overhead of up to 1.15×1.15$\times$1.15 × with 64KB pages and up to 1.03×1.03$\times$1.03 × with 2MB pages.
+In most cases, deferred reclamation eliminates the need to invoke CUDA VMM APIs for prefills because a newly arrived request can simply re-use physical memory allocated to a prior request.
 This way, deferred reclamation ensures that prefill latency is not affected by memory allocation.
-6.4.3. Memory allocation bandwidth
-Table 7 shows that even with 64KB (our smallest) page size, vAttention can allocate as much as 7.6GB per second per GPU.
-This is more than an order of magnitude higher than the maximum memory allocation rate of 600MB per second of decodes (Figure 4).
-Larger page sizes and higher TP dimensions increase the memory allocation rate proportionally.
-Therefore, memory allocation bandwidth of CUDA APIs is more than enough for LLM inference.
 Figure 13. Prefill completion time of a single prompt of 16K tokens with different memory allocation strategies.
+Figure 14. Effect of page size on the runtime of FlashAttention-2’s prefill (left) and decode (right) attention kernels (model: Llama-3-8B).
+7.6.3. Effect of page size.
+In many applications, use of smaller pages can potentially degrade performance due to TLB thrashing (Shin et al., 2018; Kwon et al., 2016; Panwar et al., 2019; Pichai et al., 2014).
+We find that this is not the case for LLM inference.
+For example, Figure 14 shows that the execution latency of an attention kernel remains largely unaffected when the KV cache is allocated using 64KB pages, compared to using 2MB pages.
+In a separate experiment, we find that these results are also consistent with very large models, e.g., Llama-3-70B and GPT-3-175B.
+We attribute this to the regular computation pattern of the attention operator as well as the hand-tuned implementations that explicitly try to avoid irregular memory accesses.
+Model Block Size (# Tokens in a page-group)
+64KB 128KB 256KB 2MB
+Yi-6B (TP-1) 64 128 256 2048
+Yi-6B (TP-2) 128 256 512 4096
+Llama-3-8B (TP-1) 32 64 128 1024
+Llama-3-8B (TP-2) 64 128 256 2048
+Yi-34B (TP-1) 32 64 128 1024
+Yi-34B (TP-2) 64 128 256 2048
+Table 8. KV cache block size as a function of page-group size and degree of tensor parallelism.
+Table 8 shows the block size i.e., number of tokens per physical page-group.
+Smaller page-groups enable fine-grained allocation, approaching vLLM’s recommended block size of 16–32 (the minimum block size in FlashAttention-2 is 256 – higher than vAttention).
+This shows that vAttention is as effective in reducing fragmentation as PagedAttention.
+In terms of end-to-end system performance, we find that 2MB pages are good enough for many online serving scenarios where latency constraints prevent use of very high batch sizes.
+In contrast, smaller page-groups are better for throughput-oriented scenarios where maximizing batch size is important to obtain peak performance.
+For example, using 2MB pages could serve batch sizes of up to 187, 203 and 56 for Yi-6B, Llama-3-8B and Yi-34B on a dynamic trace (dataset: OpenChat (Wang et al., 2023), load: 7 queries per seconds).
+In contrast, 64KB pages helped serve batch sizes of up to 240, 258 and 68 for our models as shown in Figure 15.
+Figure 15. Maximum batch size obtained with different page-group sizes for a dynamic workload.
 Config. 64KB 128KB 256KB 2MB
@@
 Table 9. Physical memory allocation bandwidth (GB per second) with varying allocation granularity.
-6.4.4. Effect of page size.
-In many applications, use of smaller pages can potentially degrade performance due to TLB thrashing (neighbourhoodawareat, ; ingens, ; hawkeye, ; architecturalsupportforgpuat, ).
-We find that this is not the case for LLM inference.
-For example, Figure 11 shows that the execution latency of an attention kernel remains largely unaffected when the page size for KV-cache is reduced to 64KB, from the original 2MB.
-In a separate experiment, we find that these results are also consistent with very large models e.g., Llama-3-70B and GPT-3-175B.
-We attribute this to the regular computation pattern of the attention operator as well as the hand-tuned implementations that explicitly try to avoid irregular memory accesses.
-Further, we find that 2MB pages are good enough for many online serving scenarios where latency constraints limit the maximum batch size.
-However, for throughput-oriented scenarios, smaller pages may be preferred.
-For example, using 2MB pages could serve batch sizes of up to 187, 203 and 56 for Yi-6B, Llama-3-8B and Yi-34B on a dynamic trace (dataset: OpenChat (wang2023openchat, ), load: 7 queries per seconds).
-In contrast, 64KB pages helped serve batch sizes of up to 240, 258 and 68 for our models (see Figure 12).
-Figure 11. Effect of page size on the performance of FlashAttention’s prefill (left) and decode (right) attention kernels (model: Llama-3-8B).
-Figure 12. Maximum batch size obtained with different page sizes for a dynamic workload. Large pages limit batch size due to internal fragmentation.
-Figure 13. Illustration of code changes needed to replace the prefill attention kernel of FlashAttention by FlashInfer when using vAttention for dynamic physical memory allocation.
-Model # Tokens in a physical memory block Max memory waste per request
-64KB 128KB 256KB 2MB 64KB 128KB 256KB 2MB
-Yi-6B (TP-1) 64 128 256 2048 4MB 8MB 16MB 128MB
-Yi-6B (TP-2) 128 256 512 4096 8MB 16MB 32MB 256MB
-[…]
-Llama-3-8B (TP-2) 64 128 256 2048 8MB 16MB 32MB 256MB
-[…]
-Yi-34B (TP-2) 64 128 256 2048 15MB 30MB 60MB 480MB
-Table 8.
-Block size (number of tokens in a physical memory page) as a function of the page size and degree of tensor parallelism.
-Columns on the right show how much physical memory can be wasted per-request in the worst-case.
+7.6.4. Memory allocation bandwidth
+Table 9 shows that even with 64KB pages (our smallest), vAttention can allocate as much as 7.6GB physical memory per second per GPU.
+This is more than an order of magnitude higher than the maximum memory allocation rate of 750MB per second of decodes (Figure 4).
+Larger page-group sizes and higher TP dimensions increase the memory allocation rate proportionally.
+Therefore, memory allocation bandwidth of CUDA VMM APIs is more than sufficient for LLM inference.
+8. Discussion
+8.1. Managing KV cache via Unified Memory
+To enable dynamic memory allocation, we also considered leveraging unified memory via cudaMallocManaged (NVIDIA, 2024; Nayak et al., 2021).
+However, we find that unified memory support is currently not suitable for serving LLMs.
+First, it does not support partial freeing, preventing reclamation of physical memory of individual requests.
+Second, it lacks support for memory aliasing which prevents de-duplication of KV cache content in physical memory (de-duplication is useful when requests share a common prefix (Kwon et al., 2023)), consequently limiting batch size.
+cudaMallocManaged also allocates 2MB pages by default which can cause severe fragmentation.
+However, our changes in NVIDIA drivers are based on unified memory: we allocate virtual tensors using cudaMallocManaged, enabling support for partial freeing and page sharing with additional APIs.
+In addition, we introduce latency hiding optimizations and support for smaller pages.
+Therefore, one could consider our extensions to NVIDIA drivers as “unified memory optimized for LLM serving”.
+Model Block size (# Tokens in a 2MB page)
+w/o Tensor Slicing w/ Tensor Slicing
+Yi-6B (TP-1) 2048 64
+Yi-6B (TP-2) 4096 128
+Llama-3-8B (TP-1) 1024 32
+Llama-3-8B (TP-2) 2048 64
+Yi-34B (TP-1) 1024 18
+Yi-34B (TP-2) 2048 36
+Table 10. KV cache block size with and without tensor slicing with 2MB pages.
+8.2. Reducing Fragmentation via Tensor Slicing
+To reduce fragmentation caused by 2MB pages, we also provide an alternative method that does not require modifying NVIDIA drivers.
+In this method, we use a single 2MB page to store the KV cache tokens of all layers for a given request.
+This can be done by allocating one virtual tensor of shape $[B,L,N,H,D]$ for K cache (and one for V cache) and slicing it across all layers, instead of allocating $2\times N$ virtual tensors of shape $[B,L,H,D]$.
+This reduces fragmentation to $1/N$ of the prior design as shown in Table 10.
+In this design, the K cache (or V cache) of a particular layer $n$ is represented by slice $[B,L,n:n+1,H,D]$ of the original tensor, and can be passed to attention kernels for computation.
+However, note that with tensor slicing, the K cache (and V cache) of each layer is no longer contiguous.
+Computing attention over such a tensor slice is possible only if the the attention kernel supports memory addressing with strides.
+While many kernels (e.g., FlashAttention-2) support strides out-of-the-box, the earlier versions of FlashInfer lacked such support (fi-, 2024).
+Therefore, relying on tensor slicing as the primary method of reducing fragmentation would have prevented vAttention from supporting FlashInfer kernels.
+Hence, to support unmodified FlashInfer kernels, we chose to add support for smaller pages in NVIDIA drivers.
+Importantly, our two solutions are compatible with each other; one could deploy smaller pages with tensor slicing to further reduce fragmentation, if required.
 8.3. Programming Effort
-vAttention makes it feasible to replace one attention kernel with another with only a few lines of code changes in the serving framework (Figure 13).
-In contrast, in PagedAttention, a developer first needs to write a paged kernel and then make significant changes in the serving framework.
-For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes in over 15 files (vllmfi1, ; vllmfi2, ; vllmfi3, ).
-Implementing the initial paging support in FlashAttention kernel also required $\approx 280$ lines of code changes (fapaging, ) and additional efforts to enable support for smaller block sizes (fasmall, ).
-7. Discussion
-Batch Size FA_Paged (Block Size = 256) FA_Paged (Block Size = 16)
-Model Ratio
-Yi-6B 1 0.085 0.085 1.00×1.$0$
-2 0.093 0.094 1.01×1.01$\times$1.01 ×
-4 0.109 0.115 1.05×1.05$\times$1.05 ×
-[…]
-2 0.097 0.105 1.07×1.07$\times$1.07 ×
-4 0.136 0.143 1.05×1.05$\times$1.05 ×
-[…]
-2 0.088 0.095 1.08×1.08$\times$1.08 ×
-4 0.116 0.117 1.01×1.01$\times$1.01 ×
-Table 9. Impact of varying block size on the execution latency (in milliseconds) of FlashAttention decode kernel.
-7.1. Further Analysis of Internal Fragmentation
-vAttention uses page size as the unit of physical memory allocation as opposed to the PagedAttention’s unit of KV-cache block size.
-Therefore, it is natural to ask what is the relationship between page size and block size.
-To compute this relationship, let $S$ be the size of a single token’s per-layer K-cache (or V-cache) on a worker (see §5.1).
-Then, page size t translates to KV-cache block size $=t/S$.
-Note that $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16).
-Therefore, block size = $t/(H\times D\times P)$.
-Table 8 shows the block size for different model configurations and physical memory allocation sizes; we show each model under two TP configurations – TP-1 and TP-2 – to highlight the effect of TP dimension on block size.
-In addition, the table also shows how much physical memory can be (theoretically) wasted by vAttention due to internal fragmentation in the worst-case.
-The worst-case occurs when a new page is allocated but remains completely unused.
-Context Length Batch Size FI_NonPaged FI_Paged Ratio
-8K 4 0.902 0.071 12.7×12.7$\times$12.7 ×
-8K 8 0.931 0.128 7.3×7.3$\times$7.3 ×
-8K 16 0.943 0.229 4.2×4.2$\times$4.2 ×
-[…]
-16K 8 1.821 0.233 7.8×7.8$\times$7.8 ×
-16K 16 1.875 0.450 4.2×4.2$\times$4.2 ×
-[…]
-32K 8 3.291 0.458 7.2×7.2$\times$7.2 ×
-32K 16 3.365 0.891 3.8×3.8$\times$3.8 ×
-Table 10.
-Latency (in milliseconds) of paged and non-paged decode attention kernels in FlashInfer.
-Ratio denotes the latency of non-paged kernel divided by the latency of paged kernel (model: Llama-3-8B, 2 A100 GPUs).
-vAttention allocates physical memory equivalent to the page size on each TP worker whereas the per-token physical memory requirement of a worker goes down as TP dimension increases (because KV heads get split across TP workers).
-Therefore, block size increases proportionally with the TP dimension.
-Table 8 shows that using 2MB pages leads to large KV-cache block sizes of 1024 (Llama-3-8B and Yi-34B, TP-1) to 4096 (Yi-6B, TP-2) which could waste significant amount of physical memory (100s of MBs, per-request).
-Use of smaller (e.g., 64KB) pages reduces internal fragmentation proportionally, resulting in KV-cache block size of 32 (Yi-34B and Llama-3-8B, TP-1) to 128 (Yi-6B, TP-2), which in turn limit the maximum theoretical waste of only 4-15MB physical memory, per request.
-This shows that controlling the granularity of physical memory allocation is essential for reducing fragmentation.
-If required, page size can be reduced further to as low as 4KB which is the minimum page size supported in almost all architectures today, including NVIDIA GPUs.
-7.2. Effect of Block Size on FlashAttention Kernel
-We find that varying the block size of KV-cache also affects the execution latency of FlashAttention’s decode kernel (in addition to vLLM wherein using a larger block size increase the latency of decode kernel by up to 3×3$\times$3 × as shown in Figure 3).
-For example, Table 9 shows that using a smaller block size of 16 increases the latency of FlashAttention decode kernel by up to $5\%$ for Yi-6B, up to $8\%$ for Llama-3-8B and up to $9\%$ for Yi-34B.
-This is likely because work distribution and shared memory usage in a paged kernel needs to explicitly account for block size.
-Hence, writing an implementation that is equally efficient for all block sizes could be non-trivial.
-In fact, it is interesting to see that vLLM decode kernel performs better with small block sizes whereas FlashAttention decode kernel performs better with large block sizes.
-While the impact of block size on kernel latency is much less pronounced in FlashAttention compared to vLLM, a degradation in performance is still undesirable for an operator as important as attention.
-vAttention is free of such implementation challenges because it does not require a block table.
-7.3. FlashInfer Non-paged Decode Attention Kernel
-Note that we did not include FlashInfer non-paged decode kernel in our evaluation (§6).
-There are two reasons for this.
-First, the non-paged kernel has a much higher latency than the paged counterpart of FlashInfer e.g., Table 10 shows that the latency of non-paged kernel is up to 14.6×14.6$\times$14.6 × higher than the paged kernel.
-Further, the latency of the paged kernel follows an expected trend i.e., increasing the context length or batch size leads to a proportional increase in the execution latency.
-In contrast, increasing the batch size does not affect the latency of the non-paged kernel (e.g., context lengths of 8K and 32K).
-This behavior of the non-paged kernel is counter-intuitive and likely due to un-optimized implementation.
-Note that the job of a memory allocator is to simply allocate memory – it cannot accelerate a kernel.
-Second, the non-paged decode kernel of FlashInfer does not support variable sequence lengths – a fundamental requirement for LLM inference.
+PagedAttention requires significant programming effort.
+For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes in over 15 files (vll, 2024c, e, d).
+Implementing the initial paging support in FlashAttention-2 kernel also required $\approx 280$ lines of code changes (fap, 2024) and additional efforts to support small block sizes (fas, 2024a).
+In contrast, vAttention makes it feasible to replace one attention kernel with another with only a few lines of code changes in the serving framework (Figure 16).
+Note that the example shown in Figure 16 has no interaction with memory management; this is precisely the goal of vAttention, i.e., when a slow kernel is replaced by a fast kernel, memory management should continue to work transparently.
+Figure 16. Illustration of code changes needed to replace the prefill attention kernel of FlashAttention-2 by FlashInfer when using vAttention for memory allocation.
 9. Related Work
-Optimizing LLM inference is an active area of research.
-Various techniques have been proposed to improve different aspects of LLM serving like batching (orca, ; vllmsosp, ; sarathiserve2024, ), disaggregation (patel2023splitwise, ; distserve2024, ; tetriinfer, ), scheduling (flexgen, ; fastserve, ).
-However, central to all these techniques is the need for efficient KV-cache memory management.
-Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trtllmgithub, ), LightLLM (lightllm:github, ), and libraries e.g., FlashAttention (fagithub, ) and FlashInfer (figithub, ).
-In contrast, vAttention offers an alternate – which we believe is also a more principled – approach to dynamic KV-cache memory management.
-In a recent work, GMLake (gmlakeasplos24, ) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size.
-In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors for large object allocations.
-In contrast, vAttention is focused on avoiding fragmentation for LLM inference.
-Different from training, LLM inference is latency sensitive and requires smaller granularity allocations.
-We proposed various LLM inference specific optimizations to meet these requirements.
-Similar to our approach, PyTorch also recently added support for expandable segments (expandable-segments, ) that dynamically attach physical memory pages to a pre-reserved virtual memory buffer, if enabled.
-However, PyTorch uses expandable segments opportunistically, using synchronous memory allocation, and allocates physical memory pages only in multiples of 2MB granularity.
-In contrast, vAttention caters to the specific requirements of the KV-cache.
-Some other concurrent works are also motivated to optimize attention kernels (flashattnhopper, ; mirage-tensor-compiler, ; thunderkittens, ; heavyhitters, ; flashdecoding++, ; leanattention, ; flash-attention-3, ).
-As one would expect, new research ideas do not start with paging support in mind.
-vAttention would make it easier to deploy them.
+Optimizing LLM inference is an active area of research  (Mei et al., 2024; Griggs et al., 2024; Miao et al., 2023; Zhu et al., 2024; Kossmann et al., 2024; Strati et al., 2024; Lee et al., 2024; Singhania et al., 2024; Zheng et al., 2024; Ma et al., 2024; Mahmud et al., 2024).
+Various techniques have been proposed to improve different aspects of LLM serving like batching (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024b), disaggregation (Patel et al., 2024; Zhong et al., 2024; Hu et al., 2024), scheduling (Sheng et al., 2023; Wu et al., 2023).
+However, central to all these techniques is the need for efficient KV cache memory management.
+Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trt, 2023), LightLLM (lig, 2023), and libraries e.g., FlashAttention-2 (fag, 2022) and FlashInfer (fig, 2023).
+Some other concurrent works have also proposed optimizations for attention kernels (Bikshandi and Shah, 2023; Wu et al., 2024; thu, 2024; Zhang et al., 2023b; Hong et al., 2024; Sanovar et al., 2024; Shah et al., 2024; Kamath et al., 2024).
+Deploying new kernels for inference is hard with PagedAttention.
+vAttention offers an alternate – which we believe is also a more principled – approach to dynamic KV cache memory management that makes it easier to deploy new kernels.
+In a recent work, GMLake (Guo et al., 2024) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size.
+In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors.
+In contrast, vAttention targets optimizing inference workloads.
 10. Conclusion
-In this paper, we propose vAttention for dynamic memory management in LLM serving systems.
-The key highlight of vAttention is that it leverages system support for demand paging instead of implementing it in user space.
-We present various examples to show that the vAttention approach reduces programming burden while improving portability and performance compared to the popular PagedAttention approach that many LLM serving systems use today.
+In this paper, we propose vAttention for dynamic KV cache memory management in LLM serving systems.
+The key highlight of vAttention is that it mitigates fragmentation in physical memory while retaining the contiguity of KV cache in virtual memory.
+We present various examples to highlight that vAttention reduces programming burden while improving portability and performance compared to the popular PagedAttention approach.
+Acknowledgements.
+We thank our shepherd Sam Ainsworth, along with the anonymous ASPLOS reviewers, for their valuable feedback.
+We thank NVIDIA for useful discussions on CUDA VMM APIs and for sharing their experiences with vAttention.
+We also thank members of the open-source community (GitHub user names izhuhaoran and tongping) for trying to upstream the vAttention approach to vLLM (vat, 2024b, a), and Zihao Ye for clarifying the role of register spilling in the runtime overhead of PagedAttention.
+Ajay Nayak is supported by the Prime Minister’s Fellowship Scheme for Doctoral Research, co-sponsored by the Confederation of Indian Industry, the Government of India, and Microsoft Research India.
+Appendix A Artifact Appendix
+A.1. Abstract
+vAttention is a simpler, portable and performant alternative to PagedAttention for dynamic memory management in LLM serving systems.
+This artifact contains instructions to install vAttention and scripts to reproduce key results of our paper (Figures 2, 3, 4, 6, 7, 8, 9, 11).
+The scripts are configured to run three large language models: Yi-6B, Llama-3-8B and Yi-34B.
+A.2. Artifact check-list (meta-information)
+  • Algorithm: Dynamic memory management for serving LLMs.
+  • Run-time environment: CUDA 12.1, Python 3.10, PyTorch 2.3.0.
+  • Hardware: Two NVLink connected NVIDIA A100 GPUs with 80GB memory each.
+  • Disk space required: 200GB.
+  • Time needed to prepare workflow: 1 hour.
+  • Time needed to complete experiments: 1-2 days.
+  • Publicly available?: Yes
+  • Archived DOI?: https://doi.org/10.5281/zenodo.14048692
+A.3. Description
+A.3.1. How to access
+Clone the artifact as follows:
+     \$ git clone https://github.com/microsoft/vattention
+A.3.2. Hardware dependencies
+Running Yi-6B requires one NVIDIA A100 GPU with 80GB memory while the other two models require two NVLink-connected A100 GPUs with 80GB memory each.
+A.3.3. Software dependencies
+Requires PyTorch 2.3.0 and CUDA 12.1 (later CUDA versions may or may not work).
+A.3.4. Expected runtime
+Figures 2, 3, and 11 should take a few minutes each. Figures 4, 7 should take about 1 hour each. Figure 6 requires approximately 3 hours. Figures 8 and 9 may consume more than 12 hours each.
+A.4. Installation
+Move to the directory containing artifact scripts, create and activate a conda environment, then install the artifact as follows:
+     \$ cd vattention/scripts/artifact˙asplos25/  \$ conda create -n vattn python=3.10  \$ conda activate vattn  \$ (vattn) ./install.sh
+vattention/scripts/artifact_asplos25/README.md file provides detailed installation instructions.
+A.5. Alternate Setup: Docker Image
+We also provide a docker image for vAttention with all its dependencies pre-installed.
+To access the docker image, you need to have Docker and NVIDIA Docker installed on your system.
+You can then launch the docker container and navigate to the folder containing vAttention artifact, as follows:
+     \$ docker run –gpus all -it “    -p 8181:8181 –rm –ipc=host –cap-add=SYS˙ADMIN “    rnp1910/vattention:asplos˙25˙pytorch˙run  \$ cd /workspace/vattention/scripts/artifact˙asplos25
+Then follow the experiment workflow detailed in A.7
+A.6. Accessing Llama-3-8B
+Accessing Llama-3-8B requires logging into huggingface with the user’s private token (HF_TOKEN below). Login as follows before any experiment:
+     \$ (vattn) huggingface-cli login –token HF˙TOKEN
+A.7. Experiment workflow
+You can launch all experiments at once or individually as:
+     \$ (vattn) ./run˙all.sh            OR  \$ (vattn) ./run˙figure˙2.sh  \$ (vattn) ./run˙figure˙3.sh
+A.8. Evaluation and expected results
+The raw output logs and the final plots will be redirected to ./logs.
+and ./plots/ subdirectories within the main artifact directory vattention/scripts/artifact_asplos25/.
+The plots are in the same format as the paper and can be compared directly.
+The expected result is that non-paged or vAttention based configurations would perform better than the PagedAttention counterparts, especially at longer context lengths.
+However, the exact numbers may differ from the paper depending on GPUs and software libraries installed on the experiment system.
+A.9. Experiment customization
+Reproducing Figures 8 and 9 of the paper requires running a large number of requests and hence these experiments can take more than 24 hours to complete.
+To reduce the experimental time, we have configured their scripts to run with a smaller number of requests (100 each) by default.
+If you want to run the full trace, execute these scripts with the --full argument, i.e., ./run_figure_8.sh --full and ./run_figure_9.sh --full.
+If two NVLink connected GPUs aren’t available, update run scripts to avoid running Llama-3-8B and Yi-34B.
+You can also configure Llama-3-8B to run on a single GPU by setting --model_tensor_parallel_degree to 1 in the ./helpers/common.sh file.
+Note that such changes would likely impact the absolute performance numbers considerably but vAttention is still expected to perform better than PagedAttention.